A Complete Bibliography of Publications in Mathematical Medicine and Biology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 April 2025 Version 1.12

Title word cross-reference

1 [BKW^{+}14]. 2 [GL15, GH05, TC06]. 3 [BKF14, WS06]. 4 [WDD^{+}17]. ^{11} [ACG^{+}15]. _{2} [EHE^{+}18, SSA89]. α [HVPC24]. β [LKN96, POB23, WBSD12]. k [HPH90]. k + 1 [HPH90]. n [SSS91].

-acetate [ACG^{+}15]. -cell [LKN96]. -FMISO-PET [GCC^{+}12]. -hydroxyphenylpyruvate [WDD^{+}17].

0 [bMO04]. 0-262-22066-0 [bMO04].

1 [BA91, BKL09, bKP04, SNHMM16, UPM^{+}15, UH96, WHT12, ZZ08]. 1- [ACG^{+}15]. 1-58488-296-4 [bKP04]. 11 [RMLC14]. 18F [GCC^{+}12]. 19 [DMO25, JPB^{+}21, MMT22, NBT^{+}22, PMAV23, PJL^{+}22, RTMT23, TG23].

2 [MDS^{+}24, PGL^{+}22, SPN22].

3D [SZB21]. 3rd [Seg03].

4 [bKP04].

5 [O’N13].

6 [RMLC14]. 6/11 [RMLC14].

978 [O’N13]. 978-0-691-15539-5 [O’N13].

abdominal [JLPB08]. ablation [CMM^{+}17, TKBD^{+}19]. Abundance [CB84]. Acanthaster [Sey90]. acceleration [PP21]. according [BCPS19]. accounts [BMB18]. Accurate [EBFR21]. acetaminophen [RSA14]. acetate [ACG^{+}15]. Achieving [LTW93]. acid [CKL^{+}05, VV24]. acid-injured [CKL^{+}05]. Acknowledgements [Ano84a, Ano85a, Ano86a, Ano96a, Ano97, Ano98a, Ano99a, Ano00a, Ano01a]. across [MDS^{+}24, Ric09]. Action [MGSS21b, MGSS21a, CS08, Ric09]. activation [POB23]. Activator [Jäg86]. active [AS12, Kro10, OKM^{+}05, RBGR^{+}14]. Activity [BA88b, GA94, LKN96, MC01, VL00]. acute [BASH20, LJR^{+}19, Noo12, TRK19]. adaptation [CBR97, SNB^{+}11]. adaptations [CVEH10, GH05]. adapted [DM25]. adaptive [PG10]. Addendum [MGSS21b]. addict [GL00]. addict-needle [GL00]. addiction [RTG21]. adding [BMHK16]. adhesion [And05, Byr97, FFK17]. adjacent [WBG22]. adjoint [Cox06]. adjustment [BRM93]. administration [GLDP17, LMMS10, Noo12, YB18]. adoptive [NO94]. adrenal [BCC08, KN18, LP05, MČM^{+}16]. adrenocorticotropic [BCC08]. adult [SMML92]. aerobic [SGBCC18]. aeruginosa [MBC^{+}14]. affinity [VM13]. affinity-based [VM13]. Africa [BA91, MDS^{+}24, Mwa02, NBT^{+}22]. after [BF03, MEB10]. against [HVPC24, MMA^{+}21, PGC^{+}00, WCM98]. Age [BCC91, Gre87, Gre88, Sch84, vdBD86, DH04, ED00, GA94, HA12, JPB^{+}21, LRW94, MMA^{+}21]. age-dependent [ED00]. Age-Structured [BCC91, Gre87, Gre88, Sch84, DH04, LRW94]. ageing [HvBB10, PP21, Thi16]. Agent [AMMJ86, CLD^{+}18, WBSD12]. agent-based [CLD^{+}18, WBSD12]. aggregates [GP12]. Aggregation [BTCE18, ILA89, VW16, DF18, EHSALR12, FMHF10, GWO^{+}18, SAEHLR98]. Aging [WDS12]. AIDS [AMMJ86, BCC91, GH97, GL00, GDL01, GEW00, Hsi96a, Hsi96b, HC00, Kak99, Kak98, LdA99, MGSS89, YBG96]. airway [BJ14, ZHF^{+}13]. airways [Don23, POB23]. algal [CSeA02]. algebra [AA00]. Algebraic [CH87]. Algebras [HPH90, Rln85]. Algorithm [AC89, WL89, BCS97, ÇÇ17, Har95, RZBR07]. alignment [DSMF00]. Allee [BWSCC11, dSCdA16]. allele [She92, SA02]. allocation [CKS20, GBSRS18, ZB02]. Allostery [Smi20]. alone [Don23]. along [BRR25, GSW87, KK15]. alopecia [CBPD21, DPC18]. Altay [LPSR24]. alterations [CVEH10]. Alternative [HT94, Phe91a]. altitude [FK02]. Altruistic [Kim84]. Alzheimer [BFM^{+}17]. amino [VV24]. amino-acid [VV24]. among [GY93, She92]. amongst [GH97]. Amplitude [AFK87]. amyloid [Kuz24]. anaerobic [SGBCC18]. Analogue [Hop84]. Analysing [DPC18, Pan89, JvdBMH98]. Analysis [BW01, Bro89, CC08, CD86, DH87, Eva00, GS84a, GS84b, KKO^{+}11, Koh13, LMM87, MM87, Poz90, RG95, SSSK91, Swa87, Tho86a, Tho86b, Tho84, Tho87b, WL89, Wei85, Alr19, Ban11, BFK17, ED00, EHE^{+}18, FKK93, HM07, HWK97, HKV07, Kak99, Kak98, KN18, LBPF15, DP08, LOT00, LRW93, MSBM04, MHMI06, MC98, MC01, NPO14, ONE^{+}15, PK21, SDA^{+}23, SW98, TKBD^{+}19, TTBL21, Tho94, VW16]. Analytic [BHB19, BHB20, GEW00, WS09, PMMA^{+}21]. Analytical [Gre87, KE95, RMJLSG10, SSS88, TRMB09, KK22, MMMW15]. Analyzing [SGZ19]. aneurysms [WVH09]. Angiogenesis [CS90, CS91, CS93, DG09, OSMA97, OC96, OC97, PS03]. angiogenic [LMMS10, PS03, SQGP12]. anisotropic [EHS16, MF21]. annual [GD98]. answer [ETO07]. ant [RZBR07]. anterior [CGDF02, TPRS18]. anti [CBR97, DG09, LMMS10, OC97, PS03, SH01]. anti-angiogenesis [DG09, OC97]. anti-angiogenic [LMMS10, PS03]. anti-cancer [DG09]. anti-Hebbian [CBR97]. antibiotic [ARCE^{+}07, Rel05, YB18]. antibiotic-resistant [ARCE^{+}07]. antibodies [PGL^{+}22]. antibody [GY19, PK21, RY22, SNHMM16]. antibody-dependent [GY19, RY22]. antigenic [NBC18]. antimalarial [TCC^{+}11]. antioxidants [MGDPC^{+}15]. antipredator [PB00]. antithrombotics [MGDPC^{+}15]. antiviral [LP08, RB15]. appearance [YF00]. Appendages [CK87]. applicable [KKO04]. Application [BCC91, CD86, dSCM06, LMM87, ALM17, BCS97, BCSS15, BCSS18, BCSS21, GH99, Krz01, LOT00, LG24, MSB10, Ric09, THH92, Tsi24, Ven93, YBG96]. Applications [Seg03, Swa85, Swa88, WW05, BT15, DBG14, RW01, SZB21, SW12, SHGB17, WDS12]. Applied [MC98, GM19, JLPB08, LTW93, Rob99, YC99]. Applying [GCC^{+}12]. Approach [Hib85, SSS88, Tho87a, BMY04, BWSCC11, BZF^{+}15, CCV05, DGN23, EHS16, ETOK03, EP01, MK20, MMT22, PLFP17, Ric09, SSS^{+}20, SPN22]. Approaches [LL00]. Approximate [Tho86a, ILA89]. approximating [HvBB10]. Approximation [AES85, LvdBTJ02, MR20, Par97, SKG11]. approximations [DH04, Fad93, FBJK10, Ren98]. Aqueous [DRRT18]. area [SAEHM03]. areata [CBPD21, DPC18]. argument [BSEO07]. Arjen [bMO04]. arrivals [BR98]. art [RA12, PFF^{+}20]. ART-induced [PFF^{+}20]. arterial [GH05, WGH03]. arteries [MF21]. artery [CVEH10]. arthritis [BDJB^{+}13]. articular [WBG22]. artificial [LK07, TPRS18]. ascending [RL13]. aspartate [BCS96]. Aspects [Cai89, Sey95]. assays [AMP^{+}23]. assess [RY22]. Assessing [Mod88, MDS^{+}24, MHGY18, YJY21]. Assessment [RPSGR^{+}07, HWG11]. assignments [CCS94, CCS99]. assimilation [CBPD21]. assisting [TH96]. associated [HA12, TRK^{+}15, SQGP12]. associations [VdC01]. assumption [ED00]. assumptions [GL00, Whi20]. asthmatic [POB23]. Astrocytic [ACG^{+}15]. Asymmetric [BM02]. asymmetries [Bra93]. asymmetry [SA02]. Asymptotic [GJ90, SK24, KNB11, Kor09a, VW16, WHT12, ZHF^{+}13]. atherogenesis [IMRW05]. atherosclerosis [MF21]. atherosclerotic [Fok12]. atrial [MGSS21b, MGSS21a]. Attracted [WP85]. Auditory [BH86]. aureus [JC19]. Australia [CCL^{+}02]. Author [Ano93a, Ano98b, Ano99b]. Authors [Ano12i, Ano12j, Ano12k, Ano13m, Ano13n, Ano13o, Ano13p, Ano14j, Ano14k, Ano14l, Ano15j, Ano15k, Ano15l]. autocatalytic [LC98]. autoimmune [DPC18, PMB14]. Automatic [AC89]. Automating [PRK05]. autonomous [LL09]. avascular [BM02, WK97, WK99]. avascular-tumour [WK97, WK99]. avian [Yam18]. AVM [SJ03]. Avoidance [DM89]. avoids [HG10]. axial [GH05]. axis [BCC08, KN18, LP05, MČM^{+}16, PMB14]. axisymmetric [WS06]. Axon [Bel86, KK15]. axonal [Kuz13, KK15]. Axons [Bel84, GS84b, KK17].

B [Lut06, bKP04, BASH20, MHMI06, She92]. B-lineage [BASH20]. B.-L [Lut06]. Bacillus [BMHK16]. Bacteria [CF84, TTBL21, WKK^{+}01]. bacterial [BRM93, BD20, CCV05, JCS15, KGG^{+}99, KKC^{+}03, WKK^{+}04]. bacteriophage [Car06]. bacteriophages [KSS16]. bacterium [Zha22]. bacterium-directed [Zha22]. Balancing [GA94]. band [KPSW09]. Barr [HA12]. Barrier [AKR^{+}89, CCV13, LVWG17]. barriers [PMN00]. basal [Noo12]. based [ACRF08, AS88, BCC08, Bla90, CPB10, CMM^{+}17, CLD^{+}18, CGS^{+}22, FBJK10, HG10, Har95, MMA^{+}21, NBGWJ18, Poz90, SAEHM03, SW98, VM13, WBSD12]. basic [LBY00]. Bayesian [BSEO07, GGA06, MSB10, UPM^{+}15, ZOZ^{+}22]. BCG [BMHK16]. be [GBBT19, Has84, KK17, NM92]. beam [NGK19]. Bearings [BBN90]. Behaviour [Bel84, BH86, Bla90, HC00, WP85, BMd19, DMS25, GSBA92, Hsi96a, Hsi96b, IMM94, KKO^{+}11, KNB11, Kro10, LBPF15, LL00, LK07, MNB06, MH97]. behavioural [BMd19]. behaviourally [SGBA92]. being [CCV13]. benefits [SNHMM16, SW98]. Bennett [VdC01]. Bennett-Binet [VdC01]. Bernstein [HPH90]. Best [JKK13, GL00]. beta [Kuz24]. better [GDFH13]. between [AALL23, ABMR21, AS86, BA91, BAJ^{+}17, BL15, BMd19, CB84, dSCD20, DMO25, FKCK17, FK03, HCWK17, JCS15, KK15, NS14, Tho10, VHT07, vdBD86]. Beyond [Tsi24]. biased [BD20, ETO07]. biaxially [ÇÇ17]. bidomain [Whi20]. Bifurcation [LM86, NPO14, TTBL21, Alr19, BWSCC11, GH99, KN18, KG99, ZHF^{+}13]. Bifurcations [MA06, Sle88]. binary [Ven93]. binding [VM13]. Binet [VdC01]. Binomial [OWL^{+}98]. bio [GCC^{+}12]. bio-mathematical [GCC^{+}12]. bioactive [KBVW19]. biochemical [FS17, LC98, TK05]. biocontrol [LvdB98]. biofilm [CDNR16, CK04, MBC^{+}14]. Biological [Ano91, BRR25, DSMF00, Hor84a, Lut06, MO84, dSCdA16, CB15, GL15, Hor96, ONE^{+}15, SKTO93, TRMB09, WW05, Woo94]. Biology [bNFB05, Mai03, Sle88, Sne03, ACRF08, FKCK17, Seg03, Bur06, bKP04, JKK13]. biomass [SAEHM03]. biomechanics [WDS12]. Biomedical [Seg03]. biometry [Ren94]. bioreactor [CW07, OWB10, PSWO14, SMGW13, WCSR06]. bioreactors [SW12]. bipartite [TC04]. Biphasic [MNB06, DM25, SGZ19]. Bird [SW98]. birds [KKO04]. Births [Hor84a]. bistable [BKL09]. bladder [BMHK16]. blast [BASH20]. bleb [DSW^{+}21, SG13]. blended [MS02]. blind [ZNB97]. blink [BF03, HBD^{+}07, WAB10]. blinking [ACH^{+}21]. block [DRRT18, MGSS21b, MGSS21a, SDA^{+}23]. blockage [SBM04]. Blood [SSS91, SSA89, BHO96, DOM^{+}09, LVWG17, LJR^{+}19, LF11, Mat08, Poz05, SSS^{+}20]. blood-brain [LVWG17]. blooms [CSeA02]. bluefin [KT93]. BMI [BvGPB24]. Board [Ano12f, Ano12g, Ano12h, Ano13i, Ano13j, Ano13k, Ano13l, Ano14g, Ano14h, Ano14i, Ano15g, Ano15h, Ano15i]. bodies [CCV05, KK22]. body [IN12]. Book [Bur06, Lut06, O’N13, bMO04, bKP04, Seg03]. bootstrap [BSEO07]. borne [LVTJ00, LVTJ01, Mwa02]. both [HH08]. Bound [Ano86c]. boundary [CS08, GMM^{+}99, HM07, JOBF23, MBHKS10]. bovine [Rob92]. Boyden [CHC13]. brain [CGG^{+}22, EBFR21, HO19, LVWG17, MGDPC^{+}15, OAG^{+}06, PK23, WDS12, WS09]. Branching [Ren85, FHR20, HM07, O’N99, OC96, SJ03]. Brazil [MMA^{+}21, YJY21]. Breaking [XDJ87]. breaks [MSZMGFS18]. breakup [BDB^{+}18, ZKB^{+}19]. breast [DBG14, FBM^{+}03, JBG^{+}19, THH92, WGCB07]. breathing [FK00, FK02, WGH03]. breeding [RKAL02]. Britton [O’N13]. bud [Sch85b]. budget [GBSRS18]. Build [Pug91]. burden [BC00, HYM^{+}24]. burn [RMJLSG10]. burn-depth [RMJLSG10]. burst [BMY04]. bypass [SEL17].

C [ACG^{+}15, CGH12]. Cable [Poz90, KE95, Krz01, PP96]. Calcium [DGS08, ACC18, BHB19, BHB20, CO19, Kee06]. calculated [BHO96]. Calculating [RTG21]. Calculation [MSjH^{+}85]. Calibration [AAVAV24, AC89, ZOZ^{+}22]. calm [BGW99]. Calmette [BMHK16]. Cambridge [bMO04]. Cameroon [NBT^{+}22]. Campbell [Car06]. Campbell-like [Car06]. Can [Sch85a, BV20, Has84, NM92]. Cancer [Swa87, Swa88, AC24, AMP^{+}23, BW01, BTCE18, BC00, BFPH17, BMHK16, DG09, DBG14, FFK17, FHR20, HR16, JBG^{+}19, Koh13, MK20, MD18, Mur95, Mur97, MSB10, OWL^{+}98, RA12, RPSGR^{+}07, SQGP12, UH96, WBSD12]. cancerous [YLS^{+}20]. cancers [NO99]. Canopy [Ren85]. Capillaries [SSSK91, SSS91]. Capillary [OC96, BHO96, CGS^{+}22, OSMA97, PRWJ07]. carbon [AGPV07, LVTJ00, LVTJ01]. Carcinogenesis [GJ90]. carcinoma [FBM^{+}03, RQA12]. carcinomas [Byr97]. cardiac [CS08, CVEH10, DGM^{+}18, MS13, PRK05, TPS^{+}08]. cardiomyocytes [RBGR^{+}14, SDA^{+}23]. cardiopulmonary [SEL17]. Cardiospheres [DGM^{+}18]. cardiovascular [DCC^{+}21, FNdO16, PTW^{+}17]. cards [ML00]. care [Tsi24]. cargo [Kuz13]. caries [FR14]. Carlo [GT91, LTW93, YB03]. carotid [CVEH10]. carriage [JC19]. carriers [ETO07]. cartilage [TPL09, WG20, WBG22]. cascade [WBSD12]. case [CBB94, GL00, KSS16, LCC^{+}17, MS05, PMAV23]. Cases [Bro89, DMNW20, Yam18]. catastrophe [GF08]. Catastrophic [Ree88]. Catch [CB84, Lud89]. Cats [GNA90]. Causative [AMMJ86]. causing [FK00]. CD8 [AALL23]. Cell [BH86, CBW89, CBWA91, GMS89, ML10, Sne03, WKK^{+}04, And05, ACRF08, BWWvB04, BWW05, BTCE18, Bra93, CPB10, CHC13, CCVV98, EVK^{+}18, FKCK17, FFK17, GP12, GWO^{+}18, HMM23, KWW^{+}13, Krz01, LL09, LJR^{+}19, LK07, LKN96, LP21, MSBM04, MHMI06, NBGWJ18, OK13, OKM^{+}05, OSMA97, Ric09, RBMH17, SZB21, SMGW13, SW12, VHT07, ZZ08]. Cell-Chemotaxis [GMS89]. Cell-cycle [ML10, MSBM04]. cell-extracellular [OSMA97]. cell-growth [BWWvB04]. cell-packed [SW12]. Cell-signalling [WKK^{+}04]. Cells [THI91, AALL23, BFPH17, BHO96, CS93, CMR^{+}04, DGM^{+}18, DSW^{+}21, EHE^{+}18, Koh13, MS01, MS05, Poz05, RNE^{+}20, RY22, TKBD^{+}19, TC06]. cellular [Egl06, FMHF10, FFK17, NO94]. centers [MHMI06]. central [FK00, RBMH17]. centred [ML00]. Centres [BA91, CKS20]. cerclage [GBPF16]. cerebellar [BMB^{+}23]. cerebral [WVH09, WS08]. cerebrospinal [ESC13]. CFD [BAJ^{+}17]. Chain [Ano86c, GR98, LTW93, YB03]. chains [dSCdA16]. chamber [CGDF02, CHC13, DRRT18, TPRS18]. change [Hsi96a, Hsi96b, HC00, LPSR24, THH92]. Changes [Poz88, DMO25, GH05]. Channel [BS87a, Cox06, SDA^{+}23]. channels [BMY00, BMY04, MH97]. Chaos [Bok93, Ren94, Sle88, CKC07]. chaotic [OOB01]. characteristics [JvdBMH98]. characterization [YLS^{+}20]. Characters [Bur86]. charge [LBY00]. chemical [CBB92, PSWO14]. chemokinesis [BCM98]. chemoreceptor [FK00]. chemorepellent [EHH12]. chemotactic [CS93, FM08, GWO^{+}18]. Chemotaxis [AES85, GMS89, BD20, BCM98, CHC13, FMHF10, PMN00]. chemotherapeutic [BG11]. Chemotherapy [Swa85, Swa87, Swa88, BC00, CBB94, MS01, MS05, Mur95, Mur97, NO99, UH96]. childhood [LJR^{+}19, NS95, OWL^{+}98, PGC^{+}00]. China [LPSR24]. Chiral [XDJ87]. CHO [EHE^{+}18]. choanoflagellates [OOB01]. cholera [Yam18]. cholesterol [MČM^{+}16]. chondrocyte [CHC13, TPL09]. chromosomes [BCS97]. Chronic [HvBB10, BZF^{+}15, BL15, RSA14]. CI857 [CCVV98]. CI857-controlled [CCVV98]. Circuits [Hop84]. Circular [MSjH^{+}85, PSS85]. Circulation [SSS88]. circulations [DCRK21]. citrate [ACC18]. city [NS14]. class [KKO04, Tho93]. Classification [PK21, RZBR07]. Clearance [BBR84, ST89]. climate [LPSR24]. clinical [BCC08]. cloning [RY22]. Close [Pug91]. closure [GMM^{+}99, GBPF16]. cloth [BT21]. Clustering [BS87a]. clusters [DGM^{+}18]. CMJ [Mod97]. co [BKL09, SSA89]. co-operativity [BKL09]. coagulation [DOM^{+}09, LF11]. coated [EHSALR12, SAEHLR98]. code [TC04]. coding [TS94]. coefficients [MBS92]. coevolution [Sey95]. coevolutionary [AALL23]. coexpression [BK19]. cohort [EVK^{+}18]. cokriging [OWL^{+}98]. coli [CCVV98, SGBCC18]. collagen [ÇÇ17, DSMF00, LLZ^{+}17]. collagen/tissue [LLZ^{+}17]. Collapse [Ree88]. Collective [GP11, BMS04]. colony [KKO^{+}11, RZBR07]. colour [RMK04]. colubriformis [DGDW87]. column [BGW99]. Combination [BZF^{+}15, Zha22]. combinatorial [LLZ^{+}17]. Combined [MGDPC^{+}15, CHC13, WBSD12]. comedo [FBM^{+}03]. commonly [RTG21]. communities [BWSCC11, MANS11]. Community [AAVAV24]. comparative [Bok93, Mon07]. Comparison [GS84a, AAVAV24, BW01, JCS15, KK15]. comparisons [EHS16]. compartment [GGA06, HCWK17, NDd17, RTG21]. Compartmental [Fad85, BSA^{+}12, Fad93, GR98, GEW00]. compartmentalized [YBG96]. competing [LC04]. Competition [MJB11, BP98, SK24]. competitor [BP98]. Complex [GT91, TG91, She92, Tho94]. complexity [Bok93, PK23]. compliance [WS08]. complicated [GDFH13]. complication [SHGB17]. Component [GT91]. Components [BH90]. composition [OS98]. compression [CGP06, GP12]. Computation [Tho86a, Tho86b]. Computational [RQA12, Sne03, SQGP12, ÇÇ17, RA12, SL16, SG13, YLS^{+}20]. computations [WG03]. Computed [LBD^{+}16, LBHKS18]. Computer [HPR^{+}23, MBEP84, Ren85, AA00]. Computing [YBG96]. concentration [KK17, PSWO14]. concept [PGPR16]. conciliating [PG10]. conditional [Tho10]. conditions [BMN^{+}93, MBHKS10, MH97]. condom [GDL01]. Conduction [Bel86, SL16]. Conference [Ano91]. configuration [CGS^{+}22]. connectivity [PK23]. Consequences [BS90]. considerations [Sch85b]. considered [GBBT19]. considering [LBY00, YC99]. Consistent [Pug91]. constants [TK05]. constituent [DGN23]. constitutive [Kro10, WS08]. Constrained [Bur86, GH05]. constraints [BG11, MS01, MS05, MSBM04, Sey95, She92, TC06, UA96]. construct [OWB10]. Contact [BA91, Pug91, RB15, ACH^{+}21, CLD^{+}18, PJL^{+}22, RTMT23, Sey95, TXAS15, dAL02]. contagious [DMO25]. containing [ACC18, EP01]. containment [BB21]. Contents [Ano12a, Ano12b, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano15a, Ano15b, Ano15c]. context [dSCD20, NBT^{+}22]. continued [Par97]. continuous [BP98, DGM^{+}18, TRK^{+}15]. continuous-external-fertilization [BP98]. Continuum [FBJK10, AALL23, BKW^{+}14, FG04, HYM^{+}24, KBVW19, SMGW13, SSS^{+}20, TPL09]. contractile [BJ14]. contraction [RBGR^{+}14]. contribution [GBBT19]. Control [AS88, HW86a, KR90, MCRC08, MC01, Swa85, Swa87, Swa88, AGPV07, ACS14, AS12, BMN^{+}93, BMB18, CLM^{+}20, CKC07, CBB92, dSCdA16, – DD22, EP01, FKK93, GDFH13, GC08, GBSRS18, Hor96, JLPB08, Lu11, ML10, MG14, NS95, PMB14, Rob92, RW01, WW05, ZB02]. Controlled [AC89, CKL^{+}05, CCVV98, LP05]. convective [SAEHLR98]. cooperation [AC24]. cooperative [KT93]. Coral [AFK87]. core [HC00]. cores [Fok12]. Corneal [KP91, GMM^{+}99]. coronavirus [YJY21]. Correlations [HC84]. Corrigendum [BCSS18, BCSS21]. cortical [TC06]. corticotropin [BCC08]. corticotropin-releasing [BCC08]. cortisol [BCC08]. costs [SW98]. could [KK17]. Counters [Ano86c]. counting [DJG06, VL00]. Counts [Tho87b]. County [PMAV23]. Couple [BBN90, TC06]. Coupled [GS84b, FNdO16, FM08, SL16]. coupling [CO19, MGSS21b, MGSS21a, SKG11]. CoV [MDS^{+}24, PGL^{+}22, SPN22]. Cover [Ano12c, Ano12d, Ano12e, Ano13e, Ano13f, Ano13g, Ano13h, Ano14d, Ano14e, Ano14f, Ano15d, Ano15e, Ano15f]. coverage [AA00, PGC^{+}00]. COVID [DMO25, JPB^{+}21, MMT22, NBT^{+}22, PMAV23, PJL^{+}22, RTMT23, TG23]. COVID-19 [DMO25, JPB^{+}21, MMT22, NBT^{+}22, PMAV23, PJL^{+}22, RTMT23, TG23]. CPR [JLPB08]. CRC [bKP04]. Critical [Mod88]. Cross [WCM98, BK19, HR16, Ven93, ZEM11]. cross-diffusion [ZEM11]. cross-inhibitory [BK19]. cross-section [Ven93]. cross-sectional [HR16]. Crossover [Tho87b]. Crown [AKR^{+}89]. Crown-of-Thorns [AKR^{+}89]. crucial [PGC^{+}00]. Cuban [LdA99]. culling [dSCM06, dSCD20]. culture [HMM23]. cumulant [Ren98]. cure [EVK^{+}18, NPO14]. current [PP96]. currents [Das08]. curvature [CO19, Poz05]. Curve [AC89]. Curves [SSA89, KGG^{+}99, PGL^{+}22]. cyanobacterial [BGW99]. Cycle [UH96, BK19, CVEH10, CW14, HBD^{+}07, LL09, ML10, MSBM04, MS13]. Cycle-nonspecific [UH96]. Cycles [Sey90, LL09, LKN96]. cycling [DPC18, Rel05]. Cylinder [Poz88, Eva00]. cylindrical [MF21]. cystic [MBC^{+}14]. cytokine [BDJB^{+}13, SH01, ZJJA19]. cytokine-mediated [BDJB^{+}13]. cytokine-receptor [SH01]. cytosol [IN12]. cytotoxic [AALL23, MCK04].

D [BCS96, bKP04, Seg03, BKW^{+}14, BKF14, FDC^{+}24, GL15, GH05, TC06, WS06]. daily [NS14]. Damage [Tho87a, BMB18, JCS15]. Damping [KR90]. Darren [Bur06]. Data [Bro89, CBPD21, Lud89, Pan89, Pug91, Wei85, BCS97, BCSS15, BASH20, BT15, CGS^{+}22, FKCK17, HKV07, KGG^{+}99, LTW93, LLZ^{+}17, MMA^{+}21, MSB10, PK21, Ree97, RZBR07, RMK04, YB18, BCSS18, BCSS21]. data-driven [BASH20, LLZ^{+}17]. DCIS [DBG14]. Death [Gre90, Gre92]. Debarquement [BMB^{+}23]. decay [PGL^{+}22]. decision [MK20]. decline [Thi16]. deconvolution [YF00]. Defensive [AS86]. deformable [SGZ19]. deformation [KPSW09, VHT07, WG20]. degenerating [KK17]. degeneration [CLM^{+}20]. degradation [Alr19, KKC^{+}03]. degree [RW01]. Delay [PGBMG^{+}15, Tam99, Alr19, Car06, CSeA02, KG99, LP05, MGDPC^{+}15]. delay-differential [LP05]. delayed [RW01]. delays [GL15, WHT12, ZZ08]. delivery [AP22, FNdO16, LKH98, PLFP17, SGZ19, VM13]. Demographic [LHZ87, MPE87]. demographics [RJ96]. demography [RTMT23]. dendrite [Kuz13]. dendrites [KB07]. Dendritic [Poz88]. dengue [GY19, MMA^{+}21, NS14, NBC18, RY22]. Densities [Hib85]. Density [Gre90, Has84, MLVHCC99, Gre92, LP21, RJ96]. Density-Dependent [Gre90, MLVHCC99, LP21, RJ96]. dental [FR14]. departures [BR98]. Dependence [Bel86, Has84, BP98, Gre92, HA12]. dependencies [FKCK17]. Dependent [Gre90, Ban11, ED00, EHE^{+}18, GY19, Kor09a, LP21, MLVHCC99, MA06, NSCO97, RJ96, RY22, Sec16, VHBCC96]. depletion [LK07]. deposition [BFK17, DSMF00, JPM^{+}05, LF11, ZHF^{+}13]. depth [RMJLSG10]. derived [BASH20, MMA^{+}21, Tho10]. dermal [DSMF00]. dermatological [BT21]. Descemet [IFP13]. Descent [Tho88b]. describe [GY19, RY22]. Describing [BASH20, DM89, DGDW87, BRM93, OAG^{+}06]. description [KKO04]. descriptions [ADCN24]. Design [AM84, SW12]. detached [IFP13]. detaching [BBPF13, LBPF15]. detachment [GBPF16]. detectable [SA02]. detecting [TS94]. Detection [DWLS87, LKN96, TKBD^{+}19]. determination [RMJLSG10, TK05]. determine [HVPC24]. Deterministic [Bla90, RMLC14, Zha22, ABMR21, BRM93, PG10]. develop [GCC^{+}12]. developing [CBB92]. Development [Egl06, NBGM85, SSA89, BLLM18, BGW99, Byr97, CGP06, CK04, JvdBMH98, TPL09, bMO04]. devices [LKH98, MHGY18, TSE24]. diabetes [NPO14, NSCO97]. Diagnosing [WDLS88]. diagnosis [DCC^{+}21, NDd17]. Diagnostic [THH92, PK21]. diagnostics [SBM04]. dialysate [ACC18]. diameter [SNB^{+}11]. diameters [AS12, HR16]. Dictyostelium [DJG06]. Diekmann [O’N13]. Dielectric [THI91]. differences [HCWK17]. Different [DWLS87, RB87, BD20]. Differential [AS86, BSA^{+}12, CSeA02, LP05, MC98, PMMA^{+}21, PDD08, bKP04]. Diffusion [AES85, AKR^{+}89, BT21, CS91, Fad93, GS84a, GS84b, KJ88, Kim84, BRR25, BMM93, CHV93, EHSALR12, GM19, HSV24, KK15, LKR96, LG24, MBS92, NM92, OAG^{+}06, PK23, SAEHLR98, ZEM11]. Diffusive [Bel86, MJB11]. diffusivity [Kuz24]. digital [PJL^{+}22]. Dilemma [Phe91a]. dilution [FHC09]. dimension [RZBR07]. Dimensional [BW87, ST85, CO19, HZ25, LOT00, OC97]. dimorphic [Woo94]. dioxide [AGPV07]. dioxygenase [WDD^{+}17]. directed [Zha22]. Directions [AS84]. directly [LBY00, Yan02]. discoideum [DJG06]. discontinuous [TRK19]. Discrete [AALL23, LHZ87, Sle88, DGM^{+}18, DSMF00, ETOK03, SSS^{+}20, SNHMM16, TRK^{+}15]. discrete-continuum [SSS^{+}20]. discrete-time [ETOK03]. Discrimination [DH87]. Disease [MA07, O’N13, BFM^{+}17, BL15, CBPD21, DMO25, ED00, EVK^{+}18, HWK97, HvBB10, HH08, HMMN97, KM05, LL00, LRW93, LRW94, MG14, Mwa02, Por99, RZBR07, Rob99, Sey95, VHBCC96, Ven02, XZT11, Yam18, dAL02]. diseases [Mod97, Mwa02, PGC^{+}00, ZJJA19, dMS08]. disintegration [CCV05]. disorder [DPC18]. disordered [NBGWJ18]. Disorders [DWLS87, WDLS88]. Dispersal [RR99, DMS25, PvdB98, XZT11]. dispersing [DH04]. Dispersion [Bla90, ALDLDMC12, ÇÇ17, FHC09, PRWJ07]. Disputed [Bro89]. dissemination [RTMT23]. Dissociation [SSA89]. distancing [RTMT23].

Distinct [NBGM85]. Distributed [BA88a, Alr19, Tho10, WHT12]. Distribution [DWLS87, GJ90, Pan89, THI91, ETO07, KK15, Par97]. Distributions [GSW87, DBG14, GC08, HvBB10, KKO04, SA02, Ven93]. divergence [AC24]. division [O’N99]. division-within-division [O’N99]. DNA [Bro89, MSZMGFS18, TS94]. does [BKL09, GP11, ZOZ^{+}22]. domain [LBD^{+}16, MBHKS10]. Dosage [HYM^{+}24, Noo12]. dose [BFK17, GC08, HVPC24, Kor09a, PR16, PGC^{+}00]. dose-dependent [Kor09a]. dose-rate [PR16]. doses [BCPS19]. double [MSZMGFS18]. doubling [Sle88]. Dougherty [PMAV23]. downs [JC19]. downstream [ABMR21]. drainage [BF03]. draining [JPM^{+}05]. Drawbacks [CBW89, CBWA91]. Drawing [Tho88a]. Drift [KR90]. drive [SW12]. driven [BASH20, LLZ^{+}17, MC98, ZKB^{+}19]. drop [TPS^{+}08]. droplets [MHGY18]. drug [BC00, CGH12, CBB92, CBB94, FNdO16, GY93, GLDP17, GH97, HYM^{+}24, LKH98, LCC^{+}17, MMWM11, MMMW15, Mur95, PLFP17, TCC^{+}11, TW00, UH96]. drug-delivery [LKH98]. drug-eluting [MMWM11, MMMW15]. drug-resistance [HYM^{+}24]. drug-resistant [UH96]. drugs [LMMS10, LKR96, MS05, Mur97, UH96]. drying [BT21]. dual [SJ03]. ductal [FBM^{+}03]. due [DH87, DRRT18, MMF24, OOB01, PP21]. dummy [KGG^{+}99]. duration [CS08]. during [AGPV07, ACC18, ASY16, CKS20, LJR^{+}19, PDR^{+}21, SEL17, Thi16, THH92, WS09]. Dynalet [DM25]. Dynamic [CD86, TH96, ZB02, CVEH10, GBSRS18, MPP05, Tsi24]. Dynamical [MHMDM17, DGN23, POB23, TW00]. Dynamics [AMMJ86, Ano91, BMS04, BWSCC11, BA88a, Bur86, GT99, JPM^{+}05, RG91, Sch85a, Sey90, AAVAV24, ARCE^{+}07, ADCN24, AALL23, ACH^{+}21, ADH^{+}05, ACG^{+}15, BCC08, BLLM18, BTCE18, BDB^{+}18, CO19, CCL^{+}02, CDNR16, DCC^{+}21, DGS08, DLM23, DPC18, ETOK03, FY19, FFK17, GP11, HWG11, JCS15, Kor09a, KSS16, LP01, LJR^{+}19, DP08, LTY15, LBD^{+}16, MBHKS10, MČM^{+}16, MLVHCC99, MS13, PMAV23, PG10, PDD08, PRK05, RNE^{+}20, Rob92, RG92, RL13, SKTO93, SBM04, SH01, SWWC09, UPM^{+}15, VDLK13, VHBCC96, WHT12, Woo94, ZJJA19, ZZ08, O’N13].

E. [CCVV98]. ear [MEB10]. Early [YB18, NDd17, RB15]. eating [vdBD86]. ECG [ZNB97]. eco [CKC07]. eco-epidemiological [CKC07]. Ecoepidemics [CCV13]. ecological [RMK04, Sey95]. Ecology [Sch85a, Por99]. Ecosystem [You90, Kor09b]. Ecosystems [ILA89]. Ed [Seg03]. edema [TRK19]. Edited [bMO04, RNE^{+}20]. Editorial [BC06, Dem25, JKK03, Ano12f, Ano12g, Ano12h, Ano13i, Ano13j, Ano13k, Ano13l, Ano14g, Ano14h, Ano14i, Ano15g, Ano15h, Ano15i]. education [HWG11]. EEG [Das08]. EEG-MEG [Das08]. Effect [BBN90, Kuz24, RL13, SW06, vdBD86, BHO96, BCM98, CBB94, HC00, LC04, MSZMGFS18, Par97, Phe95, RG92, RY22, Tam99, VdC01, ZLC03]. Effective [EHS16]. effectiveness [LP08, MHGY18, RTMT23]. Effects [BMN^{+}93, OS98, PMAV23, VHBCC96, BLLM18, BWSCC11, BSA^{+}12, dSCdA16, DJG06, FFK17, LK07, MRG98, MČM^{+}16, MDS^{+}24, Mur95, Noo12, OSS05, PGBMG^{+}15, SGBCC18, SGZ19, TW00, WPH^{+}18, Yan02]. efficiency [OOB01, SGRPSM^{+}09]. Efficient [TRK^{+}15, WG03, Har95]. Effort [Lud89]. EGFR [WBSD12]. EGFR-TGF [WBSD12]. Egg [vdBD86]. Egg-eating [vdBD86]. Eggs [LMM87, BHB19, BHB20]. egocentric [BT15]. elastic [BBN90, CB15, KPSW09]. elder [YJY21]. electric [GMM^{+}99]. electrical [KB07, LKN96, MC98]. electrochemical [Ric09]. electrodiffusion [EBFR21]. electrophysiology [Whi20]. Element [SSSK91]. Elements [Oht84]. elimination [HBHJ22, PR16]. Ellipsoids [DS91]. elliptical [RK01]. elongation [MG06]. Elucidate [Sch85a]. eluting [MMWM11, MMMW15]. embryonic [SMML92]. emergence [HBS^{+}12, MA07]. Emergent [CO19]. empirical [HF20]. encephalitis [GT99]. encompassing [FY19]. end [MSZMGFS18]. Endangered [Mod88, BWM97]. endemic [DMO25, SGBA92]. endothelial [CS93, OSMA97]. Energetics [RG95, PMMA^{+}21]. Energy [BP98, SW98]. energy-based [SW98]. engineered [TPL09]. engineering [SW12]. Enhanced [OOB01, PR16]. enhancement [GY19, RY22]. enrichment [KSS16]. entries [Tsi24]. entropy [CBR97]. entry [DPO^{+}15]. Enumeration [TC03]. Enviroments [Tan87]. Environment [Mod88, SSS88, BRM93, CLD^{+}18, CDNR16, DH04, MJB11]. environment-host [CLD^{+}18]. Environments [Has84]. Enzyme [BL86, Mat08]. Epidemic [AS84, BCC91, GP84, Gan84, Gre87, Gre88, Gre90, GSW87, Lou89, MK89, MGSS89, Pug91, Rin90, ADH^{+}05, BZ93, BB21, DLM23, DMO25, DMNW20, DMS25, ETOK03, ETO07, Gre92, GEW00, JvdBMH98, KG99, Kor04, MA07, MLVHCC99, MA06, PFF^{+}20, SKG11, VW16, ZB02]. Epidemics [Cai89, Cai90, Ven02, Bak92, BRR25, BT15, HWK97, JvdBD02, Mon07, Tsi24]. epidemiological [CKC07, MS02, MMT22, YJY21]. Epidemiology [Sch85a, GT99, RJ96]. epidermis [SMML92]. episodic [SKTO93]. epithelial [FBJK10, GMM^{+}99]. epithelium [NBGWJ18]. epizootic [LvdB98]. epizootics [Bak92]. Epstein [HA12]. Epstein-Barr [HA12]. Equation [Kim84, ST89, vdBJ84, BWW05, CSeA02, HBD^{+}07, KNB11, Krz01, LP05, RMJLSG10, TRMB09]. Equations [MM87, bKP04, BSA^{+}12, BMM93, EHS16, MC98, OAG^{+}06, PMN00, Whi20, WS09]. equatorial [GBPF16]. equilibria [RKAL02, TW00]. Equilibrium [HW86a, YB03]. Equivalent [Poz88, Poz90, Eva00]. Equivalents [MSjH^{+}85]. eradication [PGC^{+}00]. Erratum [Ano89, BHB20]. Erythrocyte [RS90]. escalation [BG11, PR16]. Escherichia [SGBCC18]. Eskimos [She92]. Essential [Mai03]. estimate [TSE24, Woo94]. estimated [ACG^{+}15]. Estimating [AA00, CMR^{+}04, DCT06, GR98, Hib85, PGL^{+}22, Ree97, CB15, OWL^{+}98]. Estimation [GJ90, GT91, Lud89, NSCO97, Ven93, WBGN89, AGPV07, Don23, MMMW15, PDD08, TRK^{+}15, TKR^{+}18, UPM^{+}15]. estimations [ESC13, SAEHM03]. ethanol [Noo12]. Evaluating [Kap01, LP08, RTG21]. Evaluation [TG91, ADCN24, BCS97, Whi20]. evaporation [TXAS15, WAB10]. evidence [BCC08]. evoked [VL00]. Evolution [DGDW87, IR85, Kim84, PM86, BASH20, CHV93, ÇÇ17, DBG14, Phe95, WS06, YBG96]. Evolutionary [Bur86, DMS25]. evolutionary-epidemic [DMS25]. evolving [HBHJ22]. Exact [YB03]. Exactly [Lut06]. example [DMO25, Mur95]. Examples [HW86a, MBEP84, CCS94]. Exchange [SSS88, ACC18, PDR^{+}21]. exemption [dMS08]. exercise [CC08]. exhibiting [RNE^{+}20]. Existence [MR20, MM87]. existing [NPO14]. Exoprotein [CF84]. expectations [YBG96]. experimental [BCM98]. Experiments [GJ90, JCS15]. explained [BMB^{+}23]. explanation [IFP13]. exploitation [KT93]. Exploited [Hor84b]. exploiting [BM02]. Exploring [DGN23, SNHMM16]. exponential [RW01]. exposure [– DD22, GBBT19]. expression [CCVV98, HKV07, RZBR07]. Extended [TKR^{+}18, THH92]. External [Hor84b, BP98, NGK19]. Extinction [ZLC03]. extracardiac [DCRK21]. extracellular [CPB10, KKC^{+}03, OSMA97]. extracorporeal [PDR^{+}21]. extrema [CC08]. Extreme [PGPR16]. extrinsic [RQA12]. eye [BAJ^{+}17, CGDF02, DRRT18, LBD^{+}16, MBHKS10, TPRS18]. eye-shaped [LBD^{+}16, MBHKS10]. eyeball [KPSW09].

Factor [CS90, CS91, CS93, DJG06, FKCK17, LLZ^{+}17]. factors [BK19, PP21, VM13]. failing [DCRK21]. failure [LVWG17]. Familial [HC84]. families [IC04]. family [DG09]. fascicles [PK23]. Fast [Tho87a, LKN96]. Fatal [dMS08, LRW93, LRW94, VHBCC96]. favoritism [PYLC09]. feature [RZBR07, TC06]. features [KB07, RPSGR^{+}07]. feedback [GL15, LP05, Mag92, PMB14, RL13, SL16, VDLK13]. feedback-controlled [LP05]. feedback-mediated [RL13, SL16]. feeding [OOB01]. feline [GGA06]. female [Thi16]. fermentation [SGBCC18]. fertilisation [BHB19, BHB20]. fertilization [BP98, LMM87]. Fibre [DWLS87, ÇÇ17, PSWO14, SW12, Ven93]. Fibres [GS84a]. Fibrin [GFK07]. fibrinolysis [BKW^{+}14, BKF14]. fibroblast [MGSS21b, MGSS21a]. fibrosis [MBC^{+}14]. Fickian [FNdO16]. field [Bra93, Ven93]. Fields [MSjH^{+}85, Das08, GMM^{+}99, Tho10]. Film [BBN90, ACH^{+}21, BF03, BDB^{+}18, HBD^{+}07, JPM^{+}05, LBD^{+}16, MBHKS10, MBB20, OKM^{+}05, TXAS15, WAB10, ZKB^{+}19]. Filtering [CK87]. filtration [NSCO97]. finding [Har95]. fingerprint [MLH92]. Fingerprinting [Bro89]. Finite [SGBCC18, SSSK91, UA96]. Finite-Element [SSSK91]. Finite-time [SGBCC18]. fire [Ree97]. firing [MC98]. first [OC96, PGC^{+}00]. Fish [Hor84b, GL15, KB07]. Fisher [EHMS22]. Fisheries [CB84, HW86a, dSCM06]. Fishery [Ree88, dSCD20]. Fishing [Hor90]. Fitness [BS90]. fitting [KGG^{+}99]. FitzHugh [MC98, ZEM11]. fixated [TPRS18]. flea [AS88]. Fleets [Hor90]. Flow [CK87, SWWC09, TPRS18, CGDF02, CVEH10, CW07, DRRT18, EHSALR12, GH05, GFK07, LF11, LBHKS18, OKM^{+}05, OOB01, PSWO14, RK01, SSS^{+}20, SJ03, TMM96]. flow-diffusion [EHSALR12]. flow-rate [TMM96]. fluctuations [AJ12]. Fluid [BLLM18, CGDF02, CK87, SW12, SN98, CDNR16, CW07, ESC13, IFP13, PSWO14]. fluorescence [LBHKS18]. FMISO [GCC^{+}12]. foetal [ZNB97]. Fokker [BFK17]. follicle [DPC18, PP21, Thi16]. Follow [BGW88]. Follow-up [BGW88]. following [BGW99]. Fontan [DCRK21]. food [CK93, dSCdA16]. foods [YF00]. foot [HWK97]. foot-and-mouth-disease [HWK97]. foragers [BR98]. foraging [RG95]. force [BJ14]. forces [MMA^{+}21]. Forecasting [DMO25, NBT^{+}22]. forest [Ree97]. forest-fire [Ree97]. Form [MO84, EHSALR12]. formalism [GMM^{+}99]. formation [BMS04, Ban11, BAJ^{+}17, DGM^{+}18, CMM92, GWO^{+}18, GFK07, JOBF23, Kuz24, MBS92, MBB20, Mat08, NW93, OC96, PRWJ07, RQA12, SG13, YRCU14]. Formulae [SSA89]. formulated [Mod97]. formulation [FG04, MSBM04]. forward [Don23]. four [RTG21]. fraction [Par97]. fractional [BCSS15, BCSS18, BCSS21]. fractional-order [BCSS15, BCSS18]. fractionated [BCPS19]. fractionation [SGK16]. frames [ABMR21]. framework [CLD^{+}18, SZB21]. free [HM07, RG92]. free-boundary [HM07]. free-living [RG92]. frequencies [Ree97]. Frequency [Hor84b, SA02]. fronted [EHMS22]. fronts [HZ25]. fully [WS06]. Function [THI91, CBR97, Don23, Sec16, YF00]. Functional [HO19, vdBD86, BWW05, OAG^{+}06]. Functionally [Bur86]. Functions [PM86, Tho86a, Tho86b, Kor04, Ren98]. fundus [ALDLDMC12]. fungal [LvdBTJ02]. fungi [CTB20, LVTJ00, LVTJ01]. fungi-initiated [CTB20]. Further [CCS99]. fusion [VHT07].

Gabon [NBT^{+}22]. Gain [Tho84]. gains [YB18]. Galerkin [TRK19]. Gambian [AG01]. game [DLM23]. Games [AS86, BP98]. Gametes [AES85]. ganglia [Noo12]. Gas [SSS88, WG03]. gastric [DPO^{+}15]. GATA [BKL09]. Geiger [Ano86c]. gel [GFK07]. Gene [Tho88b, CCVV98, RNE^{+}20, RZBR07]. gene-edited [RNE^{+}20]. genealogical [Lew92]. genealogy [CCS94, CCS99]. General [BCC91, GP84, Swa88, BZ93, CHMH98, ED00, GC08, Kor09b]. generalist [ACS14, MCRC08]. generalization [BZ93]. Generalized [ST89, EHSALR12, Eva00, PK23, SAEHLR98]. Generated [Jäg86]. generating [Ren98]. Generation [HPH90, MO84, BJ14, BGG03]. genes [LTY15]. Genesis [MMF24]. Genetic [AM84, DGDW87, Phe91b, TG91, VdC01, Wei85, BCS97, BKL09, FY19, Phe95, SGBCC18]. Genetics [CH87, HC84, Oht84, Tho93]. genome [Phe95]. Genomes [PSS85]. Genomic [TSL94]. genotypes [HT94, She92]. genotypic [CCS94, CCS99]. genotyping [IC04]. geographically [ETO07]. Geometric [GP84, MSjH^{+}85, KB07, TC06]. geometries [Sec16]. geometry [SGZ19]. Georgia [PMAV23]. Geriatric [TMM96, IMM94]. Geriatric-patient [TMM96]. germ [DSW^{+}21]. germinal [MHMI06]. GIST [LCC^{+}17]. Glagov [MF21]. glioblastoma [LG24, SZB21]. glioma [EHS16, GCC^{+}12, MPP05]. glioma-host [MPP05]. gliomas [PGBMG^{+}15, PGPR16]. glob [ZKB^{+}19]. glob-driven [ZKB^{+}19]. Global [Kor09a, Rin90, VDLK13, WHT12, Yam18, BT15, Kor04, Kor09b, DP08]. glomerular [NSCO97]. glucose [GC15, PDD08, YF00]. Go [HBS^{+}12]. Gompertz [BCSS18, BCSS21, BCSS15]. Gonococcal [GSBA92, SGBA92]. grade [PGBMG^{+}15, PGPR16]. Gradient [Tan87, KK17]. Gradient-like [Tan87]. granulocyte [BASH20]. granulocyte-monocyte-blast [BASH20]. graph [Lew92, Ven93]. graphical [HT94, TC04]. graphs [DCT06, Tho93, TC03]. Great [Hib85, AKR^{+}89]. Green [Sec16]. grid [MBDKS08]. Grodins [FKK93]. Group [Kim84, BMS04, BR98, RG95]. groups [HC00]. Grow [HBS^{+}12, LF11]. growing [SGBA92]. Growth [BH90, CBW89, CBWA91, Fok12, NBGM85, Ren85, Tan87, BV20, BRM93, BWWvB04, BWW05, BTCE18, BCSS15, BCSS18, BCSS21, BFPH17, BB21, BM02, BP03, CGG^{+}22, CDNR16, DGM^{+}18, CCVV98, CW07, DBG14, GSBA92, GF08, GCC^{+}12, JOBF23, KGG^{+}99, Koh13, KK22, LVTJ00, LVTJ01, LvdBTJ02, LLZ^{+}17, MR20, MBC^{+}14, MF21, Muñ16, OWB10, ONE^{+}15, OS98, PSWO14, SGBA92, VM13, WK97, WK99, WCSR06, WVH09, WA20]. Guérin [BMHK16]. gynaecological [NO99].

H [O’N13]. habitat [HZ25]. haematological [NO99]. Haematopoiesis [GNA90, BM01]. haematopoietic [RNE^{+}20]. haemodiafiltration [GM19]. haemodialysis [ACC18]. Haemofilters [BW87]. Haemoglobin [SSS91]. hair [DPC18]. Haldane [Phe91a]. Haptenated [Ano86c]. Hardy [YB03]. Hardy-Weinberg [YB03]. harmful [CSeA02]. Harvest [HW86b]. Harvested [Hor90]. Harvesting [BS90, Ree88, RB87, CCV13, Hor96, Mag92]. Hashimoto [PMB14]. Hausdorff [ALDLDMC12]. HBV [BZF^{+}15]. head [WWWW^{+}14]. head-up [WWWW^{+}14]. heal [ZOZ^{+}22]. healing [OSMA97, SMML92, ZOZ^{+}22]. health [BMd19, HWG11, RT01, Tsi24]. healthcare [JPB^{+}21, TRK^{+}15]. healthcare-associated [TRK^{+}15]. healthy [DCRK21, VHT07]. heart [BLLM18]. heartbeat [CC08]. heat [AP22, TRMB09]. heating [TRMB09]. Hebbian [CBR97]. Heesterbeek [O’N13]. hematocrit [BLLM18]. hematopoiesis [GGA06]. hemodynamical [MMF24]. hemorrhagic [DCC^{+}21]. heparin [VM13]. heparin-binding [VM13]. hepatic [CMM^{+}17, HR16, HCWK17]. hepatitis [CGH12, DGN23]. HER2 [JBG^{+}19]. Herbivores [AS86]. Heterogeneity [AM84, MPE87, GSBA92, MANS11]. Heterogeneous [BA88b, Cai89, Cai90, DH04, HvBB10, VW16]. heterologous [GY19, RY22]. heterosexual [GA94]. HFV [BA88c]. hidden [GGA06, KLG^{+}14]. hierarchical [HVPC24]. high [BKL09, FK02, PGC^{+}00, RZBR07]. high-dimension [RZBR07]. Higher [LOT00, PSS85, Ban11]. Higher-dimensional [LOT00]. higher-order [Ban11]. Hindmarsh [MC01]. histology [TKBD^{+}19]. historical [Ree97]. Histories [Cuz84, MBEP84, Woo94]. History [Tho85, HR16]. HIV [AMMJ86, BA88a, ALM17, BA88b, BA91, – DD22, GY93, GA94, GH97, GEW00, GBSRS18, Hsi96a, Hsi96b, HC00, HWG11, Kak99, Kak98, LOT00, LdA99, MHMDM17, PFF^{+}20, RNE^{+}20, SNHMM16, UPM^{+}15, WHT12, YC99, ZZ08]. HIV-1 [BA91, SNHMM16, UPM^{+}15, WHT12, ZZ08].

HIV/AIDS [GH97, GEW00, Hsi96a, Hsi96b, HC00, Kak99, Kak98, LdA99]. holds [Kap01]. hollow [PSWO14, SW12]. Homogeneity [BGW88]. homogeneous [PM86, SGBA92]. homologous [MSZMGFS18]. Homosexual [BA88b]. Hopf [Alr19, KN18, KG99, LM86]. horizontal [CKC07]. Hormonal [AS88]. hormone [BCC08]. Host [AM84, CS91, LM86, ARCE^{+}07, CLD^{+}18, GBBT19, JCS15, KKO^{+}11, KKC^{+}03, MPP05, MA07, WW05, YB18, Zha22]. host-directed [Zha22]. Host-Parasitoid [LM86]. host-pathogen [WW05]. Hosts [Agu87, MCRC08]. house [ML00]. house-of-cards [ML00]. HPA [MČM^{+}16]. HPT [PMB14]. HTLV [ETOK03, ETO07]. HTLV-I [ETOK03, ETO07]. Human [AMMJ86, BA88a, BH90, DWLS87, HC84, ST85, BCS97, CGDF02, CW14, HO19, PP21, YB18, RMLC14]. humans [LP05]. humour [DRRT18]. hybrid [And05, BCS97, MDS^{+}24, SSS^{+}20]. hydrocephalus [SBM04, SW06, WDS12, WS06]. hydroxyphenylpyruvate [WDD^{+}17]. Hyperbaric [SSS88]. hyperbolic [TRMB09]. hyperemia [OAG^{+}06]. hypertension [ALDLDMC12, HPR^{+}23, SPL^{+}07]. hypothalamic [BCC08, KN18, MČM^{+}16]. hypothalamic-pituitary-adrenal [BCC08, KN18]. hypothalamus [LP05, PMB14]. hypothalamus-pituitary-adrenal [LP05]. Hypothesis [Bla90, MLH92, Sch85b].

identically [Tho10]. Identification [FM08, SGBCC18, Tsi24]. Identifying [DMNW20, WBSD12]. Identity [Tho88b]. Idiopathic [SPL^{+}07]. IFN [HVPC24]. ignored [Tho10]. II [Seg03, Cai90, CCS99, CW07, EHSALR12, GSBA92, GB86, ILA89, NSCO97, WK99]. IL-2 [BMHK16]. Illustrations [bMO04]. Illustrative [MBEP84]. IMA [Ano91, JKK13]. images [ALDLDMC12, GCC^{+}12]. imaging [BDB^{+}18]. imitation [DLM23]. Immune [SGRPSM^{+}09, FDC^{+}24, FFK17, JCS15, JBG^{+}19, RPSGR^{+}07, SNHMM16, WPH^{+}18]. immunity [WCM98, Yam18]. Immunization [AM84]. immuno [CH20]. immuno-oncology [CH20]. Immunoassay [AC89]. Immunodeficiency [AMMJ86, BA88a]. Immunological [DH87, LBY00]. immunomodulation [JCS15]. immunotherapy [BMHK16, NO94, RPSGR^{+}07, TG23]. Impact [LPSR24, PFF^{+}20, ZZ08, DOM^{+}09, MF21, RMLC14, RTG21, WDS12]. impacting [TKBD^{+}19]. impacts [YJY21]. implanted [TPRS18]. implications [DSMF00, PGBMG^{+}15, TH96]. implicit [TRK19]. Importance [FHC09, And05, Byr97, GLDP17]. impossibility [BB21]. Improved [KLG^{+}14, SAEHM03]. Improving [BMHK16, RT01]. impulses [MC98]. Impulsive [ACS14, LC04, ZLC03]. in-host [Zha22]. in-patient [IMM94]. in-vivo [SBM04]. incidence [BL15, CKC07, KM05, Kor09a, MA06]. including [Kro10]. inclusion [CCV05, WG20]. Incorporating [BS87a, LvdB98]. Incubation [BA88a, Cai90, MK89, GL00]. incus [MEB10]. indentation [YLS^{+}20]. independences [Tho10]. independent [Tho10]. Index [Ano93a, Ano93b, Ano98b, Ano99b, Ano00b, Ano00c, Ano01b, Ano01c, Ano04, SAEHM03, SGRPSM^{+}09, Ano84b, Ano85b, Ano86b, Ano96b]. Indicator [THI91, FHC09, NDd17]. Individual [BvGPB24, FBJK10]. individual-based [FBJK10]. induced [FMHF10, FS17, JBG^{+}19, MD18, MS02, PFF^{+}20, PR16, WGH03]. inducible [LTY15]. Inducing [GF08]. Induction [CBWA91, TTBL21]. inequities [JPB^{+}21]. infected [HG10, Tsi24]. Infection [HSV24, ASY16, BZF^{+}15, Car06, – DD22, ETOK03, GSBA92, GBBT19, GY19, HH08, LvdB98, LOT00, MMA^{+}21, MHMDM17, MRG98, MA06, MHGY18, NBC18, PGL^{+}22, RY22, SNHMM16, SGBA92, TKR^{+}18, Yan02, ZZ08]. infection-dependent [MA06]. Infections [RG91, BMN^{+}93, Kap01, KLG^{+}14, LBY00, NS95, RG92, Yan02]. Infectious [BA88a, O’N13, HA12, KM05, LL00, Par97, Rob99, Yam18, ZJJA19]. Infectiousness [BA88c, Cai90]. Inference [LG24, BSEO07, GGA06, HVPC24]. Inferring [BT15]. infertility [GSBA92, SGBA92]. infestations [MG14]. inflammation [BDJB^{+}13, WA20]. inflammatory [ADCN24, IMRW05, JCS15, SH01, ZJJA19]. Influence [MS05, ADCN24, BG11, CDNR16, FDC^{+}24, GSBA92, GBPF16, GD98, MBB20, MS01, PSWO14]. influenza [AAVAV24, LP08, VHT07, Yam18]. Information [CBR97, Tho84]. Information-entropy [CBR97]. Informed [HF20]. infrared [RMJLSG10]. infusion [ESC13, WS09]. inhalable [MHGY18]. Inhibition [KH87, AMP^{+}23, VDLK13, WDD^{+}17]. Inhibitor [Jäg86]. inhibitory [BK19]. inhomogeneities [RL13]. inhomogeneous [MBS92, PDD08]. initiated [CTB20]. initiation [DPO^{+}15]. injecting [CGH12, GH97]. injured [CKL^{+}05]. injury [FS17, RSA14]. injury-induced [FS17]. input [Don23, Lu11, PP96]. insect [LvdB98]. insect-pathogen [LvdB98]. Insects [CK87]. inserting [MEB10]. insertion [EHSALR12, SAEHLR98]. Insights [RBMH17, AAVAV24, dSCdA16]. Instability [WBGN89, FY19]. instigator [CH20]. instillation [SGZ19]. Instructions [Ano12i, Ano12j, Ano12k, Ano13m, Ano13n, Ano13o, Ano13p, Ano14j, Ano14k, Ano14l, Ano15j, Ano15k, Ano15l]. insulating [CS08]. insulin [GC15, NSCO97, SHO^{+}94]. Interacting [RB87, BMY00, LG24]. Interaction [vdBJ84, BKL09, CHMH98, DCT06, GL00, JCS15, TTBL21, VHT07]. Interactions [FK03, vdBD86, CTB20, FG04, FKCK17, MPP05, MANS11, MSB10, OSMA97]. intercellular [Byr97]. interface [TSE24]. interleukin [BMHK16, WPH^{+}18]. interleukin-2 [WPH^{+}18]. internal [RKAL02]. interplay [BMd19, dSCD20]. interpretation [BCM98]. Interval [BS87a]. intervention [BMd19]. interventions [HPR^{+}23, MDS^{+}24]. Intestinal [KJ88, VJ94, YRCU14]. intestine [TRK19]. intimal [FS17, MMF24]. intra [NDd17]. intra-compartment [NDd17]. Intracellular [THI91, BV20]. intracranial [SPL^{+}07]. Intransigent [DLM23]. intravenous [GY93]. intravesical [SGZ19]. intrinsic [RQA12]. Introduction [Seg03, bNFB05]. intrusion [KKO^{+}11]. invading [ACS14]. invalidity [CB15]. Invasion [AJ12, And05, BTCE18, HBS^{+}12, MPP05, MJB11, MNB06, OS98, Lut06]. invasions [BRR25]. invasive [MCRC08]. Inverse [Has84, WS08]. inverted [KK17]. investigate [HCWK17, Lew92]. investigating [KB07]. investigation [RQA12, RTMT23]. Investigations [BSA^{+}12]. Involving [Pan89]. Ion [BS87a, BMY00, BMY04, BHB19, BHB20, MH97, SDA^{+}23, VJ94]. Ionizing [Tho87a]. Iontophoretic [PLFP17]. iridotomy [DRRT18]. iris [TPRS18]. iris-fixated [TPRS18]. iron [CTB20]. irreducibility [LTW93]. Irreducible [Lud89]. irritant [BT21]. ISBN [O’N13, bMO04, bKP04]. island [Mat08]. isogenic [AC24]. isolated [RBGR^{+}14]. isolating [YJY21]. isolation [OK13]. Isotopy [CH87]. Italy [PFF^{+}20].

J [Bur06]. J. [Seg03]. Japan [Kak99, Kak98]. Japanese [GT99]. joining [MSZMGFS18]. Joint [Tho84, EHE^{+}18, HvBB10]. Joints [BBN90]. Jones [bKP04]. Journal [JKK13]. July [Ano91]. Jump [CD86]. juxtacrine [OK13].

Kermack [Rob99]. Kermack-McKendrick [Rob99]. key [HBS^{+}12]. kinetic [DLM23, LP21, Mat08, TK05, SPN22]. kinetic-type [SPN22]. Kinetics [SSS91, BSA^{+}12, CGG^{+}22, GM19, HMM23, PDR^{+}21]. Kirchhoff [CO19]. Knight [GSW87]. Kolmogorov [EHMS22]. KPP [HZ25]. Krukenberg [BAJ^{+}17].

L [Lut06]. L. [SAEHM03]. lac [Alr19]. lactate [CGG^{+}22, PMMA^{+}21]. lake [BGW99]. Landfill [You90]. Laplace [CB15]. larch [Sch85b]. larch-larch-bud-moth [Sch85b]. Large [AFK87, AKR^{+}89, GT91, DMS25, GH05, SDA^{+}23, SJ03, YBG96]. Large-Amplitude [AFK87]. Large-Scale [AKR^{+}89]. larger [BSEO07]. laser [TRMB09]. last [Ree97]. latency [BMd19]. Latent [Rin90, Mon07]. Lattice [KBVW19, MG14]. laundered [BT21]. law [CB15, WS08]. Laws [PM86]. layered [PLFP17]. layers [WBG22]. LDL [EHSALR12, SAEHLR98]. leachable [TSE24]. Lead [KJ88]. Leading [Sle88]. leaf [SAEHM03]. left [Bra93, MS13, TH96]. left-right [Bra93]. leg [ZOZ^{+}22]. lens [ACH^{+}21, TXAS15, TPRS18]. Lethal [Tho85]. leucine [BSA^{+}12]. leucocytes [BASH20]. leukaemia [BASH20]. leukemia [LJR^{+}19]. leukocyte [BCM98]. leukocytes [OSS05]. Levamisole [DGDW87]. level [WG20]. levels [XZT11]. Lewy [KK22]. Li [Lut06]. lid [HBD^{+}07]. Life [Cuz84, RT01, Woo94, Yam18]. life-table [RT01]. life-time [Yam18]. Lifespans [RB87]. like [Car06, Tan87]. Likelihood [Tho87b, TG91, PDD08]. limb [RL13]. Limit [BK19, LKN96]. Limit-cycle [BK19]. Limited [BS87b, IR85]. limiting [RY22]. line [BRR25, GSW87]. lineage [BK19, BASH20]. Linear [KJ88, BSA^{+}12, CCS99, KM05, Kor09a, MA06, PRWJ07, SW06, WGCB07, WS09, dAL02]. Linkage [Tho94, Tho84, Tho87b]. Linked [Tho85]. lipid [BGG03, MBB20]. lipid-mediated [BGG03]. liver [LCC^{+}17, RSA14, YLS^{+}20]. Living [THI91, XDJ87, RG92]. Load [Phe91b]. loads [Phe95]. local [BFPH17, LP21, MPP05, MR20, THH92]. localization [Cox06, PRK05]. loci [BCS97]. lockdown [SPN22]. locomotion [BCM98]. locus [Tho88b]. Long [Sey90]. longitudinal [FKCK17]. loop [TC03, TC04]. loss [CGP06]. Lotka [LC04, PMN00, SK24]. low [MS13, Noo12, PGBMG^{+}15, PGPR16]. low-dosage [Noo12]. low-grade [PGBMG^{+}15, PGPR16]. low-order [MS13]. LQ [BCPS19]. Lumped [DCRK21]. lung [HM07, MBC^{+}14]. Lyapunov [Kor04]. Lyme [Por99]. lymphoblastic [BASH20, EVK^{+}18, LJR^{+}19]. lymphocytes [AALL23, MCK04]. lymphoma [EVK^{+}18].

M [EHE^{+}18]. M. [ARCE^{+}07]. macroparasite [KKO^{+}11]. macrophage [ADCN24, OS98]. magnocellular [MC98]. maintenance [DPO^{+}15, LJR^{+}19]. major [RB15]. make [GP11]. Mal [BMB^{+}23]. Malaria [Nås86]. Male [BA88b, WP85]. malignant [GF08, MNB06, PR16]. malleus [MEB10]. malleus-incus [MEB10]. Management [RG91, dSCM06, PYLC09]. MANF [CLM^{+}20]. manoeuvre [PTW^{+}17]. mapping [OWL^{+}98]. maps [TC06]. marginal [HvBB10, MDS^{+}24, YBG96]. marina [SAEHM03]. Markov [GR98, KLG^{+}14, LTW93, MK20, Tho10, Ven93, YB03]. marriage [TC03]. Mass [GLDP17, SW12]. Massachusetts [bMO04]. Master [vdBJ84]. material [FK03]. Maternal [ZNB97]. Maternity [MBEP84]. Mathematical [ACC18, Ano91, BDJB^{+}13, BBR84, CS90, CS91, CGP06, CTB20, CF84, CKL^{+}05, EVK^{+}18, FBM^{+}03, GH97, HM07, HVPC24, HA12, JBG^{+}19, LVWG17, LKH98, Mai03, MCK04, MBC^{+}14, MBEP84, NBT^{+}22, bKP04, PP21, PTW^{+}17, RSA14, RS90, RTMT23, RBGR^{+}14, Sch85b, SSS91, SMML92, SSA89, Thi16, WK97, WK99, WKK^{+}01, ZKB^{+}19, And05, BCC08, BZF^{+}15, BFM^{+}17, BCS96, BSEO07, BGG03, BMHK16, CH20, CW14, DGM^{+}18, CCVV98, CCV05, DCC^{+}21, DSMF00, DSW^{+}21, EHE^{+}18, ETOK03, ETO07, FR14, FKK93, GT99, GY19, GDL01, GCC^{+}12, HMMN97, IMRW05, Kak99, Kak98, KKO^{+}11, KPSW09, Koh13, LJR^{+}19, LMMS10, MPP05, MSBM04, MG06, MMT22, MHGY18, NPO14, OK13, OSMA97, OC96, PGBMG^{+}15, Por99, PRK05, RNE^{+}20, RY22, RBMH17, SZB21, SHO^{+}94, TG23, TCC^{+}11, TPS^{+}08, VJ94, WDD^{+}17, YRCU14]. mathematical [ZOZ^{+}22, Seg03, bNFB05, JKK13, O’N13]. Mathematically [CGH12, DJG06, WA20]. mating [VdC01]. Matrices [Pug91, UA96]. Matrix [GP84, Pug91, CPB10, CK04, OSMA97]. Matrix-Geometric [GP84]. matter [PK23]. maturation [BM01]. maximization [CBR97]. Maximum [Ano86c, PDD08]. may [GD98]. McKendrick [Rob99]. Mean [PM86]. Measles [Sch84, Bok93]. Measurement [CHC13, ST89]. measurements [AGPV07, ACG^{+}15]. Measures [GB86, SGB86, RTMT23, SPN22]. Mechanical [YLS^{+}20, CHMH98, FS17, IFP13, LBPF15, NBGWJ18, WCSR06]. mechanics [BBPF13]. mechanism [BK19, CS93, CPB10, EHSALR12, MMMW15, SAEHLR98]. Mechanisms [Sch85a, DGN23, KK17, MHMI06, RQA12]. Mechanochemical [LMM87]. mediated [BDJB^{+}13, BGG03, CHC13, MPP05, RL13, SH01, SL16, VHT07]. medical [TSE24]. Medicine [JKK13, ACRF08]. medullary [SEL17]. Meeting [Gre88]. MEG [Das08]. mellitus [NSCO97]. Membrane [Poz88, EHSALR12, IFP13, PSWO14, Poz05, PDR^{+}21, Ric09, VHT07]. memory [KN18, OSS05]. Men [BA91]. menstrual [CW14]. Menten [HMM23]. messenger [EHE^{+}18]. metabolic [HCWK17, SGBCC18]. metabolism [RSA14]. Metapopolation [IR85]. metapopulations [RR99]. metastases [HR16, LCC^{+}17]. metastasis [FHR20, SN98]. metastatic [EHH12]. Method [DWLS87, Kim84, ST89, CCS94, CB15, Cox06, LTW93, MBDKS08, Sec16, TRK19]. Methodological [MGSS89]. Methods [CBW89, CBWA91, DH87, GP84, GR98, ZNB97]. methyl [BCS96]. mice [HCWK17]. Michaelis [HMM23]. microarray [HKV07]. microbiome [FDC^{+}24]. Microcirculation [ST89]. microcirculatory [SSS^{+}20]. Microorganisms [Tan87]. microparasite [WCM98]. microsatellite [SA02]. microvascular [PRWJ07, Sec16]. microvessel [AS12, SNB^{+}11]. middle [MEB10]. migration [BD20, CPB10, CHC13, LG24, LP21, RBMH17, SZB21, SW98]. migrations [BWSCC11]. minimal [HVPC24, KWW^{+}13]. Minimizing [VM13]. mismatch [TSL94]. Mitochondrial [PSS85]. mixed [BSA^{+}12, VdC01]. Mixing [BCC91, UA96, ED00, IN12, OOB01]. mixture [DF18, GH05, PK23]. mixtures [BP03]. modalities [MK20]. modality [NGK19]. Model [AFK87, AS88, AKR^{+}89, BH86, Bel86, BvGPB24, BBR84, CS90, CS91, CMM^{+}17, CF84, DGDW87, Gre88, Gre90, GMS89, GNA90, ILA89, KH87, LMM87, MBEP84, MGSS89, Pan89, PSS85, Phe91b, Poz88, Poz90, RG91, Sch84, Sey90, SSS91, ST85, TH88, WP85, AGPV07, Alr19, And05, ABMR21, ACC18, AS12, ADH^{+}05, ASY16, AG01, ALM17, BP98, Ban11, BV20, BKF14, BRM93, BFK17, BWWvB04, BD20, BWW05, BM01, BSA^{+}12, BFM^{+}17, BTCE18, BZ93, BMB18, BKL09, BK19, BCSS15, BCSS18, BASH20, BCSS21, BFPH17, BHO96, Bra93, BCS96, BCPS19, BGG03, BMHK16, BMd19, CO19, Car06, CH20, CS93, CSeA02, CPB10, CHC13, CW14, CGG^{+}22, CGS^{+}22, CDNR16, DGM^{+}18, CHMH98, CMM92, DCC^{+}21, DSMF00, DOM^{+}09, DLM23, DH04, DCT06, DSW^{+}21, DPC18, Don23]. model [DF18, ED00, EHMS22, ETOK03, ETO07, Eva00, EVK^{+}18, FR14, FNdO16, FKK93, FS17, FFK17, FHR20, GA94, GH05, GGA06, GY19, GM19, GC08, GC15, Gre92, GCC^{+}12, HYM^{+}24, HG10, HM07, HCWK17, Hsi96a, Hsi96b, HH08, IMRW05, JPB^{+}21, JBG^{+}19, JvdBMH98, JOBF23, JLPB08, KN18, KPSW09, KE95, KLG^{+}14, KWW^{+}13, Koh13, Kor09b, Kro10, Krz01, LvdB98, LvdBTJ02, LJR^{+}19, LKR96, LF11, LBY00, LKN96, LOT00, LP05, LC04, LdA99, LP21, LLZ^{+}17, MPP05, MR20, MHMDM17, MRG98, Mat08, MSBM04, MG06, MLVHCC99, MC01, MF21, Mod97, MA06, MGSS21b, MGSS21a, MS13, Muñ16, Mwa02, MHGY18, NBGWJ18, NS14, NPO14, NW93, OWB10, OK13, O’N99, OSS05, OSMA97, OC96, PMB14, PYLC09, PGBMG^{+}15, PMMA^{+}21, Phe95, PDD08, PvdB98, PS03, Por99, PP96, PB00, PK23, POB23, RNE^{+}20]. model [RJ96, Rob99, RW01, RY22, RG95, RL13, SBM04, SGBCC18, SL16, SK24, SMGW13, SHO^{+}94, SNHMM16, SW06, SHGB17, SG13, Tam99, TSE24, TG23, TCC^{+}11, TTBL21, TS94, TC06, TKR^{+}18, THH92, TPS^{+}08, TPL09, TRMB09, Tsi24, VJ94, VDLK13, VW16, WBSD12, WDD^{+}17, WG20, WAB10, WS06, WS08, WS09, YC99, YRCU14, ZOZ^{+}22, ZHF^{+}13, dAL02]. Model-based [CMM^{+}17, HG10]. Modeling [FY19, YJY21, MDS^{+}24, NBT^{+}22, RTMT23, TRK19]. modelled [Yan02]. Modelling [Bak92, BKW^{+}14, BKF14, BF03, BR98, Bur06, BP03, CVEH10, CD86, CCL^{+}02, EHE^{+}18, FMHF10, FDC^{+}24, GMM^{+}99, GY93, GP12, GS84a, GS84b, JCS15, JvdBD02, KP91, KKC^{+}03, KB07, Kuz13, LVTJ00, LVTJ01, MD18, MČM^{+}16, MMWM11, NO94, NO99, NBC18, OAG^{+}06, OS98, RS90, SPN22, TXAS15, Tho87a, VHT07, WVH09, WG20, WBG22, XZT11, AP22, BDJB^{+}13, CGP06, CTB20, CGH12, CCVV98, DJG06, FHC09, FBM^{+}03, GBBT19, GH97, HVPC24, HvBB10, HA12, JPB^{+}21, JBG^{+}19, KBVW19, LVWG17, LL00, LKH98, LCC^{+}17, LK07, MCK04, MBC^{+}14, PSWO14, PJL^{+}22, PP21, PTW^{+}17, RA12, RSA14, RMLC14, Ric09, RTG21, RBMH17, RBGR^{+}14, SZB21, SW12, SSS^{+}20, SJ03, SPL^{+}07, TMM96, Thi16, TC04, WK97, WK99, WKK^{+}01, WGCB07, WA20, WWWW^{+}14, ZKB^{+}19, bMO04]. Models [AS84, Bel84, Bla90, BA88a, BA88b, BA88c, bNFB05, BCC91, Cuz84, DM89, DH87, Fad85, Gan84, Gre87, GT91, HW86a, LM86, LHZ87, Lut06, MK89, NBGM85, PDR^{+}21, Pug91, Ren85, Rin90, Seg03, Sle88, TG91, WDLS88, WBGN89, ZJJA19, AAVAV24, AALL23, ACRF08, BMY00, BKW^{+}14, BSA^{+}12, Bok93, BG11, BM02, CKL^{+}05, DG09, DCRK21, DMNW20, Egl06, Fad93, FG04, FBJK10, FKCK17, GL15, GR98, GDFH13, GWO^{+}18, GEW00, HBD^{+}07, HvBB10, HMMN97, IMM94, KKO^{+}11, KG99, Kor04, KM05, Kor09a, KK15, LL09, LMMS10, LRW93, Lu11, LP08, ML00, MBS92, MS02, MA07, MMT22, MSB10, OKM^{+}05, OC97, PRK05, Rel05, RT01, RKAL02, SMML92, Smi20, SQGP12, TRK^{+}15, TKR^{+}18, Tsi24, TW00, UH96, Ven02, Woo94, YBG96, Yam18, Zha22]. modes [Don23, OK13]. modifications [DRRT18]. modulated [LTY15]. modulus [CB15]. Molecular [Wei85, CCV05]. monocyte [BASH20]. monocytes [SH01]. monolayers [FBJK10]. monomers [Kuz24]. mononucleosis [HA12]. monotherapy [BZF^{+}15]. monotone [ABMR21]. Monte [GT91, LTW93, YB03]. morphogenesis [CMM92, HM07]. Morphogenetic [LMM87]. Mortality [WBGN89, PG10, Woo94]. mortem [HR16]. most [RTG21]. moth [Sch85b]. Moths [WP85]. motility [CO19, FFK17, KWW^{+}13, SZB21]. motion [ACH^{+}21, BMS04, HBD^{+}07, OKM^{+}05, WBG22]. mouth [HWK97]. movement [EBFR21, NS14]. movements [BAJ^{+}17]. moving [GMM^{+}99, JOBF23]. MRI [PK23]. Multi [AP22, AJ12, ADH^{+}05, CGS^{+}22, HvBB10, MA07, PLFP17, SJ03]. multi-branching [SJ03]. multi-host [MA07]. multi-layered [PLFP17]. multi-physics [CGS^{+}22]. Multi-scale [AP22, CGS^{+}22]. multi-scale/multi-physics [CGS^{+}22]. multi-species [ADH^{+}05]. multi-state [HvBB10]. multi-type [AJ12]. multicellularity [AC24]. Multicohort [HW86b]. Multicomponent [Phe91a]. multiconductance [BMY04]. multicyEinder [KE95]. Multidimensional [VL00, CDNR16]. Multihost [GH99]. multimodal [BW01]. Multiparasite [GH99]. Multiphase [LK07, PSWO14, GWO^{+}18, JOBF23, OWB10]. Multiple [Hor90, KWW^{+}13, PGC^{+}00, Swa85, Eva00, FHC09, Kap01, LKR96, MK20, SGK16, WCM98, ZB02]. multiple-pathway [LKR96]. Multipoint [Tho87b]. Multipopulation [Pug91]. Multiscale [LLZ^{+}17, BKF14, EHS16, ONE^{+}15, Ric09, WBSD12]. Multispecies [MM87, dSCM06, dSCD20]. multistate [RT01]. Multistep [OSS05]. Multistrain [DP08]. mumps [CKS20]. murine [JBG^{+}19]. Murray [Seg03]. muscle [BJ14, Kro10, Ven93]. Muscles [DWLS87]. Mutant [Tho85]. mutation [ML00]. mutational [SA02]. mutations [FY19, DP08]. mutualism [SK24]. Myelinated [Bel84, Bel86]. myeloid [BASH20]. myeloid-derived [BASH20]. Myeloma [Swa85]. myeloproliferative [HVPC24]. myocyte [MGSS21b, MGSS21a]. myocyte-fibroblast [MGSS21b, MGSS21a]. Myxomatosis [AS88].

N [BCS96]. N-methyl-D-aspartate [BCS96]. Nagumo [MC98, ZEM11]. nanoparticle [AP22]. nasal [JC19]. nation [KT93]. national [LdA99]. natural [HR16, Rob99]. Nature [TW00]. navigation [OSS05]. Near [Hor90, Sey90, CS08, EHSALR12, SAEHLR98]. Near-Optimal [Hor90, Sey90]. Necessary [UA96]. necrosis [LP01]. necrotic [Fok12]. needle [GL00]. negative [LTY15, PMB14, VDLK13]. negative-feedback [PMB14]. Neighbourhood [HW86a]. Nematode [DGDW87, RG91, MRG98, RG92]. nematodes [KKO04]. neoplasms [HVPC24]. nephrons [SL16]. Nerve [GS84a, GS84b, Krz01]. nervous [RBMH17]. network [BT15, DAS15, EHE^{+}18, PRWJ07, SKG11]. Networks [BL86, Hop84, XZT11]. Neural [bMO04, vdBJ84, MC01]. neurite [MG06]. neuro [PMMA^{+}21]. neuro-energetics [PMMA^{+}21]. neurochemical [Noo12]. Neurogenic [DWLS87, WDLS88]. Neuron [Hop84, KB07, PP96]. Neuronal [Das08, KH87, VL00]. neurones [KE95]. neurons [CBR97, MC98]. neutralizing [PGL^{+}22]. Niche [GB86, SGB86]. nitisinone [WDD^{+}17]. nitric [OAG^{+}06]. nitrogen [LVTJ00, LVTJ01]. NMDA [BCS96]. node [TC03]. nodes [SKG11]. noise [LC98, PP96, RMK04]. noisy [MJB11]. Non [EHMS22, KM05, WGCB07, ABMR21, BSA^{+}12, BFPH17, FNdO16, GC08, Kor09a, LP21, MR20, MA06, MSZMGFS18, MDS^{+}24, NSCO97, PRWJ07, SW06, dAL02]. non-Fickian [FNdO16]. non-homologous [MSZMGFS18]. non-insulin [NSCO97]. Non-linear [KM05, WGCB07, BSA^{+}12, Kor09a, MA06, PRWJ07, SW06, dAL02]. non-local [BFPH17, LP21, MR20]. non-monotone [ABMR21]. non-pharmaceutical [MDS^{+}24]. non-uniform [GC08]. Non-vanishing [EHMS22]. nonautonomous [BRM93]. Nonexponential [Cai90]. noninvasive [AGPV07]. Nonlinear [Fad85, MGSS89, Sch85a, Alr19, CBR97, LOT00]. Nonlinearities [DH87]. nonlocal [BRR25]. nonpassive [CHMH98]. Nonrandom [HC84]. Nonself [DH87]. nonspecific [UH96]. nonswelling [LKH98]. nonuniform [LK07]. normal [MS01, MS05, SGK16, SHGB17]. nosocomial [KLG^{+}14, TKR^{+}18]. Note [DS91]. novel [ÇÇ17, CBPD21, KKO^{+}11, MMT22, NPO14, PGPR16, Tsi24]. NTCP [SHGB17]. null [HCWK17]. Number [Ano86c, BCPS19, CCS94, CCS99, CMR^{+}04, DMNW20, KKO04]. numbers [BP98]. Numerical [Car06, CCVV98, DH87, DH04, GBSRS18, KP91, Muñ16, Bok93, BHO96, EBFR21, FG04, MR20, MSBM04, PMMA^{+}21]. nursery [JvdBD02]. Nutrient [MANS11, CW07, LK07, ONE^{+}15, SMGW13]. Nutrient-rich [MANS11].

O [O’N13, SSA89]. occlusion [LLZ^{+}17]. occurrence [CKC07]. occurring [Ric09]. OCTA [CGS^{+}22]. ocular [ALDLDMC12, LBHKS18]. oedema [LVWG17, YRCU14]. Offensive [AS86]. Offensive-Defensive [AS86]. offs [BD20]. oleic [CKL^{+}05]. oligomeric [Smi20]. Omission [BS87a]. Omsk [ALM17]. oncogenesis [FY19]. oncological [SGRPSM^{+}09]. oncology [CH20]. One [BW87, SSS91, GP11, HZ25, ZLC03]. One-Dimensional [BW87, HZ25]. one-predator [ZLC03]. One-Step [SSS91]. onset [BFM^{+}17]. Ooyen [bMO04]. Open [BL86, ABMR21, AG01]. Openings [BS87a]. operation [TH96]. operative [TPS^{+}08]. operativity [BKL09]. Operon [Alr19]. opinions [DLM23]. opioids [RTG21]. optic [TRMB09]. Optimal [BS87b, BCPS19, BMd19, Cai89, CLM^{+}20, CKS20, CBB92, CBB94, HW86a, HW86b, Hor90, JLPB08, Lu11, MS01, MG14, NGK19, Ree88, SGK16, Sey90, Swa85, Swa87, Swa88, Tho86b, BC00, CBR97, CMM^{+}17,

– DD22, GBSRS18, Hor96, LMMS10, MMA^{+}21, MS05, Mur95, Mur97]. Optimality [TK05, HO19]. Optimization [KP91, HYM^{+}24, SGRPSM^{+}09]. optimizing [MK20]. Order [HPH90, SSS91, Ban11, BCSS15, BCSS18, BCSS21, KSS16, MS13, PK23, Tho10]. ordering [ALDLDMC12, BCS97]. ordinary [BSA^{+}12]. organisms [BMS04]. organization [CCV05, GL15]. orientation [TC06]. origin [She92]. original [NBC18]. Oscillations [Hop84, AS12, BV20, Kee06, LOT00, MS02, Noo12, WGH03]. Oscillator [KH87, BMB^{+}23]. Oscillatory [ZEM11, BK19]. osmolarity [LBD^{+}16]. other [PP21]. outbreak [CKS20, RB15]. Outbreaks [AKR^{+}89]. outcome [RPSGR^{+}07]. overcoming [CCV13]. Overlap [GB86, SGB86]. overset [MBDKS08]. Overview [MBEP84]. Oxford [Ano91]. oxidative [PP21]. oxide [OAG^{+}06]. Oxygen [SSSK91, SSS91, SEL17, WGH03]. Oxygenation [SSS91, BHO96, PDR^{+}21].

p53 [DAS15]. packed [SW12]. paediatric [BASH20]. Page [Ano12d, Ano12a, Ano12e, Ano12b, Ano13e, Ano13a, Ano13f, Ano13b, Ano13g, Ano13c, Ano13h, Ano13d, Ano14d, Ano14a, Ano14e, Ano14b, Ano14f, Ano14c, Ano15d, Ano15a, Ano15e, Ano15b, Ano15f, Ano15c]. Pages [bMO04]. pairwise [Tho10]. pancreatic [LKN96]. pandemic [LP08, NBT^{+}22, RB15]. Paper [JKK13]. Papillomavirus [RMLC14]. paracrine [TKBD^{+}19]. Paradox [KSS16, BFPH17]. Parameter [AGPV07, ESC13, BM01, DCRK21, MMMW15, MSBM04, TRK^{+}15, TKR^{+}18]. parameters [DCT06, GR98, Mod97, Ven93]. Parametric [Bel86, Cuz84, MSB10, SGBCC18]. Parasite [DGDW87, Kor09a]. Parasitism [Has84]. parasitizing [KKO04]. Parasitoid [LM86]. parasitoids [MCRC08]. Part [CW07, UH96, WCSR06]. Partial [PMMA^{+}21, WGH03]. particle [ZHF^{+}13]. Particles [Tho87a]. partnership [GA94, Mod97]. Passive [Poz88, KE95, Kro10, VM13]. patch [ASY16, HZ25]. Patchy [Has84]. Paternity [Bro89]. Pathogen [CLD^{+}18, LvdB98, TRK^{+}15, WW05]. Pathogens [Agu87, KKC^{+}03, PvdB98]. pathway [LKR96, MSZMGFS18]. pathways [AMP^{+}23]. Patient [CHMH98, WWWW^{+}14, GCC^{+}12, IMM94, PMB14, TMM96]. Patient-specific [WWWW^{+}14, GCC^{+}12, PMB14]. Patient-ventilator [CHMH98]. patients [EVK^{+}18, NSCO97]. Pattern [GWO^{+}18, JOBF23, MBS92, MO84, Ban11, Bra93, CMM92, HO19, NW93]. patterning [OK13]. Patterns [BA91, GMS89, Jäg86, EHSALR12, LPSR24, MLH92, OOB01, RTMT23, SAEHLR98]. Paulo [YJY21]. Pea [WP85]. Pedigree [Tho86a, Tho86b, WL89, HT94, Har95, LTW93]. pedigree-based [Har95]. Pedigrees [GT91, Tho88a, Tho88b, She92, TSL94, Tho94, TC03, TC04]. peeling [Har95]. Pennes [RMJLSG10]. Penrose [MLH92]. perfect [Tho93]. performance [SBM04]. perfusion [OWB10, SMGW13]. Period [Rin90, Sey90, Sle88, BMd19, GL00, Mon07, Par97]. Period-doubling [Sle88]. Periodic [AFK87, FK02, KH87, RG91, WGH03, Bak92, DM25, FK00, LC04]. periodicity [SKTO93]. Periods [BA88a, Cai90, MK89]. Peristaltic [IN12]. permanence [ZLC03]. permeability [SW06]. persons [YJY21]. Perturbation [BL86]. Perturbations [Hor84b, RG91, LC98, SK24]. Pest [PYLC09, dSCdA16]. pests [ACS14]. PET [ACG^{+}15, GCC^{+}12]. Petrovskii [Lut06]. phage [TTBL21]. phage-bacteria [TTBL21]. pharmaceutical [MDS^{+}24]. pharmacokinetic [FHC09]. phase [DF18, OKM^{+}05, PRK05]. phases [DMO25]. phasic [Noo12]. Phenomena [LMM87, DM25]. Phenomenological [SDA^{+}23]. phenomenon [RNE^{+}20]. Phenotype [AC24, ADCN24, HYM^{+}24]. phenotype-structured [HYM^{+}24]. Phenotypic [Bur86, AJ12, CPB10, She92]. Pheromone [WP85]. photoreceptor [CLM^{+}20]. phototaxis [FMHF10]. phylogenetic [BSEO07]. Physarum [IN12]. physics [CGS^{+}22]. Physiological [RK01, GMM^{+}99]. pigmentation [GL15]. pigs [CKL^{+}05]. pits [EHSALR12, SAEHLR98]. pituitary [BCC08, KN18, LP05, MČM^{+}16, PMB14]. Planci [Sey90]. Planck [BFK17]. plankton [BWWvB04, MANS11]. planning [CMM^{+}17]. plant [HMMN97, JvdBMH98, JvdBD02, PvdB98]. plant-disease [HMMN97]. plant-virus [JvdBMH98]. plantation [JvdBD02]. plantation-nursery [JvdBD02]. Plants [AS86, PSS85]. plaque [EHSALR12, Fok12]. plaque-form [EHSALR12]. plaques [Kuz24]. plasma [Kap01, PDR^{+}21, RY22]. plasmodium [IN12]. Platelet [FG04, TH88, DF18, LF11, LLZ^{+}17]. Platelet-wall [FG04]. plexuses [CGS^{+}22]. plus [GP11]. points [KGG^{+}99]. Poisson [Rln85]. Polar [She92]. Policies [Cai89, HF20]. policy [dSCM06, Tsi24]. Polymer [Ano86c, TSE24]. Polymer-interface-tissue [TSE24]. polynomial [RW01]. Polyploids [HPH90]. pool [Thi16]. Population [ETOK03, Gan84, Hib85, KR90, LHZ87, Lou89, MM87, MS02, MPE87, Mod88, Oht84, PM86, RG91, Sey90, WBGN89, ARCE^{+}07, BRM93, BGW99, BWSCC11, BTCE18, DH04, GSBA92, HYM^{+}24, Hsi96a, Hsi96b, LRW94, NW93, RNE^{+}20, RG92, Rob99, RT01, SGBA92, TW00, VDLK13, Woo94, YC99]. Population-induced [MS02]. Populations [AM84, BA88b, Cai89, Cai90, NBGM85, AJ12, AG01, CBB92, GA94, GD98, HvBB10, Hor96, HH08, SDA^{+}23, SGBA92, UA96, VW16, Woo94, ZB02]. porcine [YLS^{+}20]. Poro [BBN90]. Poro-elastic [BBN90]. poroelastic [WS06, WS09]. porosity [LK07]. porous [KBVW19, LKH98]. poroviscoelastic [WG20, WBG22]. posedness [Yam18]. possums [Rob92]. Post [LMM87, Sch84, HR16, TC06, TPS^{+}08, WAB10]. post-blink [WAB10]. Post-fertilization [LMM87]. post-mortem [HR16]. post-operative [TPS^{+}08]. post-synaptic [TC06]. Post-Vaccination [Sch84]. posterior [DRRT18, GBPF16]. Postsynaptic [Poz90]. Potential [PJL^{+}22, CS08, MGSS21b, MGSS21a, NPO14]. potentials [Ric09]. power [WS08]. pp [O’N13]. practical [CB15]. Pre [Sch84, – DD22, TC06, VHT07]. Pre- [Sch84, TC06]. pre-exposure [– DD22]. pre-fusion [VHT07]. precorneal [BF03, WAB10]. Predation [AFK87, BWM97]. Predator [HH08, KR90, vdBD86, dSCM06, dSCD20, HG10, Kor09b, MANS11, PYLC09, RKAL02, Ven02, ZLC03]. predators [ACS14, Ven02]. predictability [SKTO93]. Predicting [HBHJ22, WDD^{+}17, CBPD21, KK15]. Prediction [YF00]. Prefecture [LPSR24]. preference [TC06]. preferred [CO19]. Preliminary [AMMJ86]. Prepatterns [NM92]. prescribed [RTG21]. prescription [RTG21]. presence [CSeA02, GMM^{+}99, Kap01, LC98, SPN22]. Press [O’N13, bMO04, bKP04]. Pressure [BS87b, CVEH10, CKL^{+}05, GH05, MBHKS10, NDd17, WGH03, WG20]. Pressure-Limited [BS87b]. prevalence [Bak92, BL15, VHBCC96]. prevalence-dependent [VHBCC96]. prevention [GBSRS18]. previous [EHS16]. previously [IFP13]. Prey [KR90, vdBD86, BWSCC11, HG10, HH08, Kor09b, MANS11, PYLC09, RKAL02, Ven02, ZLC03]. prey-predator [PYLC09]. primordial [DSW^{+}21]. Princeton [O’N13]. principle [LOT00, TK05]. Principles [GS84a]. prions [CMR^{+}04]. prison [YBG96]. Prize [JKK13]. Pro [SH01]. Pro-inflammatory-anti-inflammatory [SH01]. probabilistic [ÇÇ17, TS94]. probabilities [RTG21]. Probability [Pan89, Tho86a, Tho86b, WDLS88, BMB18, GDFH13, GC08, ML10, RB15, SHGB17]. Probe [BBR84]. probing [Mag92]. probiotic [FDC^{+}24]. Problem [BCC91, Lou89, Swa87, BCS97, KT93, Mur97]. Problems [RS90, CW07]. Process [CD86, SSS88, FHR20, MK20, Mod97]. Processes [NBGM85, PM86, Ren98, Ric09, SGBCC18, Sey95, SQGP12, VL00]. processing [DM25]. produced [NM92]. Production [CD86, CS90, CF84, TH88, ZZ08]. progenitor [DGM^{+}18, RNE^{+}20]. prognosis [DCC^{+}21]. program [HWG11]. programme [LdA99]. Programmes [AM84, BMN^{+}93]. progress [HMMN97]. progression [BFM^{+}17, CBPD21, DBG14, EHH12, FR14, HBS^{+}12, MGDPC^{+}15]. Projection [PM86]. Projections [MPE87, HvBB10]. proliferating [FK03]. proliferation [EHS16, SMGW13]. promiscuity [BK19]. promoter [BKL09, WA20]. promoting [EHH12]. proof [PGPR16]. propagation [MGSS21b, MGSS21a]. Properties [GJ90, BMY04, BRM93, BT15, FK03, FFK17, Kor04, Kor09a, Kor09b, KB07, MHMDM17, SK24, YLS^{+}20]. prophage [TTBL21]. prophylaxis [– DD22, RB15]. proportionate [ED00]. proposed [BCSS15, BCSS18, BCSS21]. prosthesis [MEB10]. protection [YC99]. protective [MHGY18]. protein [DCT06, KK17, VHT07]. protein-mediated [VHT07]. protein-protein [DCT06]. proteolysis [MPP05]. proteolytic [CCV05]. protocol [BG11]. protocols [LMMS10]. proton [MD18]. protraction [PGPR16]. PRRSv [GBBT19]. Prüfer [TC04]. Pseudomonas [MBC^{+}14]. psychobiology [FKCK17]. PU.1 [BKL09]. PU.1-GATA-1 [BKL09]. public [BMd19, HWG11]. publication [LG24]. puffs [DGS08]. Pulmonary [SSS91, TH88, HPR^{+}23]. pulse [NS95]. pulses [ZEM11]. pupillary [DRRT18].

Quadratic [AC89]. Qualitative [GS84b, HWG11]. Quantifying [AMP^{+}23, FKCK17]. quarantine [SPN22]. Quasi [LvdBTJ02]. Quasi-steady [LvdBTJ02]. queueing [CLD^{+}18]. quickly [ZOZ^{+}22]. quorum [WKK^{+}01, WKK^{+}04].

R [WBSD12]. Rabbit [AS88]. Rabbit-flea [AS88]. rabies [EP01]. radially [SAEHLR98]. Radiation [MSjH^{+}85, Tho87a, BCS97]. radiofrequency [CMM^{+}17]. Radiotherapy [MSjH^{+}85, BFK17, BCPS19, NGK19, PGBMG^{+}15, SGK16, SGRPSM^{+}09]. Radiotracer [ST89]. rainsplash [PvdB98]. Rall [Krz01]. Random [Mod88, Ree88, PS03, RG95, Tho10, Ven93]. random-walk [RG95]. Randomly [Ano86c]. Rank [Phe91a]. rat [OAG^{+}06, SEL17]. Rate [Gre88, Gre90, SSS91, Alr19, Gre92, Kor09a, MA06, NSCO97, PR16, TMM96]. Rates [CB84, BMN^{+}93, BR98, Lu11, PG10, RTG21, SJ03, Woo94]. ratio [Ban11, LBY00, SGRPSM^{+}09]. ratio-dependent [Ban11]. rational [dMS08]. Ratios [TG91]. rats [WDD^{+}17]. Ray [DS91]. re [DPO^{+}15, Yan02]. re-entry [DPO^{+}15]. re-infection [Yan02]. Reaction [AKR^{+}89, BL86, GS84a, GS84b, BMM93, HSV24, KK15, MBS92, NM92, OAG^{+}06, Sec16]. reactions [TK05, SPN22]. reactive [OKM^{+}05]. reading [ABMR21]. realistic [HBD^{+}07, Sec16, TC04]. Realizable [LMMS10]. reattachment [IFP13]. Receptor [BH86, CHC13, EHE^{+}18, GC15, MPP05, SH01, Smi20]. receptor-dependent [EHE^{+}18]. receptor-mediated [CHC13, MPP05]. Receptors [Ano86c, BCS96, EHSALR12, SAEHLR98]. recombinant [CCVV98]. reconstructed [MEB10]. Reconstruction [DS91]. Recruitment [MGSS89, VHBCC96]. Rectangular [MSjH^{+}85]. recurrence [MD18]. Recurrent [Gre87]. red [BHO96, Poz05]. reducing [HYM^{+}24, MHGY18]. reductase [HCWK17]. reduction [KSS16, PJL^{+}22]. Reef [AKR^{+}89]. Reference [BBN90]. reflex [MBDKS08]. refractory [HPR^{+}23]. regimen [BC00]. regimens [MS01, MS05]. Regional [ST89]. regions [TS94]. registry [MSB10]. regular [Egl06]. regulate [Mag92]. regulation [BV20, Bra93, DSW^{+}21, LRW94, WPH^{+}18]. regulations [LTY15]. regulatory [SGBCC18]. reinforced [PS03]. relapse [EVK^{+}18]. Related [WBGN89, BMN^{+}93, OC96]. Relating [NBGWJ18, Tsi24]. relation [BL15]. Relationship [BAJ^{+}17, CB84, ABMR21]. relationships [Don23]. relative [WBG22]. Relaxation [THI91, LOT00]. Release [MMMW15, DGS08, LKH98, OAG^{+}06, TSE24, VM13]. releasing [BCC08]. reliable [NDd17]. remains [PGC^{+}00]. remodeling [MF21]. remodelling [ÇÇ17]. removal [MA06]. remove [BV20]. Renal [SEL17]. renewable [Mag92]. Renewal [Ano86c]. reorganization [BJ14, GP12]. repair [BMB18, DSMF00, MSZMGFS18]. replication [Tam99]. replicator [KNB11]. representations [HT94]. representing [SGBCC18]. repression [WKK^{+}04]. Reproduction [AS88, Kor09a, LBY00]. reproductive [Thi16]. require [BKL09]. rerouting [Kuz13]. Research [AS84]. reservoir [MBB20]. Resilience [Agu87]. Resistance [DGDW87, BC00, CBB92, CBB94, HYM^{+}24, LCC^{+}17, Mur95, Mur97, Sey95, TCC^{+}11]. resistant [ARCE^{+}07, UH96, WW05]. resolution [Muñ16]. Resource [Bla90, Mag92, ZB02]. respiration [SGBCC18]. respiratory [AGPV07, FKK93, WG03]. Response [Hor84b, Hor84a, vdBD86, CS93, FDC^{+}24, IMRW05, JCS15, JBG^{+}19, LVTJ00, LVTJ01, LBY00, MCK04, PGBMG^{+}15, PP96, PTW^{+}17, RPSGR^{+}07, SNHMM16, WPH^{+}18, ZJJA19]. responses [BR98]. responsible [KK17]. responsiveness [CGP06]. Resting [Poz05]. restraining [Tsi24]. Results [Gre87, Gre88, Rin90, BCM98, CCS99, Gre92, PMMA^{+}21]. resuscitation [TRK19]. retina [BBPF13, Egl06, LBPF15]. retinal [CGS^{+}22, GBPF16]. retrograde [EHSALR12]. reveals [BD20, NPO14]. Review [bNFB05, Bur06, Cuz84, Lut06, O’N13, bMO04, bKP04, Seg03]. Revisited [Gan84, BZ93]. Revisiting [ASY16, MANS11]. Rewards [Hor90]. rheumatoid [BDJB^{+}13]. rich [MANS11]. right [Bra93]. rigid [RK01]. Risk [BCC91, Hor96, MHGY18, OWL^{+}98]. Risk- [BCC91]. Risk-sensitive [Hor96]. risks [MD18]. River [CCL^{+}02]. Robust [LL09]. robustness [AJ12]. rod [CO19]. Role [CKC07, MK89, WP85, AJ12, BCS96, BJ14, CGP06, CK04, EHH12, FK00, HWG11, MANS11, RTMT23]. Root [Ren85]. Rose [MC01]. Rose-Hindmarsh [MC01]. Rosenberg [NO94]. Rosenberg-type [NO94]. Ross [CCL^{+}02]. rotating [CW07, WCSR06]. Rotation [KP91]. routes [LvdB98]. Ruminants [RG91, MRG98, RG92].

S [Lut06, bKP04]. S. [JC19]. saccadic [BAJ^{+}17]. Saddlepoint [Ren98]. Saharan [NBT^{+}22]. sample [Ree97]. Sampling [HC84, She92]. São [YJY21]. SARS [MDS^{+}24, PGL^{+}22, SPN22]. SARS-CoV-2 [MDS^{+}24, PGL^{+}22, SPN22]. saturation [CBB94, WK99]. scaffold [KBVW19]. scaffolds [LK07]. Scale [AKR^{+}89, AP22, GBBT19]. scale/multi [CGS^{+}22]. scales [VV24]. scanning [TSL94]. scenario [GL00]. schedule [Mur95]. scheduling [Mur97]. Scheme [Poz90]. Schemes [WBGN89]. season [AAVAV24]. seasonal [GD98]. seasonality [BMd19]. Seasonally [Mon07, RG92]. second [MD18]. secondary [EHE^{+}18, GY19, OC96, RY22]. Secration [CS90]. secreted [EHH12]. secretion [BCC08, RW01, SHO^{+}94]. section [Ven93]. sectional [HR16]. Sedimentation [RS90, CHC13]. SEI [MS02, RJ96]. SEIR [AAVAV24, Gre92, Kor04]. SEIR/SLIR [AAVAV24]. SEIS [Kor04]. Selection [CBW89, Kim84, Phe91a, HVPC24, MHMI06, NGK19, RZBR07, TKR^{+}18]. Selective [BS90, CK93, BB21]. Selector [VV24]. Self [DH87, EHH12, GL15]. self-metastatic [EHH12]. Self-Nonself [DH87]. self-organization [GL15]. Semi [Cuz84, MSB10]. Semi-parametric [Cuz84, MSB10]. senile [Kuz24]. sensing [WKK^{+}01, WKK^{+}04]. sensitive [Hor96]. Sensitivity [BFK17, BM01, CDNR16, FM08, GL15]. sensory [CBR97]. separate [BHO96]. separation [LOT00, ZNB97]. sequences [Har95, TS94]. Sequential [CMM92, IC04, BZF^{+}15, SGBCC18]. Series [DH87, CC08, FKCK17, RMK04, YB18]. serological [MMA^{+}21]. serotype [MMA^{+}21]. serotype-specific [MMA^{+}21]. set [VV24]. settings [EHS16]. settlement [FHR20]. Several [Hor90, MS01, MS05]. sex [Hsi96a, Hsi96b]. Sexual [BA88b, BA91, GSBA92, GA94, PYLC09]. sexually [Mod97, dAL02]. shape [NBGWJ18, Poz05]. shaped [LBD^{+}16, MBHKS10]. sharing [ALM17]. sharp [EHMS22]. sharp-fronted [EHMS22]. shear [GFK07, MMF24]. Sheep [DGDW87]. Shock [PRWJ07, DCC^{+}21]. short [FKCK17, NBT^{+}22, RMK04]. short-term [NBT^{+}22]. Shortening [CS08]. should [Don23, GBBT19]. shunt [KE95, SBM04]. SI [MS02, VW16]. SI-epidemic [VW16]. SI-SEI [MS02]. Siberia [ALM17]. sickness [AG01]. siderophore [CTB20]. siderophore-iron [CTB20]. Sighting [Hib85]. Signal [CBR97, DM25, SL16]. signal-processing [DM25]. signalling [AMP^{+}23, EHE^{+}18, GC15, LLZ^{+}17, OK13, TKBD^{+}19, WBSD12, WKK^{+}04]. signed [HKV07]. silico [MGDPC^{+}15]. Simple [MK89, MF21, Rel05, CCS94, GY19, Mod97, Mur97, SDA^{+}23, TW00]. Simulated [TKBD^{+}19]. Simulating [TC04, KK22]. Simulation [CBW89, CBWA91, MBEP84, Ren85, WP85, Car06, EBFR21, HPR^{+}23, LLZ^{+}17, Mon07, NO94, NO99, Sec16, Tho94, TC03, YLS^{+}20]. simulations [PMMA^{+}21, PG10]. Simultaneous [TC06, PDR^{+}21]. sin [NBC18]. since [Ree97]. Single [ACRF08, HBD^{+}07, BMY04, Mur95]. Single-cell-based [ACRF08]. Single-equation [HBD^{+}07]. Singular [BL86, Don23]. singularity [PRK05]. sinus [SPL^{+}07]. SIR [AAVAV24, CKS20, DMNW20, SKG11, dMS08]. SIRD [Tsi24]. SIS [ASY16]. site [DGS08]. Sitka [Ren85]. situ [FBM^{+}03]. Sixth [Ano91]. Size [BS90, MSjH^{+}85, Mod88, DBG14, Hsi96a, Hsi96b, UA96]. Size-Selective [BS90]. sizes [BCPS19]. Skeletal [DWLS87]. skin [CMM92, LKR96, NM92, THH92]. SLAIR [AAVAV24]. Sleeman [bKP04]. sleep [AGPV07]. sleeping [AG01]. slice [OAG^{+}06]. SLIR [AAVAV24]. slow [DPO^{+}15, KK15, LKN96]. slow-fast [LKN96]. small [TRK19]. smooth [BJ14, Kro10]. Social [Bla90, RTMT23]. Sodium [VJ94]. soil [LVTJ00, LVTJ01]. soil-borne [LVTJ00, LVTJ01]. solid [And05, BM02, BP03, CGP06, MCK04]. Solute [HMM23, Sec16, SGZ19]. Solution [BCC91, FG04, KT93, Krz01, KK22, RMJLSG10, WS09]. Solutions [AFK87, MM87, BHB19, BHB20, DMS25, EHMS22, GEW00, KNB11, KE95, KWW^{+}13, MR20, MMMW15, MSBM04, SJ03]. Solvable [Lut06]. somatic [KE95]. Some [AS84, BRM93, Bel84, BS90, Gan84, Gre92, RS90, Sey95, UH96, WDLS88, dSCdA16, TW00, Whi20]. somitogenesis [MSBM04]. source [BMM93, ZNB97]. southern [KT93]. Southwestern [CCL^{+}02]. Space [IR85, GC15, HMMN97, LL00, MSB10]. Space-Limited [IR85]. space-time [LL00, MSB10]. spacing [Egl06]. Spain [AAVAV24]. sparks [DGS08]. Spatial [AM84, Ban11, GMS89, LCC^{+}17, NW93, Ren85, Seg03, ACS14, ADH^{+}05, CHV93, DMS25, LF11, LRW93, NS14, OK13, PB00]. spatial-temporal [LF11]. spatially [DH04, LK07, MBS92, NBGWJ18]. spatio [MCK04]. spatio-temporal [MCK04]. Special [BBN90]. Species [Hor90, Mod88, RB87, ADH^{+}05, BWM97, CCV13]. specific [GCC^{+}12, MMA^{+}21, PMB14, WWWW^{+}14]. Spectral [HKV07]. Speed [Bel86, LP21]. Sperm [BP98, CO19]. Sperm-competition [BP98]. spherical [DP16]. spherically [SW06, WS09]. spheroid [DGM^{+}18]. spindle [BAJ^{+}17]. spines [KE95]. Spline [Woo94]. Splines [AC89]. Spontaneous [AS12, IFP13, Poz05]. spots [NM92]. Spread [BA91, BCC91, ALM17, CGH12, EHS16, EP01, GY93, GH97, HC00, JPB^{+}21, Kak99, Kak98, LRW93, MMT22, XZT11]. spreading [HSV24, SPN22]. sprout [OC96]. Spruce [Ren85]. Square [MSjH^{+}85]. Squeeze [BBN90]. squid [GD98]. Stability [Alr19, ED00, Gre87, Gre88, HPH90, KN18, Kor09b, LC98, LP21, MG06, MC98, Rin90, SNB^{+}11, Ban11, CDNR16, DP16, KM05, LvdB98, LP01, LRW93, RR99, RKAL02, VDLK13, WHT12, WCSR06]. stabilization [WVH09]. stabilized [MANS11]. Stabilizing [Has84]. Stable [Jäg86]. Stage [NBGM85, WBGN89, GL00, Woo94]. Stage-Structure [NBGM85, WBGN89]. stages [RG92, RB15]. Starfish [AFK87, AKR^{+}89]. State [GMS89, RA12, YJY21, GC15, HvBB10, LvdBTJ02, MEB10]. states [NW93, SGBA92]. Statistical [Wei85]. Statistically [Mod88]. Stature [BH90]. Steady [GMS89, LvdBTJ02, NW93]. Steady-State [GMS89]. Stein [Lu11]. stem [BFPH17]. stenosis [SPL^{+}07]. stent [MMMW15]. stented [SWWC09]. stents [MMWM11]. Step [SSS91]. steps [OC96]. stiffness [TH96]. stimulated [OAG^{+}06]. stimuli [FKCK17]. Stochastic [BMY00, BV20, Bla90, – DD22, Fad85, GP84, GSW87, GNA90, IMM94, Kee06, MRG98, MH97, MGSS89, MMT22, TH88, BKF14, BSA^{+}12, BZ93, BMB18, Car06, Fad93, FHR20, GR98, GGA06, Mod97, OSS05, PR16, PG10, PDD08, Ren98, SK24, SHGB17, Zha22, Bur06]. stochasticity [MSZMGFS18, Phe95]. Stock [HW86b]. Stocks [Hor84b]. strain [CB15, MEB10, PMAV23]. strain-stress [MEB10]. strained [ÇÇ17]. strains [WCM98]. strand [MSZMGFS18]. strategies [BD20, CK93, NPO14, OC97, PS03]. Strategy [PR16, CBPD21, Mag92, PGPR16]. stratification [BGW99]. stratified [GA94]. stress [CB15, DAS15, MEB10, NBGWJ18, PP21]. Stresses [BBN90, MMF24]. stretch [GH05]. stripes [NM92]. strongly [BD20]. structural [CK04, SNB^{+}11]. Structure [NBGM85, PSS85, Ren85, Rln85, WBGN89, vdBD86, GD98, Lew92, Don23]. Structure-function [Don23]. Structured [BCC91, Gre87, Gre88, Sch84, DH04, HYM^{+}24, JPB^{+}21, LRW94, Woo94]. Studied [Kim84]. Studies [BGW88]. Study [AMMJ86, AFK87, KJ88, MGSS89, BCC08, Bok93, BAJ^{+}17, CGS^{+}22, CBB94, DPO^{+}15, GT99, HG10, KSS16, MBDKS08, Mon07, MDS^{+}24, PMAV23, PJL^{+}22, SPL^{+}07]. Studying [ALDLDMC12]. sub [NBT^{+}22]. sub-Saharan [NBT^{+}22]. Subject [Ano93b, Ano00c, Ano01c, ST85, CCV13, Ree88, SNB^{+}11]. sublethal [BMB18]. Subpopulations [BCC91, IR85]. substances [CSeA02]. Substitutional [Phe91b]. substrate [BMM93]. Subthreshold [PP96]. suggests [NPO14]. Suicide [BMM93]. sulcus [HO19]. super [MHMDM17]. super-infection [MHMDM17]. Superinfection [Nås86, MLVHCC99, Par97]. Superparasitism [DM89]. supplementation [FDC^{+}24]. support [BSEO07]. suppression [PFF^{+}20, RKAL02]. Surface [EHSALR12, SAEHLR98, BHB19, BHB20, LP01, LBHKS18]. surgery [TPS^{+}08]. surgical [HPR^{+}23]. Surname [Pan89]. Surrounding [CS91]. surroundings [FK03]. Surveys [Hib85]. Survival [LHZ87, PG10, BWM97]. Survivals [Hor84a]. swelling [WG20]. switch [BKL09]. switching [CPB10]. symmetric [SW06, WS09]. symmetries [BM02]. Symmetry [XDJ87]. Synapse [BH86]. Synaptic [KH87, TC06]. Synchronization [BB21, DAS15]. Synchronized [Hop84]. Synchronous [CBW89, CBWA91]. synchrony [GLDP17]. syndrome [NDd17, BMB^{+}23]. Synovial [BBN90]. synthetic [CBPD21]. syringe [ALM17]. syringe-sharing [ALM17]. System [KR90, CKC07, DMS25, FNdO16, FM08, GC15, JvdBD02, KSS16, LvdB98, LG24, LC98, NM92, RBMH17, SGRPSM^{+}09, VM13, YBG96, ZLC03]. system-tumour [SGRPSM^{+}09]. systematic [ADCN24]. Systemic [SSSK91, SSS88]. Systems [Ano91, Bur06, GS84a, GS84b, Hor84a, Jäg86, XDJ87, BMY00, CCS99, DGN23, GH99, LOT00, MMMW15, PJL^{+}22, VdC01, WW05, ZEM11].

T [O’N13, AALL23, EVK^{+}18, MCK04, RBMH17]. T-cell [EVK^{+}18]. T-lymphocytes [MCK04]. Table [Ano12l, RT01]. taper [Eva00]. Tapering [Poz88, Poz90, PP96]. targets [WBSD12]. tau [KK15, KK17]. Taylor [FHC09]. TBU [BDB^{+}18]. TDT [IC04]. Tear [ACH^{+}21, MBHKS10, BF03, BDB^{+}18, HBD^{+}07, JPM^{+}05, LBD^{+}16, MBB20, TXAS15, WAB10, ZKB^{+}19]. tearing [MBDKS08]. Techniques [CBW89, CBWA91, CCVV98, Lew92]. temperature [GD98, TPS^{+}08]. template [LG24]. Temporal [AM84, BS87a, LF11, MCK04, THH92]. tension [LP01, SEL17]. term [NBT^{+}22]. Terminal [GSW87]. termination [DPO^{+}15]. terms [RPSGR^{+}07]. Test [BGW88, CC08, ESC13, WS09, YB03]. testing [PK21]. testosterone [RW01]. TGF [POB23, WBSD12]. TGF- [POB23]. th [HPH90, SSS91]. th-Order [HPH90, SSS91]. thalamic [KB07]. their [RTG21]. theoretic [CLD^{+}18, EP01]. Theoretical [KJ88, CGS^{+}22, DPO^{+}15, Egl06, HG10, PGPR16]. theories [OKM^{+}05]. Theory [Ano86c, Ano91, BL86, BW87, Swa88, BP03, GH99, Lew92]. therapeutic [PDR^{+}21, WBSD12]. therapeutical [PGBMG^{+}15, PGPR16]. therapies [MGDPC^{+}15]. Therapy [BC00, DG09, HVPC24, HF20, LJR^{+}19, MK20, MD18, Zha22]. Thermal [BBR84, THH92, TRMB09]. thermal-optic [TRMB09]. thermographs [RMJLSG10]. Thermoregulation [ST85]. thick [RL13]. thickening [FS17, MMF24]. thickness [SNB^{+}11]. Thin [OKM^{+}05]. Thin-film [OKM^{+}05]. third [RW01]. thoraco [JLPB08]. Thorns [AKR^{+}89]. Three [FKCK17, GL00, Tho88b, CO19, HCWK17, KT93, LvdB98]. three-compartment [HCWK17]. three-dimensional [CO19]. Three-factor [FKCK17]. Three-locus [Tho88b]. three-nation [KT93]. Three-stage [GL00]. Threshold [Gre88, Mod97, BMN^{+}93, dSCM06, KKO^{+}11, RNE^{+}20]. threshold-type [KKO^{+}11]. Thresholds [LRW93, BWM97, MHMDM17, PG10]. thrombin [BGG03]. thrombosis [FG04]. thrombus [LLZ^{+}17]. throughout [MSZMGFS18, MS13]. thyroid [PMB14]. thyroiditis [PMB14]. tick [Mwa02]. tick-borne [Mwa02]. Ticks [Mwa02]. Tiller [CD86]. tilt [WWWW^{+}14]. Time [BS87a, PM86, CC08, DBG14, DMS25, ETOK03, FKCK17, HMMN97, KG99, LL00, MSB10, PDD08, Ree97, RPSGR^{+}07, RMK04, Sec16, SGBCC18, TRK^{+}15, Yam18, YB18]. time-dependencies [FKCK17]. time-dependent [Sec16]. time-evolution [DBG14]. Time-homogeneous [PM86]. time-inhomogeneous [PDD08]. time-since-last-fire-sample [Ree97]. times [ML10]. timescale [BKL09].

Tissue [CS91, CW07, WCSR06, BV20, BHO96, CGP06, ÇÇ17, EBFR21, HSV24, KKC^{+}03, KBVW19, LLZ^{+}17, MMMW15, MGSS21b, MGSS21a, OWB10, ONE^{+}15, PSWO14, PRK05, SW12, SHGB17, TSE24, TPL09, TRMB09, WGCB07, WG20, YLS^{+}20]. tissue-engineered [TPL09]. tissues [SGK16]. tone [Kro10]. tool [DM25]. Tools [O’N13]. top [dSCD20]. total [CB15]. toxic [CSeA02]. Toxicity [BGW99, BG11, CBB94, MS01, MS05]. toxin [GM19]. trabeculation [BLLM18]. tracer [ACG^{+}15]. tracing [PJL^{+}22, PK23, RB15, dAL02]. trade [BD20]. trade-offs [BD20]. Trait [Kim84]. Trajectories [BvGPB24]. transcoding [CBR97]. Transcription [LTY15, BK19]. transdermal [LKH98, PLFP17]. transfer [CBR97, CLD^{+}18, TRMB09]. Transform [DS91, DM25]. Transient [SKTO93, ST85, BWW05]. Transition [THI91, HZ25]. translation [ABMR21]. Transmission [AMMJ86, BA88a, BA88b, BA88c, Sch84, BMN^{+}93, CCL^{+}02, GA94, HWG11, JvdBMH98, MDS^{+}24, Mwa02, NS14, PMAV23, PJL^{+}22, RMLC14, Ric09, Sey95, TRK^{+}15, YC99, YJY21]. transmitted [ED00, LBY00, Mod97, Yan02, dAL02]. Transport [AKR^{+}89, SSSK91, AP22, ASY16, CW07, GC15, HMM23, IN12, KK15, KK17, MMMW15, ONE^{+}15, Sec16, SMGW13, SW12, SN98, SGZ19, VJ94, WG03]. Transposable [Oht84]. transverse [SPL^{+}07]. Traps [WP85]. trastuzumab [JBG^{+}19]. trastuzumab-induced [JBG^{+}19]. Travelling [SKG11, BTCE18, EHMS22, JOBF23, KWW^{+}13, PMN00]. travelling-wave [KWW^{+}13]. treat [Don23]. treated [BASH20, RNE^{+}20]. treatment [BC00, CLM^{+}20, EVK^{+}18, GDL01, HPR^{+}23, Hsi96a, Hsi96b, HC00, LMMS10, LP08, MBC^{+}14, NPO14, SGRPSM^{+}09, VHBCC96, YC99, YB18]. treatments [BW01, GM19, TW00]. tree [WG03]. Trees [Poz88]. Trials [WL89]. Tribulations [WL89]. Trichostrongylus [DGDW87]. triggered [LLZ^{+}17]. tritrophic [dSCdA16]. Triumphs [WL89]. Truncated [Pan89, Ren98]. trypanosomiasis [Bak92]. tube [SJ03]. tuberculosis [Rob92, Zha22, ARCE^{+}07]. tubular [RL13]. tumor [CGG^{+}22, HYM^{+}24, MR20]. Tumorous [THI91]. Tumour [CS90, CS91, GJ90, LP01, AALL23, And05, BMB18, BCSS15, BCSS18, BCSS21, BFPH17, BP03, CS93, CGP06, DP16, EHH12, FK03, FFK17, GDFH13, GC08, HBS^{+}12, JOBF23, LCC^{+}17, ML10, MGDPC^{+}15, MCK04, Muñ16, OC96, OC97, PR16, PS03, SGRPSM^{+}09, WK97, WK99, WA20]. tumour-immune [FFK17]. tumour-related [OC96]. tumour-secreted [EHH12]. Tumours [CS90, AP22, BM02, MS01, MS05, OS98, SN98]. tuna [KT93]. Tune [GJ90]. Tune-to-Tumour [GJ90]. Turing [GL15]. Two [Lou89, OC97, ST85, Tho88b, ASY16, BTCE18, BKL09, CKS20, DCRK21, DF18, GGA06, GP11, HZ25, HKV07, Hsi96a, Hsi96b, LdA99, Mod97, Mur97, OKM^{+}05, PMAV23, Tho10, WHT12, WBG22, XZT11, ZLC03]. two-compartment [GGA06]. Two-Dimensional [ST85, OC97]. Two-locus [Tho88b]. two-patch [ASY16, HZ25]. two-phase [DF18, OKM^{+}05]. two-population [BTCE18]. two-prey [ZLC03]. two-sex [Hsi96a, Hsi96b]. two-signed [HKV07]. two-strain [PMAV23]. two-timescale [BKL09].

Two-type [Lou89, LdA99, Mod97]. two-ventricle [DCRK21]. type [AJ12, HCWK17, KKO^{+}11, Lou89, LdA99, Mod97, NO94, NSCO97, NPO14, Ven93, ZEM11, SPN22]. Types [DWLS87]. tyrosinaemia [WDD^{+}17].

UK [HWK97]. ulcer [ZOZ^{+}22]. ultrasound [THH92]. ultrastructure [DGS08]. Uncertainty [Lud89, MPE87, DOM^{+}09, PK21]. uncoupled [SDA^{+}23]. Uncovering [HR16]. underlying [RQA12]. underpinning [Whi20]. understandability [BCC08]. Understanding [O’N13, TPS^{+}08, WP85]. unidimensional [GM19]. Unified [DM89, BMY04]. uniform [FK03, GC08]. uniformly [FK03]. Unifying [Phe91b]. uninformed [HF20]. Uniqueness [MM87]. unit [BJ14]. units [Tsi24]. University [Ano91, O’N13]. Unmyelinated [GS84a, GS84b]. unreported [DMNW20]. unsteady [WS09]. ups [JC19]. upstream [ABMR21]. Uptake [SSS91]. Urban [BA91]. uremic [GM19]. ureter [SWWC09]. urinary [SEL17]. USA [PMAV23]. usage [HCWK17]. Use [Pan89, GDL01, KGG^{+}99, Lew92, RT01]. users [CGH12, GY93, GH97]. Using [AC89, AA00, ACC18, ALDLDMC12, BCS97, BSA^{+}12, BP03, CHC13, GR98, GL00, HYM^{+}24, LLZ^{+}17, MK20, NBGWJ18, RTG21, SSS91, SNHMM16, SGZ19, TC04, TRMB09, WBSD12, YF00, ZNB97].

Vaccination [Cai89, Sch84, BMN^{+}93, BMd19, ED00, JCS15, KG99, MMA^{+}21, NS95, PMAV23, PGC^{+}00, RMLC14, WCM98, Yan02, dMS08]. vaccine [AA00, CKS20]. Valencian [AAVAV24]. Validating [HMMN97]. Valsalva [PTW^{+}17]. vanishing [EHMS22]. Variability [Agu87, BCC08, Hor84a, ML00]. Variable [BM01, BA88c, Cai90, MK89]. variables [Tho10]. Variance [GT91, Hor84a, PM86]. variants [HBHJ22]. Variation [BH90, LHZ87]. variations [GD98]. Varying [BMM93, Hsi96a, Hsi96b, MH97, Mon07]. vascular [CGS^{+}22, ÇÇ17]. vascularised [AP22]. vascularized [SN98]. vector [AG01]. vegetation [LPSR24]. VEGF [SQGP12]. VEGF-associated [SQGP12]. velocity [BM01]. venous [ZOZ^{+}22]. Ventilation [BS87b, CHMH98, CKL^{+}05]. ventilator [CHMH98]. ventricle [DCRK21, MS13, TH96]. versus [BZF^{+}15, FKCK17]. vertex [NBGWJ18]. vertex-based [NBGWJ18]. vertical [MANS11]. vertically [ED00]. Very [Sey90]. vessel [LLZ^{+}17, RK01]. vessels [CB15]. vestibulo [BMB^{+}23]. vestibulo-cerebellar [BMB^{+}23]. via [BFK17, GMM^{+}99, MANS11]. View [DM89]. vigilance [BR98, PB00]. villages [NS14]. viral [CH20, NBC18, NS95, PFF^{+}20, TW00, ZJJA19]. Virtual [MM87, EVK^{+}18, GCC^{+}12]. virulence [Kor09a, Sey95]. Virus [AMMJ86, BA88a, CCL^{+}02, HBHJ22, HA12, JvdBMH98, JvdBD02, Kor09a, DP08, LBY00, Tam99, VHT07, ZZ08]. viscoelastic [Kro10]. vitamin [FDC^{+}24]. vitro [ONE^{+}15, HMM23, LG24]. vivo [JBG^{+}19, SBM04, UPM^{+}15]. volatile [DLM23]. Voltage [Poz88]. Volterra [LC04, PMN00, SK24]. Volume [Ano00b, Ano01b, Ano04, Seg03, SGZ19, Ano84b, Ano85b, Ano86b, Ano96b]. vs [CB15, DLM23].

walk [PS03, RG95]. Wall [KJ88, FG04, SNB^{+}11, VJ94]. walls [PMN00]. water [BGW99, EBFR21]. Wave [LMM87, BTCE18, DPO^{+}15, EHMS22, HSV24, KWW^{+}13]. waveform [RK01]. Waveforms [BS87b]. Waves [AKR^{+}89, LMM87, BHB19, BHB20, JOBF23, PMN00, SKG11]. weak [Krz01, SKG11]. webs [CK93]. Weinberg [YB03]. well [Yam18]. well-posedness [Yam18]. Western [ALM17]. wetting [WAB10]. Whales [Hib85]. Wheat [CD86]. Which [Don23, BKL09, NM92]. white [LJR^{+}19, PP96, PK23]. white-noise [PP96]. whole [IN12]. wild [HCWK17]. wild-type [HCWK17]. Wilkinson [Bur06]. wind [BCS96]. wind-up [BCS96]. Winter [CD86]. Within [ARCE^{+}07, CPB10, GBBT19, IC04, O’N99, YB18]. Within-host [ARCE^{+}07, GBBT19]. without [Mon07, PR16, Yam18]. Women [BA91]. wound [DSMF00, GMM^{+}99, OSMA97, SMML92, ZOZ^{+}22]. wound-healing [OSMA97].

X [DS91, Tho85]. X-Linked [Tho85]. X-Ray [DS91]. xvi [bMO04].

yeast [CMR^{+}04]. Yield [BS90, dSCD20]. young [YJY21]. Yule [Pan89].

zebrafish [DSW^{+}21]. Zero [Pan89, TC03, TC04]. zero-loop [TC03]. Zero-Truncated [Pan89]. zeros [RW01]. ZIP [MSB10]. Zones [THI91]. zoonotic [ZJJA19]. Zostera [SAEHM03].

References

Altmann:2000:EVC

[AA00] Doris Altmann and Klaus Altmann. Estimating vaccine coverage by using computer algebra. Mathematical Medicine and Biology, 17(2):137–146, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

              academic.oup.com/imammb/article/17/2/137/736862.

Almeida:2023:DCM

[AALL23] Luís Almeida, Chloe Audebert, Emma Leschiera, and Tommaso Lorenzi. Discrete and continuum models for the coevolutionary dynamics between CD8+ cytotoxic T lymphocytes and tumour cells. Mathematical Medicine and Biology, 40(2): 141–174, June 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/40/2/141/6873764.

Adenane:2024:CCS

[AAVAV24] Rim Adenane, Carlos Andreu-Vilarroig, Florin Avram, and Rafael-Jacinto Villanueva. Calibration and comparison of SIR, SEIR/SLIR and SLAIR models for influenza dynamics: insights from the 2016–2017 season in the Valencian Community, Spain. Mathematical Medicine and Biology, 41(4):277–303, December 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/41/4/277/7759225.

Andreev:2021:DMN

[ABMR21] D. E. Andreev, P. V. Baranov, A. Milogorodskii, and D. Rachinskii. A deterministic model for non-monotone relationship between translation of upstream and downstream open reading frames. Mathematical Medicine and Biology, 38 (4):490–515, December 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/38/4/490/6413973.

Anthony:1989:AAI

[AC89] Helen M. Anthony and M. G. Cox. An automatic algorithm for immunoassay curve calibration using controlled quadratic splines. Mathematical Medicine and Biology, 6(2):91–110, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/6/2/91/651333.

Alvarez:2024:PDC

[AC24] Frank Ernesto Alvarez and Jean Clairambault. Phenotype divergence and cooperation in isogenic multicellularity and in cancer. Mathematical Medicine and Biology, 41(2):135–155, June 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/41/2/135/7708721.

Aniort:2018:MMC

[ACC18] Julien Aniort, Laurent Chupin, and Nicolae Cîndea. Mathematical model of calcium exchange during haemodialysis using a citrate containing dialysate. Mathematical Medicine and Biology, 35(S1):??, April 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL https://academic.

oup.com/imammb/article/35/Supplement_1/i87/4558631.

Arnold:2015:ATD

[ACG^{+}15] Andrea Arnold, Daniela Calvetti, Albert Gjedde, Peter Iversen, and Erkki Somersalo. Astrocytic tracer dynamics estimated from [1-^{11} c]-acetate PET measurements. Mathematical Medicine and Biology, 32(4):367–382, December 1, 2015. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
32/4/367/2874294.

Anderson:2021:TFD

[ACH^{+}21] Daniel M. Anderson, Maria Corsaro, Jonathan Horton, Tim Reid, and Padmanabhan Seshaiyer. Tear film dynamics with blinking and contact lens motion. Mathematical Medicine and Biology, 38(3):355–395, September 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/38/3/355/6323822.

Anderson:2008:SCB

[ACRF08] A. R. A. Anderson, M. A. J. Chaplain, K. A. Rejniak, and J. A. Fozard. Single-cell-based models in biology and medicine. Mathematical Medicine and Biology, 25(2):185–186, June 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/2/185/751933.

Anita:2014:ISC

[ACS14] Sebastian Anita, Jérôme Casas, and Christelle Suppo. Impulsive spatial control of invading pests by generalist predators. Mathematical Medicine and Biology, 31(3):284–301, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/31/3/284/710040.

Almansour:2024:SEI

[ADCN24] Suliman Almansour, Joanne L. Dunster, Jonathan J. Crofts, and Martin R. Nelson. A systematic evaluation of the influence of macrophage phenotype descriptions on inflammatory dynamics. Mathematical Medicine and Biology, 41(2):81–109, June 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/41/2/81/7644306.

Arino:2005:MSE

[ADH^{+}05] Julien Arino, Jonathan R. Davis, David Hartley, Richard Jordan, Joy M. Miller, and P. van den Driessche. A multi-species epidemic model with spatial dynamics. Mathematical Medicine and Biology, 22(2):129–142, June 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/22/2/129/770960.

Alt:1985:CGD

[AES85] Wolfgang Alt, Theodor Eisele, and Renate Schaaf. Chemotaxis of gametes: a diffusion approximation. Mathematical Medicine and Biology, 2(2):109–129, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/2/2/109/765046.

Antonelli:1987:SLA

[AFK87] P. L. Antonelli, K. D. Fuller, and N. D. Kazarinoff. A study of large-amplitude periodic solutions in a model for starfish predation of coral. Mathematical Medicine and Biology, 4(3): 207–214, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/4/3/207/874258.

Artzrouni:2001:MGS

[AG01] Marc Artzrouni and Jean-Paul Gouteux. A model of Gambian sleeping sickness with open vector populations. Mathematical Medicine and Biology, 18(2):99–117, June 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/18/2/99/
672229.

Aittokallio:2007:PER

[AGPV07] T. Aittokallio, M. Gyllenberg, O. Polo, and A. Virkki. Parameter estimation of a respiratory control model from noninvasive carbon dioxide measurements during sleep. Mathematical Medicine and Biology, 24(2):225–249, June 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/2/225/685959.

Agur:1987:RVP

[Agu87] Zvia Agur. Resilience and variability in pathogens and hosts. Mathematical Medicine and Biology, 4(4):295–307, ???? 1987.

CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
4/4/295/751257.

Alarcon:2012:IMT

[AJ12] Tomás Alarcón and Henrik Jeldtoft Jensen. Invasion in multitype populations: the role of phenotypic robustness and fluctuations. Mathematical Medicine and Biology, 29(1):3–20, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/29/1/3/753888.

Antonelli:1989:DRT

[AKR^{+}89] P. L. Antonelli, N. D. Kazarinoff, R. E. Reichelt, R. H. Bradbury, and P. J. Moran. A diffusion–reaction–transport model for large-scale waves in crown-of-thorns starfish outbreaks on the Great Barrier Reef. Mathematical Medicine and Biology, 6(2):81–89, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/6/2/81/651321.

Ayala:2012:SHO

[ALDLDMC12] Guillermo Ayala, María Concepción López-Díaz, Miguel López-Díaz, and Lucía Martínez-Costa. Studying hypertension in ocular fundus images using Hausdorff dispersion ordering. Mathematical Medicine and Biology, 29(2):131–143, June 2012. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/2/131/662199.

Artzrouni:2017:SSM

[ALM17] Marc Artzrouni, Vasiliy N. Leonenko, and Thierry A. Mara. A syringe-sharing model for the spread of HIV: application to Omsk, Western Siberia. Mathematical Medicine and Biology, 34(1):15–37, March 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/34/1/15/2885284.

Alrikaby:2019:SHB

[Alr19] Zenab Alrikaby. Stability and Hopf bifurcation analysis of lac Operon model with distributed delay and nonlinear degradation rate. Mathematical Medicine and Biology, 36(4):489– 512, December 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/36/4/489/5261128.

Anderson:1984:STG

[AM84] Roy M. Anderson and Robert M. May. Spatial, temporal, and genetic heterogeneity in host populations and the design of immunization programmes. Mathematical Medicine and Biology, 1(3):233–266, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/1/3/233/846432.

Anderson:1986:PST

[AMMJ86] R. M. Anderson, G. F. Medley, R. M. May, and A. M. Johnson. A preliminary study of the transmission dynamics of the human immunodeficiency virus (HIV), the causative agent of AIDS. Mathematical Medicine and Biology, 3(4):229–263, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/3/4/229/688610.

Anguelov:2023:QAI

[AMP^{+}23] Roumen Anguelov, G. Manjunath, Avulundiah E. Phiri, Trevor T. Nyakudya, Priyesh Bipath, June C. Serem, and Yvette N. Hlophe. Quantifying assays: inhibition of signalling pathways of cancer. Mathematical Medicine and Biology, 40 (3):266–290, September 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/40/3/266/7260841.

Anderson:2005:HMM

[And05] Alexander R. A. Anderson. A hybrid mathematical model of solid tumour invasion: the importance of cell adhesion. Mathematical Medicine and Biology, 22(2):163–186, June 2005. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
22/2/163/770979.

Anonymous:1984:A

[Ano84a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 1(4):407, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/1/4/407/768105.

Anonymous:1984:IV

[Ano84b] Anonymous. INDEX TO VOLUME 1. Mathematical Medicine and Biology, 1(4):409–412, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/1/4/409/768109.

Anonymous:1985:A

[Ano85a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 2(4):311, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/2/4/311/833998.

Anonymous:1985:IV

[Ano85b] Anonymous. INDEX TO VOLUME 2. Mathematical Medicine and Biology, 2(4):313–316, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/2/4/313/834000.

Anonymous:1986:A

[Ano86a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 3(4):341, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/3/4/341/688638.

Anonymous:1986:IV

[Ano86b] Anonymous. INDEX TO VOLUME 3. Mathematical Medicine and Biology, 3(4):343, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/3/4/343/688643.

Anonymous:1986:RTG

[Ano86c] Anonymous. Renewal theory, Geiger counters, and the maximum number of receptors bound to a randomly haptenated polymer chain. Mathematical Medicine and Biology, 3(2): 71–97, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/2/71/752720.

Anonymous:1989:E

[Ano89] Anonymous. Erratum. Mathematical Medicine and Biology, 6(1):1, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/6/1/1/666275.

Anonymous:1991:SIC

[Ano91] Anonymous. The Sixth IMA Conference on The Mathematical Theory of the Dynamics of Biological Systems 1–3 July, 1992, University of Oxford. Mathematical Medicine and Biology, 8(3):221, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/8/3/221/689712.

Anonymous:1993:AI

[Ano93a] Anonymous. Author index. Mathematical Medicine and Biology, 10(4):303–304, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/10/4/303/678673.

Anonymous:1993:SI

[Ano93b] Anonymous. Subject index. Mathematical Medicine and Biology, 10(4):305–308, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/10/4/305/678699.

Anonymous:1996:A

[Ano96a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 13(4):309, December 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/13/4/309/771433.

Anonymous:1996:IV

[Ano96b] Anonymous. INDEX TO VOLUME 13. Mathematical Medicine and Biology, 13(4):311–315, December 1996. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/4/311/771439.

Anonymous:1997:A

[Ano97] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 14(4):325, December 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/14/4/325/655365.

Anonymous:1998:A

[Ano98a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 15(4):401, December 1998. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/15/4/401/678316.

Anonymous:1998:AI

[Ano98b] Anonymous. Author index. Mathematical Medicine and Biology, 15(4):403–404, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/15/4/403/678318.

Anonymous:1999:A

[Ano99a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 16(4):407, December 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/16/4/407/652983.

Anonymous:1999:AI

[Ano99b] Anonymous. Author index. Mathematical Medicine and Biology, 16(4):409–410, December 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/16/4/409/652984.

Anonymous:2000:A

[Ano00a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 17(4):415, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/17/4/415/664767.

Anonymous:2000:IV

[Ano00b] Anonymous. Index to volume 17. Mathematical Medicine and Biology, 17(4):417–418, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/17/4/417/664774.

Anonymous:2000:SI

[Ano00c] Anonymous. Subject index. Mathematical Medicine and Biology, 17(4):419–422, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/17/4/419/664782.

Anonymous:2001:A

[Ano01a] Anonymous. Acknowledgements. Mathematical Medicine and Biology, 18(4):409, December 2001. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/18/4/409/747277.

Anonymous:2001:IV

[Ano01b] Anonymous. Index to volume 18. Mathematical Medicine and Biology, 18(4):411–412, December 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/18/4/411/747280.

Anonymous:2001:SI

[Ano01c] Anonymous. Subject index. Mathematical Medicine and Biology, 18(4):413–415, December 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/18/4/413/747283.

Anonymous:2004:IV

[Ano04] Anonymous. Index to volume 21. Mathematical Medicine and Biology, 21(4):397–398, December 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/21/4/397/741367.

Anonymous:2012:CPb

[Ano12a] Anonymous. Contents page. Mathematical Medicine and Biology, 29(3):??, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/29/3/NP/798057.

Anonymous:2012:CPd

[Ano12b] Anonymous. Contents page. Mathematical Medicine and Biology, 29(4):??, December 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/29/4/NP/821102.

Anonymous:2012:C

[Ano12c] Anonymous. Cover. Mathematical Medicine and Biology, 29 (1):??, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/29/1/NP/753913.

Anonymous:2012:CPa

[Ano12d] Anonymous. Cover page. Mathematical Medicine and Biology, 29(3):??, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/29/3/NP/798039.

Anonymous:2012:CPc

[Ano12e] Anonymous. Cover page. Mathematical Medicine and Biology, 29(4):??, December 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/29/4/NP/821036.

Anonymous:2012:EBa

[Ano12f] Anonymous. Editorial Board. Mathematical Medicine and Biology, 29(1):??, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/29/1/NP/753909.

Anonymous:2012:EBb

[Ano12g] Anonymous. Editorial Board. Mathematical Medicine and Biology, 29(3):??, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/29/3/NP/798029.

Anonymous:2012:EBc

[Ano12h] Anonymous. Editorial Board. Mathematical Medicine and Biology, 29(4):??, December 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/29/4/NP/821014.

Anonymous:2012:IAa

[Ano12i] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 29(1):??, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/29/1/NP/753959.

Anonymous:2012:IAb

[Ano12j] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 29(3):??, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/29/3/NP/798051.

Anonymous:2012:IAc

[Ano12k] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 29(4):??, December 2012. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/29/4/NP/821081.

Anonymous:2012:TC

[Ano12l] Anonymous. Table of contents. Mathematical Medicine and Biology, 29(1):??, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/29/1/NP/753961.

Anonymous:2013:CPb

[Ano13a] Anonymous. Contents page. Mathematical Medicine and Biology, 30(1):??, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/30/1/NP/710657.

Anonymous:2013:CPd

[Ano13b] Anonymous. Contents page. Mathematical Medicine and Biology, 30(2):??, June 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/2/NP/890440.

Anonymous:2013:CPf

[Ano13c] Anonymous. Contents page. Mathematical Medicine and Biology, 30(3):??, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/30/3/NP/684896.

Anonymous:2013:CPh

[Ano13d] Anonymous. Contents page. Mathematical Medicine and Biology, 30(4):??, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/30/4/NP/718831.

Anonymous:2013:CPa

[Ano13e] Anonymous. Cover page. Mathematical Medicine and Biology, 30(1):??, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/1/NP/710617.

Anonymous:2013:CPc

[Ano13f] Anonymous. Cover page. Mathematical Medicine and Biology, 30(2):??, June 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/2/NP/890429.

Anonymous:2013:CPe

[Ano13g] Anonymous. Cover page. Mathematical Medicine and Biology, 30(3):??, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/30/3/NP/684890.

Anonymous:2013:CPg

[Ano13h] Anonymous. Cover page. Mathematical Medicine and Biology, 30(4):??, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/4/NP/718716.

Anonymous:2013:EBa

[Ano13i] Anonymous. Editorial Board. Mathematical Medicine and Biology, 30(1):??, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/30/1/NP/710612.

Anonymous:2013:EBb

[Ano13j] Anonymous. Editorial Board. Mathematical Medicine and Biology, 30(2):??, June 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/2/NP/890427.

Anonymous:2013:EBc

[Ano13k] Anonymous. Editorial Board. Mathematical Medicine and Biology, 30(3):??, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/30/3/NP/684881.

Anonymous:2013:EBd

[Ano13l] Anonymous. Editorial Board. Mathematical Medicine and Biology, 30(4):??, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/30/4/NP/718708.

Anonymous:2013:IAa

[Ano13m] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 30(1):??, March 2013. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/30/1/NP/710651.

Anonymous:2013:IAb

[Ano13n] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 30(2):??, June 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/30/2/NP/890434.

Anonymous:2013:IAc

[Ano13o] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 30(3):??, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/30/3/NP/684894.

Anonymous:2013:IAd

[Ano13p] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 30(4):??, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/30/4/NP/718767.

Anonymous:2014:CPb

[Ano14a] Anonymous. Contents page. Mathematical Medicine and Biology, 31(1):??, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/31/1/NP/860114.

Anonymous:2014:CPd

[Ano14b] Anonymous. Contents page. Mathematical Medicine and Biology, 31(3):??, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/31/3/NP/710069.

Anonymous:2014:CPf

[Ano14c] Anonymous. Contents page. Mathematical Medicine and Biology, 31(4):??, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/31/4/NP/646941.

Anonymous:2014:CPa

[Ano14d] Anonymous. Cover page. Mathematical Medicine and Biology, 31(1):??, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/31/1/NP/860045.

Anonymous:2014:CPc

[Ano14e] Anonymous. Cover page. Mathematical Medicine and Biology, 31(3):??, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/31/3/NP/710046.

Anonymous:2014:CPe

[Ano14f] Anonymous. Cover page. Mathematical Medicine and Biology, 31(4):??, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/31/4/NP/646910.

Anonymous:2014:EBa

[Ano14g] Anonymous. Editorial Board. Mathematical Medicine and Biology, 31(1):??, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/31/1/NP/860100.

Anonymous:2014:EBb

[Ano14h] Anonymous. Editorial Board. Mathematical Medicine and Biology, 31(3):??, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/31/3/NP/710055.

Anonymous:2014:EBc

[Ano14i] Anonymous. Editorial Board. Mathematical Medicine and Biology, 31(4):??, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/31/4/NP/646934.

Anonymous:2014:IAa

[Ano14j] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 31(1):??, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/31/1/NP/860107.

Anonymous:2014:IAb

[Ano14k] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 31(3):??, September 2014. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/3/NP/710065.

Anonymous:2014:IAc

[Ano14l] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 31(4):??, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/4/NP/646936.

Anonymous:2015:CPb

[Ano15a] Anonymous. Contents page. Mathematical Medicine and Biology, 32(1):??, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/32/1/NP/661617.

Anonymous:2015:CPd

[Ano15b] Anonymous. Contents page. Mathematical Medicine and Biology, 32(2):??, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/2/NP/830202.

Anonymous:2015:CPf

[Ano15c] Anonymous. Contents page. Mathematical Medicine and Biology, 32(3):??, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/32/3/NP/646345.

Anonymous:2015:CPa

[Ano15d] Anonymous. Cover page. Mathematical Medicine and Biology, 32(1):??, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/1/NP/661589.

Anonymous:2015:CPc

[Ano15e] Anonymous. Cover page. Mathematical Medicine and Biology, 32(2):??, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/2/NP/830148.

Anonymous:2015:CPe

[Ano15f] Anonymous. Cover page. Mathematical Medicine and Biology, 32(3):??, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/32/3/NP/646324.

Anonymous:2015:EBa

[Ano15g] Anonymous. Editorial Board. Mathematical Medicine and Biology, 32(1):??, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/32/1/NP/661612.

Anonymous:2015:EBb

[Ano15h] Anonymous. Editorial Board. Mathematical Medicine and Biology, 32(2):??, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/2/NP/830171.

Anonymous:2015:EBc

[Ano15i] Anonymous. Editorial Board. Mathematical Medicine and Biology, 32(3):??, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/32/3/NP/646341.

Anonymous:2015:IAa

[Ano15j] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 32(1):??, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/32/1/NP/661614.

Anonymous:2015:IAb

[Ano15k] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 32(2):??, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/32/2/NP/830188.

Anonymous:2015:IAc

[Ano15l] Anonymous. Instructions to authors. Mathematical Medicine and Biology, 32(3):??, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/32/3/NP/646343.

AlSariri:2022:MSM

[AP22] Tahani Al Sariri and Raimondo Penta. Multi-scale modelling of nanoparticle delivery and heat transport in vascularised tumours. Mathematical Medicine and Biology, 39(4): 332–367, December 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/39/4/332/6647792.

Alavez-Ramirez:2007:WHP

[ARCE^{+}07] Justino Alavez-Ramírez, J. Rogelio Avendano Castellanos, Lourdes Esteva, José Antonio Flores, José Luis Fuentes-Allen, Gisela García-Ramos, Guillermo Gómez, and Jesús López-Estrada. Within-host population dynamics of antibioticresistant M. tuberculosis. Mathematical Medicine and Biology, 24(1):35–56, March 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/24/1/35/692202.

Aron:1984:SND

[AS84] Joan L. Aron and Ira B. Schwartz. Some new directions for research in epidemic models. Mathematical Medicine and Biology, 1(3):267–276, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/1/3/267/846441.

Antonelli:1986:DOD

[AS86] P. L. Antonelli and J. M. Skowronski. Differential offensivedefensive games between plants and herbivores. Mathematical Medicine and Biology, 3(4):319–340, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/3/4/319/
688634.

Antonelli:1988:MMB

[AS88] P. L. Antonelli and R. M. Seymour. A model of myxomatosis based on hormonal control of rabbit-flea reproduction. Mathematical Medicine and Biology, 5(1):65–80 (??), ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

Arciero:2012:SOM

[AS12] J. C. Arciero and T. W. Secomb. Spontaneous oscillations in a model for active control of microvessel diameters. Mathematical Medicine and Biology, 29(2):163–180, June 2012. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/2/163/662207.

Arino:2016:RTP

[ASY16] Julien Arino, Chengjun Sun, and Wei Yang. Revisiting a twopatch SIS model with infection during transport. Mathematical Medicine and Biology, 33(1):29–55, March 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/33/1/29/
2363506.

Blythe:1988:DII

[BA88a] S. P. Blythe and R. M. Anderson. Distributed incubation and infectious periods in models of the transmission dynamics of the human immunodeficiency virus (HIV). Mathematical Medicine and Biology, 5(1):1–19, ???? 1988. CODEN MM-BABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL

http://academic.oup.com/imammb/article/5/1/1/763964.

Blythe:1988:HSA

[BA88b] S. P. Blythe and R. M. Anderson. Heterogeneous sexual activity models of HIV transmission in male homosexual populations. Mathematical Medicine and Biology, 5(4):237–260, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/5/4/237/700637.

Blythe:1988:VIH

[BA88c] S. P. Blythe and R. M. Anderson. Variable infectiousness in HFV transmission models. Mathematical Medicine and Biology, 5(3):181–200, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/5/3/181/692258.

Boily:1991:SCP

[BA91] Marie-Claude Boily and Roy M. Anderson. Sexual contact patterns between men and women and the spread of HIV-1 in urban centres in Africa. Mathematical Medicine and Biology, 8 (4):221–247, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/8/4/221/666747.

Boushehrian:2017:RBS

[BAJ^{+}17] Hamidreza Hajiani Boushehrian, Omid Abouali, Khosrow Jafarpur, Alireza Ghaffarieh, and Goodarz Ahmadi. Relationship between saccadic eye movements and formation of the Krukenberg’s spindle — a CFD study. Mathematical Medicine and Biology, 34(3):293–312, September 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/34/3/293/2885297.

Baker:1992:MTP

[Bak92] R. D. Baker. Modelling trypanosomiasis prevalence and periodic epidemics and epizootics. Mathematical Medicine and Biology, 9(4):269–287, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/9/4/269/730280.

Banerjee:2011:SPF

[Ban11] Malay Banerjee. Spatial pattern formation in ratio-dependent model: higher-order stability analysis. Mathematical Medicine and Biology, 28(2):111–128, June 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/28/2/111/679773.

Bolton:2020:DEM

[BASH20] Larisse Bolton, Thomas M. Acho, David K. Stones, and Cang Hui. Describing the evolution of myeloid-derived leucocytes in treated B-lineage paediatric acute lymphoblastic leukaemia with a data-driven granulocyte-monocyte-blast model. Mathematical Medicine and Biology, 37(4):433–468, December 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/4/433/5821094.

Budich:2021:SEG

[BB21] Jan C. Budich and Emil J. Bergholtz. Synchronization in epidemic growth and the impossibility of selective containment. Mathematical Medicine and Biology, 38(4):467–473, December 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/38/4/467/6409757.

Bujurke:1990:ECS

[BBN90] N. M. Bujurke, S. G. Bhavi, and N. B. Naduvinamani. The effect of couple stresses in squeeze film poro-elastic bearings with special reference to synovial joints. Mathematical Medicine and Biology, 7(4):231–243, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/7/4/231/2268157.

Bottega:2013:MDR

[BBPF13] William J. Bottega, Peter L. Bishay, Jonathan L. Prenner, and Howard F. Fine. On the mechanics of a detaching retina. Mathematical Medicine and Biology, 30(4):287–310, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
30/4/287/718676.

Brttton:1984:MMT

[BBR84] N. F. Brttton, J. R. Barker, and E. F. J. Ring. A mathematical model for a thermal clearance probe. Mathematical Medicine and Biology, 1(1):95–105, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/1/1/95/754953.

Boldrini:2000:TBD

[BC00] José L. Boldrini and Michel I. S. Costa. Therapy burden, drug resistance, and optimal treatment regimen for cancer chemotherapy. Mathematical Medicine and Biology, 17(1): 33–51, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/17/1/33/674986.

Bressloff:2006:E

[BC06] Paul C. Bressloff and Stephen Coombes. Editorial. Mathematical Medicine and Biology, 23(2):77–78, June 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/23/2/77/
671415.

Busenberg:1991:GSP

[BCC91] Stavros Busenberg and Carlos Castillo-Chavez. A general solution of the problem of mixing of subpopulations and its application to risk- and age-structured epidemic models for the spread of AIDS. Mathematical Medicine and Biology, 8(1):1–29, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/8/1/1/788776.

Bairagi:2008:VSC

[BCC08] N. Bairagi, Samrat Chatterjee, and J. Chattopadhyay. Variability in the secretion of corticotropin-releasing hormone, adrenocorticotropic hormone and cortisol and understandability of the hypothalamic-pituitary-adrenal axis dynamics — a mathematical study based on clinical evidence. Mathematical Medicine and Biology, 25(1):37–63, March 2008. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/25/1/37/745864.

Byrne:1998:ECC

[BCM98] H. M. Byrne, G. Cave, and D. L. S. Mcelwain. The effect of chemotaxis and chemokinesis on leukocyte locomotion: a new interpretation of experimental results. Mathematical Medicine and Biology, 15(3):235–256, September 1998. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
15/3/235/795644.

Bruni:2019:ONS

[BCPS19] C. Bruni, F. Conte, F. Papa, and C. Sinisgalli. Optimal number and sizes of the doses in fractionated radiotherapy according to the LQ model. Mathematical Medicine and Biology, 36 (1):1–53, March 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/36/1/1/4807290.

Britton:1996:RMD

[BCS96] N. F. Britton, M. A. J. Chaplain, and Suzanne M. Skevington. The role of N-methyl-D-aspartate (NMDA) receptors in wind-up: a mathematical model. Mathematical Medicine and Biology, 13(3):193–205, September 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/13/3/193/651313.

Bansal:1997:EAG

[BCS97] A. Bansal, C. Cannings, and N. Sheehan. An evaluation of the application of the genetic algorithm to the problem of ordering genetic loci on human chromosomes using radiation hybrid data. Mathematical Medicine and Biology, 14(3):161– 187, September 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/14/3/161/716270.

Bolton:2015:PFO

[BCSS15] Larisse Bolton, Alain H. J. J. Cloot, Schalk W. Schoombie, and Jacobus P. Slabbert. A proposed fractional-order Gompertz model and its application to tumour growth data. Mathematical Medicine and Biology, 32(2):187–209, June 2015. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/

32/2/187/830133. See corrigenda [BCSS18, BCSS21].

Bolton:2018:CPF

[BCSS18] Larisse Bolton, Alain H. J. J. Cloot, Schalk W. Schoombie, and Jacobus P. Slabbert. Corrigendum to: A proposed fractionalorder Gompertz model and its application to tumour growth data. Mathematical Medicine and Biology, 35(2):273–277, June 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/35/2/273/4969411. See [BCSS15, BCSS21].

Bolton:2021:CPF

[BCSS21] Larisse Bolton, Alain H. J. J. Cloot, Schalk W. Schoombie, and Jacobus P. Slabbert. Corrigendum: A proposed fractional order Gompertz model, and its application to tumour growth data. Mathematical Medicine and Biology, 38(1):132–135, March 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/38/1/132/5883547. See [BCSS15, BCSS18].

Bearon:2020:MSB

[BD20] R. N. Bearon and W. M. Durham. A model of strongly biased chemotaxis reveals the trade-offs of different bacterial migration strategies. Mathematical Medicine and Biology, 37(1): 83–116, March 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/37/1/83/5426812.

Braun:2018:TFB

[BDB^{+}18] Richard J. Braun, Tobin A. Driscoll, Carolyn G. Begley, P. Ewen King-Smith, and Javed I. Siddique. On tear film breakup (TBU): dynamics and imaging. Mathematical Medicine and Biology, 35(2):145–180, June 2018. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/2/145/3038033.

Baker:2013:MMC

[BDJB^{+}13] Michelle Baker, Sarah Denman-Johnson, Bindi S. Brook, Ian Gaywood, and Markus R. Owen. Mathematical modelling of cytokine-mediated inflammation in rheumatoid arthritis. Mathematical Medicine and Biology, 30(4):311–337, December 2013.

CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
30/4/311/718682.

Bell:1984:BSM

[Bel84] Jonathan Bell. Behaviour of some models of myelinated axons. Mathematical Medicine and Biology, 1(2):149–167, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/1/2/149/643664.

Bell:1986:PDC

[Bel86] Jonathan Bell. Parametric dependence of conduction speed for a diffusive model of a myelinated axon. Mathematical Medicine and Biology, 3(4):289–300, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/3/4/289/688622.

Braun:2003:MDP

[BF03] R. J. Braun and A. D. Fitt. Modelling drainage of the precorneal tear film after a blink. Mathematical Medicine and Biology, 20(1):1–28, March 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/20/1/1/663514.

Barnard:2017:SAD

[BFK17] Richard C. Barnard, Martin Frank, and Kai Krycki. Sensitivity analysis for dose deposition in radiotherapy via a Fokker– Planck model. Mathematical Medicine and Biology, 34(1): 109–123, March 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/34/1/109/2885288.

Bertsch:2017:ADM

[BFM^{+}17] Michiel Bertsch, Bruno Franchi, Norina Marcello, Maria Carla Tesi, and Andrea Tosin. Alzheimer’s disease: a mathematical model for onset and progression. Mathematical Medicine and Biology, 34(2):193–214, June 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/34/2/193/2885293.

Borsi:2017:NLM

[BFPH17] I. Borsi, A. Fasano, M. Primicerio, and T. Hillen. A non-local model for cancer stem cells and the tumour growth paradox.

Mathematical Medicine and Biology, 34(1):59–75, March 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
34/1/59/2885286.

Boston:2011:ITC

[BG11] Eleanor A. J. Boston and Eamonn A. Gaffney. The influence of toxicity constraints in models of chemotherapeutic protocol escalation. Mathematical Medicine and Biology, 28(4):357– 384, December 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/28/4/357/658332.

Bungay:2003:MML

[BGG03] Sharene D. Bungay, Patricia A. Gentry, and Rodney D. Gentry. A mathematical model of lipid-mediated thrombin generation. Mathematical Medicine and Biology, 20(1):105–129, March 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/20/1/105/663529.

Berger:1988:HTF

[BGW88] Agnes Berger, Guadalupe Gómez, and Sylvan Wallenstein. A homogeneity test for follow-up studies. Mathematical Medicine and Biology, 5(2):101–112, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/5/2/101/731014.

Belov:1999:TWC

[BGW99] A. P. Belov, J. D. Giles, and R. J. Wiltshire. Toxicity in a water column following the stratification of a cyanobacterial population development in a calm lake. Mathematical Medicine and Biology, 16(1):93–110, March 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/16/1/93/729549.

Bell:1986:BMS

[BH86] Jonathan Bell and Mark H. Holmes. Behaviour of a model for the synapse of an auditory receptor cell. Mathematical Medicine and Biology, 3(4):301–317, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/3/4/301/688629.

Brush:1990:CGV

[BH90] G. Brush and G. A. Harrison. Components of growth variation in human stature. Mathematical Medicine and Biology, 7 (2):77–92, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/7/2/77/698802.

Bradshaw-Hajek:2019:ASC

[BHB19] Bronwyn H. Bradshaw-Hajek and Philip Broadbridge. Analytic solutions for calcium ion fertilisation waves on the surface of eggs. Mathematical Medicine and Biology, 36(4): 549–562, December 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/36/4/549/5320215. See erratum

[BHB20].

Bradshaw-Hajek:2020:ASC

[BHB20] Bronwyn H. Bradshaw-Hajek and Philip Broadbridge. Analytic solutions for calcium ion fertilisation waves on the surface of eggs — erratum. Mathematical Medicine and Biology, 37(3): 429–432, September 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/37/3/429/5809230. See [BHB19].

Bos:1996:ESR

[BHO96] Cees Bos, Louis Hoofd, and Thom Oostendorp. The effect of separate red blood cells on capillary tissue oxygenation calculated with a numerical model. Mathematical Medicine and Biology, 13(4):259–274, December 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/13/4/259/771418.

Brook:2014:RCU

[BJ14] B. S. Brook and O. E. Jensen. The role of contractile unit reorganization in force generation in airway smooth muscle. Mathematical Medicine and Biology, 31(2):99–124, June 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
31/2/99/660578.

Bokes:2019:LCO

[BK19] Pavol Bokes and John R. King. Limit-cycle oscillatory coexpression of cross-inhibitory transcription factors: a model mechanism for lineage promiscuity. Mathematical Medicine and Biology, 36(1):113–137, March 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/36/1/113/4904094.

Bannish:2014:MFS

[BKF14] Brittany E. Bannish, James P. Keener, and Aaron L. Fogelson. Modelling fibrinolysis: a 3D stochastic multiscale model. Mathematical Medicine and Biology, 31(1):17–44, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
31/1/17/860024.

Bokes:2009:BGS

[BKL09] Pavol Bokes, John R. King, and Matthew Loose. A bistable genetic switch which does not require high co-operativity at the promoter: a two-timescale model for the PU.1-GATA-1 interaction. Mathematical Medicine and Biology, 26(2):117–132, June 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/26/2/117/727670.

Painter:2004:BRD

[bKP04] Reviewed by Kevin Painter. Book review: Differential Equations and Mathematical Biology by D. S. Jones and B. D. Sleeman CRC Press, ISBN 1-58488-296-4. Mathematical Medicine and Biology, 21(2):167–168, June 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/21/2/167/698813.

Bannish:2014:MFC

[BKW^{+}14] Brittany E. Bannish, James P. Keener, Michael Woodbury, John W. Weisel, and Aaron L. Fogelson. Modelling fibrinolysis: 1D continuum models. Mathematical Medicine and Biology, 31 (1):45–64, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/31/1/45/860031.

Battelli:1986:SPT

[BL86] Flaviano Battelli and Claudio Lazzari. Singular perturbation theory for open enzyme reaction networks. Mathematical Medicine and Biology, 3(1):41–51, ???? 1986. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
3/1/41/684216.

Brinks:2015:NRB

[BL15] Ralph Brinks and Sandra Landwehr. A new relation between prevalence and incidence of a chronic disease. Mathematical Medicine and Biology, 32(4):425–435, December 1, 2015. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
32/4/425/2874297.

Blackwell:1990:DSM

[Bla90] Paul Blackwell. Deterministic and stochastic models of social behaviour based on the resource dispersion hypothesis. Mathematical Medicine and Biology, 7(4):261–279, ???? 1990. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
7/4/261/744282.

Battista:2018:FDH

[BLLM18] Nicholas A. Battista, Andrea N. Lane, Jiandong Liu, and Laura A. Miller. Fluid dynamics in heart development: effects of hematocrit and trabeculation. Mathematical Medicine and Biology, 35(4):493–516, December 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/4/493/4637493.

Belair:2001:VMV

[BM01] Jacques Bélair and Joseph M. Mahaffy. Variable maturation velocity and parameter sensitivity in a model of haematopoiesis. Mathematical Medicine and Biology, 18(2):193–211, June 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
18/2/193/672239.

Byrne:2002:AGM

[BM02] Helen Byrne and Paul Matthews. Asymmetric growth of models of avascular solid tumours: exploiting symmetries. Mathematical Medicine and Biology, 19(1):1–29, March 2002. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
19/1/1/708233.

Bobadilla:2018:SMT

[BMB18] Ana Victoria Ponce Bobadilla, Philip K. Maini, and Helen Byrne. A stochastic model for tumour control probability that accounts for repair from sublethal damage. Mathematical Medicine and Biology, 35(2):181–202, June 2018. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/2/181/3055078.

Burlando:2023:MDS

[BMB^{+}23] Bruno Burlando, Viviana Mucci, Cherylea J. Browne, Serena Losacco, Iole Indovina, Lucio Marinelli, Franco Blanchini, and Giulia Giordano. Mal de Debarquement Syndrome explained by a vestibulo-cerebellar oscillator. Mathematical Medicine and Biology, 40(1):96–110, March 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/40/1/96/6873556.

Buonomo:2019:OPH

[BMd19] Bruno Buonomo, Rossella Della Marca, and Alberto d’Onofrio. Optimal public health intervention in a behavioural vaccination model: the interplay between seasonality, behaviour and latency period. Mathematical Medicine and Biology, 36(3):297– 324, September 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/36/3/297/5057753.

Bunimovich-Mendrazitsky:2016:IBC

[BMHK16] Svetlana Bunimovich-Mendrazitsky, Sarel Halachmi, and Natalie Kronik. Improving bacillus Calmette–Guérin (BCG) immunotherapy for bladder cancer by adding interleukin 2 (IL-2): a mathematical model. Mathematical Medicine and Biology, 33(2):159–188, June 1, 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/33/2/159/2875344.

Burke:1993:SSR

[BMM93] Meghan A. Burke, P. K. Maini, and J. D. Murray. Suicide substrate reaction–diffusion equations: Varying the source. Mathematical Medicine and Biology, 10(2):97–114, ???? 1993. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
10/2/97/654162.

BezerraCoutinho:1993:EVP

[BMN^{+}93] Françisco Antonio Bezerra Coutinho, Eduardo Massad, Marcelo Nascimento Burattini, Hyur Mo Yang, and Raymundo Soares De Azevedo Neto. Effects of vaccination programmes on transmission rates of infections and related threshold conditions for control. Mathematical Medicine and Biology, 10(3): 187–206, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/3/187/666912.

Owen:2004:BRM

[bMO04] Reviewed by Markus Owen. Book review, Modelling Neural Development, Edited by Arjen van Ooyen, xvi + 336 Pages, 96 Illustrations, The MIT Press, Cambridge, Massachusetts, 2003, ISBN 0-262-22066-0. Mathematical Medicine and Biology, 21 (1):73–74, March 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/21/1/73/689405.

Babak:2004:DGF

[BMS04] Petro Babak, Kjartan G. Magnússon, and Sven Sigurdsson. Dynamics of group formation in collective motion of organisms. Mathematical Medicine and Biology, 21(4):269–292, December 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/21/4/269/741351.

Ball:2000:SMS

[BMY00] Frank G. Ball, Robin K. Milne, and Geoffrey F. Yeo. Stochastic models for systems of interacting ion channels. Mathematical Medicine and Biology, 17(3):263–293, ???? 2000. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/3/263/667738.

Ball:2004:UAB

[BMY04] Frank G. Ball, Robin K. Milne, and Geoffrey F. Yeo. A unified approach to burst properties of multiconductance single ion channels. Mathematical Medicine and Biology, 21(3):205– 245, September 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/21/3/205/660057.

Britton:2005:RMM

[bNFB05] Reviewed by N. F. Britton. Review: Mathematical Models in Biology: an Introduction. Mathematical Medicine and Biology, 22(3):289–290, September 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/22/3/289/652728.

Bokler:1993:CCM

[Bok93] Benjamin Bokler. Chaos and complexity in measles models: a comparative numerical study. Mathematical Medicine and Biology, 10(2):83–95, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/10/2/83/654157.

Ball:1998:SCG

[BP98] M. A. Ball and G. A. Parker. Sperm-competition games: Energy dependence and competitor numbers in the continuousexternal-fertilization model. Mathematical Medicine and Biology, 15(1):87–96, March 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/15/1/87/683947.

Byrne:2003:MST

[BP03] Helen Byrne and Luigi Preziosi. Modelling solid tumour growth using the theory of mixtures. Mathematical Medicine and Biology, 20(4):341–366, December 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/20/4/341/714032.

Broom:1998:MRV

[BR98] M. Broom and G. D. Ruxton. Modelling responses in vigilance rates to arrivals to and departures from a group of foragers. Mathematical Medicine and Biology, 15(4):387–400, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/15/4/387/678312.

Brandts:1993:FML

[Bra93] Wendy A. M. Brandts. A field model of left-right asymmetries in the pattern regulation of a cell. Mathematical Medicine and Biology, 10(1):31–50, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/10/1/31/749594.

Baranyi:1993:SPN

[BRM93] József Baranyi, Terry A. Roberts, and Peter Mcclure. Some properties of a nonautonomous deterministic growth model describing the adjustment of the bacterial population to a new environment. Mathematical Medicine and Biology, 10(4):293–299, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/10/4/293/678668.

Brookfield:1989:ADF

[Bro89] J. F. Y. Brookfield. Analysis of DNA fingerprinting data in cases of disputed paternity. Mathematical Medicine and Biology, 6(2):111–131, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/6/2/111/651309.

Berestycki:2025:BIE

[BRR25] Henri Berestycki, Jean-Michel Roquejoffre, and Luca Rossi. Biological invasions and epidemics with nonlocal diffusion along a line. Mathematical Medicine and Biology, 42(1):4–37, March 2025. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/42/1/4/7727261.

Ball:1987:TCI

[BS87a] Frank Ball and Mark Sansom. Temporal clustering of ion channel openings incorporating time interval omission. Mathematical Medicine and Biology, 4(4):333–361, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/4/4/333/
751269.

Boyarsky:1987:OPW

[BS87b] Abraham Boyarsky and John Senez. Optimal pressure waveforms for pressure-limited ventilation. Mathematical Medicine and Biology, 4(3):201–205, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/4/3/201/874252.

Blythe:1990:SCS

[BS90] S. P. Blythe and T. K. Stokes. Some consequences of sizeselective harvesting on fitness and on yield. Mathematical Medicine and Biology, 7(1):41–53, ???? 1990. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
7/1/41/715240.

Berglund:2012:ICM

[BSA^{+}12] Martin Berglund, Mikael Sunnåker, Martin Adiels, Mats Jirstrand, and Bernt Wennberg. Investigations of a compartmental model for leucine kinetics using non-linear mixed effects models with ordinary and stochastic differential equations. Mathematical Medicine and Biology, 29(4):361–384, December 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/29/4/361/821007.

Britton:2007:BSL

[BSEO07] Tom Britton, Bodil Svennblad, Per Erixon, and Bengt Oxelman. Bayesian support is larger than bootstrap support in phylogenetic inference: a mathematical argument. Mathematical Medicine and Biology, 24(4):401–411, December 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/4/401/661380.

Britton:2015:IGN

[BT15] Tom Britton and Pieter Trapman. Inferring global network properties from egocentric data with applications to epidemics. Mathematical Medicine and Biology, 32(1):101–114, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/32/1/101/661588.

Broadbridge:2021:DDI

[BT21] P. Broadbridge and B. S. Tilley. Diffusion of dermatological irritant in drying laundered cloth. Mathematical Medicine and Biology, 38(4):474–489, December 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/38/4/474/6410104.

Bitsouni:2018:ATW

[BTCE18] Vasiliki Bitsouni, Dumitru Trucu, Mark A. J. Chaplain, and Raluca Eftimie. Aggregation and travelling wave dynamics in a two-population model of cancer cell growth and invasion. Mathematical Medicine and Biology, 35(4):541–577, December 2018.

CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/4/541/4808586.

Burger:1986:EDF

[Bur86] Reinhard Burger. Evolutionary dynamics of functionally constrained phenotypic characters. Mathematical Medicine and Biology, 3(4):265–287, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/3/4/265/688616.

Burrage:2006:SMS

[Bur06] Kevin Burrage. Book review: Stochastic Modelling for Systems Biology by Darren J. Wilkinson. Mathematical Medicine and Biology, 23(4):391–392, December 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/23/4/391/689804.

Banerjee:2020:SIR

[BV20] M. Banerjee and V. Volpert. Stochastic intracellular regulation can remove oscillations in a model of tissue growth. Mathematical Medicine and Biology, 37(4):551–568, December 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/4/551/5879256.

Bogaardt:2024:MIB

[BvGPB24] Laurens Bogaardt, Anoukh van Giessen, H. Susan J. Picavet, and Hendriek C. Boshuizen. A model of individual BMI trajectories. Mathematical Medicine and Biology, 41(1):1–18, March 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/41/1/1/7505278.

Britton:1987:ODT

[BW87] N. F. Britton and J. Waniewski. One-dimensional theory of haemofilters. Mathematical Medicine and Biology, 4(1):59–68, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/4/1/59/875673.

Beil:2001:ACM

[BW01] Damian R. Beil and Lawrence M. Wein. Analysis and comparison of multimodal cancer treatments. Mathematical Medicine and Biology, 18(4):343–376, December 2001. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
18/4/343/747272.

Basse:1997:PTS

[BWM97] B. Basse, G. C. Wake, and J. A. Mclennan. Predation thresholds for survival of endangered species. Mathematical Medicine and Biology, 14(3):241–250, September 1997. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
14/3/241/716280.

Berezovskaya:2011:DPC

[BWSCC11] Faina Berezovskaya, S. Wirkus, B. Song, and C. Castillo-Chavez. Dynamics of population communities with prey migrations and Allee effects: a bifurcation approach. Mathematical Medicine and Biology, 28(2):129–152, June 2011. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
28/2/129/679772.

Begg:2005:FEM

[BWW05] Ronald Begg, David J. N. Wall, and Graeme C. Wake. On a functional equation model of transient cell growth. Mathematical Medicine and Biology, 22(4):371–390, December 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
22/4/371/650320.

Basse:2004:CGM

[BWWvB04] B. Basse, G. C. Wake, D. J. N. Wall, and B. van Brunt. On a cell-growth model for plankton. Mathematical Medicine and Biology, 21(1):49–61, March 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/21/1/49/689389.

Byrne:1997:IIA

[Byr97] H. M. Byrne. The importance of intercellular adhesion in the development of carcinomas. Mathematical Medicine and Biology, 14(4):305–323, December 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/14/4/305/655361.

Blllard:1993:SGE

[BZ93] L. Blllard and Zhen Zhao. The stochastic general epidemic model revisited and a generalization. Mathematical Medicine and Biology, 10(1):67–75, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/10/1/67/749601.

Bertacchi:2015:CVS

[BZF^{+}15] Daniela Bertacchi, Fabio Zucca, Sergio Foresti, Davide Mangioni, and Andrea Gori. Combination versus sequential monotherapy in chronic HBV infection: a mathematical approach. Mathematical Medicine and Biology, 32(4):383–403, December 1, 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/4/383/2874295.

Cairns:1989:EHP

[Cai89] Andrew J. G. Cairns. Epidemics in heterogeneous populations: Aspects of optimal vaccination policies. Mathematical Medicine and Biology, 6(3):137–159, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/6/3/137/684124.

Cairns:1990:EHP

[Cai90] Andrew J. G. Cairns. Epidemics in heterogeneous populations: II. Nonexponential incubation periods and variable infectiousness. Mathematical Medicine and Biology, 7(4):219–230, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/7/4/219/744268.

Carletti:2006:NSC

[Car06] Margherita Carletti. Numerical simulation of a Campbell-like stochastic delay model for bacteriophage infection. Mathematical Medicine and Biology, 23(4):297–310, December 2006. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/4/297/689793.

Cooke:1984:RBC

[CB84] J. G. Cooke and J. R. Beddington. The relationship between catch rates and abundance in fisheries. Mathematical Medicine and Biology, 1(4):391–405, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/1/4/391/768096.

Costanzo:2015:ILL

[CB15] Francesco Costanzo and James G. Brasseur. The invalidity of the Laplace law for biological vessels and of estimating elastic modulus from total stress vs. strain: a new practical method. Mathematical Medicine and Biology, 32(1):1–37, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
32/1/1/661585.

Costa:1992:OCC

[CBB92] M. I. S. Costa, J. L. Boldrini, and R. C. Bassanezi. Optimal chemical control of populations developing drug resistance. Mathematical Medicine and Biology, 9(3):215–226, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
9/3/215/649559.

Costa:1994:OCC

[CBB94] M. I. S. Costa, J. L. Boldrini, and R. C. Bassanezi. Optimal chemotherapy: a case study with drug resistance, saturation effect, and toxicity. Mathematical Medicine and Biology, 11 (1):45–59, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/11/1/45/667155.

Cogan:2021:DAS

[CBPD21] Ng Cogan, Feng Bao, Ralf Paus, and Atanaska Dobreva. Data assimilation of synthetic data as a novel strategy for predicting disease progression in alopecia areata. Mathematical Medicine and Biology, 38(3):314–332, September 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/3/314/6292138.

Chapeau-Blondeau:1997:STN

[CBR97] François Chapeau-Blondeau and Frédéric Raguin. Signal transcoding by nonlinear sensory neurons: Information-entropy maximization, optimal transfer function, and anti-Hebbian adaptation. Mathematical Medicine and Biology, 14(3):227– 239, September 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/14/3/227/716278.

Chadha-Boreham:1989:DSM

[CBW89] Harbajan K. Chadha-Boreham and Norman Westwood. Drawbacks of selection methods for synchronous cell growth: Simulation techniques. Mathematical Medicine and Biology, 6(4): 243–255, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/6/4/243/794019.

Chadha-Boreham:1991:DIM

[CBWA91] Harbajan K. Chadha-Boreham, Norman Westwood, and Malcolm Andrew. Drawbacks of induction methods for synchronous cell growth: Simulation techniques. Mathematical Medicine and Biology, 8(4):273–286, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/8/4/273/666759.

Cammarota:2008:AEH

[CC08] Camillo Cammarota and Mario Curione. Analysis of extrema of heartbeat time series in exercise test. Mathematical Medicine and Biology, 25(1):87–97, March 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/25/1/87/745873.

Coban:2017:NCR

[ÇÇ17] Gürsan Çoban and M. Serdar Çelebi. A novel computational remodelling algorithm for the probabilistic evolution of collagen fibre dispersion in biaxially strained vascular tissue. Mathematical Medicine and Biology, 34(4):433–467, December 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
34/4/433/2885302.

Choi:2002:MTD

[CCL^{+}02] Y. H. Choi, C. Comiskey, M. D. A. Lindsay, J. A. Cross, and M. Anderson. Modelling the transmission dynamics of Ross River virus in Southwestern Australia. Mathematical Medicine and Biology, 19(1):61–74, March 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/19/1/61/708243.

Camp:1994:NGA

[CCS94] N. Camp, C. Cannings, and N. Sheehan. The number of genotypic assignments on a genealogy I. The method and simple examples. Mathematical Medicine and Biology, 11(2):95–106, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/11/2/95/771450.

Camp:1999:NGA

[CCS99] N. J. Camp, C. Cannings, and N. A. Sheehan. The number of genotypic assignments on a genealogy II. Further results for linear systems. Mathematical Medicine and Biology, 16(3):213– 236, September 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/3/213/657142.

Cubarsi:2005:MAM

[CCV05] R. Cubarsi, M. M. Carrió, and A. Villaverde. A mathematical approach to molecular organization and proteolytic disintegration of bacterial inclusion bodies. Mathematical Medicine and Biology, 22(3):209–226, September 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/22/3/209/652725.

Chaudhuri:2013:EOS

[CCV13] S. Chaudhuri, A. Costamagna, and E. Venturino. Ecoepidemics overcoming the species barrier and being subject to harvesting. Mathematical Medicine and Biology, 30(1):73–93, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
30/1/73/710604.

Cubarsi:1998:NTM

[CCVV98] R. Cubarsi, J. L. Corchero, P. Vila, and A. Vilaverde. Numerical techniques and mathematical modelling for CI857controlled gene expression and cell growth in recombinant E. coli. Mathematical Medicine and Biology, 15(3):257–278, September 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/15/3/257/795653.

Chalabi:1986:ADJ

[CD86] Z. S. Chalabi and W. Day. Application of dynamic jump process analysis to modelling tiller production in winter wheat. Mathematical Medicine and Biology, 3(1):23–40, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
3/1/23/684213.

Clarelli:2016:FDM

[CDNR16] F. Clarelli, C. Di Russo, R. Natalini, and M. Ribot. A fluid dynamics multidimensional model of biofilm growth: stability, influence of environment and sensitivity. Mathematical Medicine and Biology, 33(4):371–395, December 2016. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
33/4/371/2674333.

Coleman:1984:MME

[CF84] K. D. Coleman and A. C. Fowler. A mathematical model of exoprotein production in bacteria. Mathematical Medicine and Biology, 1(1):77–94, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/1/1/77/754947.

Canning:2002:FFA

[CGDF02] C. R. Canning, M. J. Greaney, J. N. Dewynne, and A. D. Fitt. Fluid flow in the anterior chamber of a human eye. Mathematical Medicine and Biology, 19(1):31–60, March 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/19/1/31/
708239.

Cherfils:2022:TGM

[CGG^{+}22] Laurence Cherfils, Stefania Gatti, Carole Guillevin, Alain Miranville, and Rémy Guillevin. On a tumor growth model with brain lactate kinetics. Mathematical Medicine and Biology, 39 (4):382–409, December 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/39/4/382/6661797.

Corson:2012:MMS

[CGH12] S. Corson, D. Greenhalgh, and S. Hutchinson. Mathematically modelling the spread of hepatitis C in injecting drug users.

Mathematical Medicine and Biology, 29(3):205–230, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/29/3/205/798023.

Chaplain:2006:MML

[CGP06] M. A. J. Chaplain, L. Graziano, and L. Preziosi. Mathematical modelling of the loss of tissue compression responsiveness and its role in solid tumour development. Mathematical Medicine and Biology, 23(3):197–229, September 2006. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/3/197/672277.

Chiaravalli:2022:MSM

[CGS^{+}22] Greta Chiaravalli, Giovanna Guidoboni, Riccardo Sacco, Jake Radell, and Alon Harris. A multi-scale/multi-physics model for the theoretical study of the vascular configuration of retinal capillary plexuses based on OCTA data. Mathematical Medicine and Biology, 39(1):77–104, March 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/39/1/77/
6439246.

Campos:1987:AIG

[CH87] Tania M. M. Campos and Philip Holgate. Algebraic isotopy in genetics. Mathematical Medicine and Biology, 4(3):215–222, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/4/3/215/874264.

Cassidy:2020:MMV

[CH20] Tyler Cassidy and Antony R. Humphries. A mathematical model of viral oncology as an immuno-oncology instigator. Mathematical Medicine and Biology, 37(1):117–151, March 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/37/1/117/5480296.

Chen:2013:MCC

[CHC13] Chih-Yuan Chen, Kuan-Chih Hsiau, and C. A. Chung. Measurement of chondrocyte chemotaxis using a Boyden chamber: a model of receptor-mediated cell migration combined with cell sedimentation. Mathematical Medicine and Biology, 30(3):213– 239, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/3/213/684873.

Crooke:1998:PVI

[CHMH98] P. S. Crooke, J. D. Head, J. J. Marini, and J. R. Hotchkiss. Patient-ventilator interaction: a general model for nonpassive mechanical ventilation. Mathematical Medicine and Biology, 15 (4):321–337, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/15/4/321/678303.

Cannings:1993:ESD

[CHV93] C. Cannings, V. Hutson, and G. T. Vickers. The evolution of spatial diffusion. Mathematical Medicine and Biology, 10(3): 169–177, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/3/169/666909.

Cheer:1987:FFT

[CK87] A. Y. L. Cheer and M. A. R. Koehl. Fluid flow through filtering appendages of insects. Mathematical Medicine and Biology, 4 (3):185–199, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/4/3/185/874240.

Colombo:1993:SSF

[CK93] Rinaldo M. Colombo and Vlastimil Krivan. Selective strategies in food webs. Mathematical Medicine and Biology, 10(4): 281–291, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/4/281/678661.

Cogan:2004:RBM

[CK04] N. G. Cogan and James P. Keener. The role of the biofilm matrix in structural development. Mathematical Medicine and Biology, 21(2):147–166, June 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/21/2/147/698810.

Chatterjee:2007:RHI

[CKC07] Samrat Chatterjee, Kusumika Kundu, and J. Chattopadhyay. Role of horizontal incidence in the occurrence and control of chaos in an eco-epidemiological system. Mathematical Medicine and Biology, 24(3):301–326, September 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/3/301/712094.

Crooke:2005:MMP

[CKL^{+}05] P. S. Crooke, K. Kongkul, Y. Lenbury, A. B. Adams, C. S. Carter, J. J. Marini, and J. R. Hotchkiss. Mathematical models for pressure controlled ventilation of oleic acid-injured pigs. Mathematical Medicine and Biology, 22(1):99–112, March 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
22/1/99/657687.

Chernov:2020:OVA

[CKS20] Alexey A. Chernov, Mark Y. Kelbert, and Aleksandr A. Shemendyuk. Optimal vaccine allocation during the mumps outbreak in two SIR centres. Mathematical Medicine and Biology, 37(3):303–312, September 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/37/3/303/5526742.

Chen:2018:PTT

[CLD^{+}18] Shi Chen, Suzanne Lenhart, Judy D. Day, Chihoon Lee, Michael Dulin, and Cristina Lanzas. Pathogen transfer through environment-host contact: an agent-based queueing theoretic framework. Mathematical Medicine and Biology, 35(3):409– 425, September 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/35/3/409/4585724.

Camacho:2020:OCM

[CLM^{+}20] Erika T. Camacho, Suzanne Lenhart, Luis A. Melara, M. Cristina Villalobos, and Stephen Wirkus. Optimal control with MANF treatment of photoreceptor degeneration. Mathematical Medicine and Biology, 37(1):1–21, March 2020. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/1/1/5363832.

Cruywagen:1992:SPF

[CMM92] Gerhard C. Cruywagen, P. K. Maini, and J. D. Murray. Sequential pattern formation in a model for skin morphogenesis.

Mathematical Medicine and Biology, 9(4):227–248, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
9/4/227/730250.

Chen:2017:MBO

[CMM^{+}17] Qiyong Chen, Sinan Müftü, Faik Can Meral, Kemal Tuncali, and Murat Akçakaya. Model-based optimal planning of hepatic radiofrequency ablation. Mathematical Medicine and Biology, 34(3):415–431, September 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/34/3/415/2885301.

Cole:2004:ENP

[CMR^{+}04] D. J. Cole, B. J. T. Morgan, M. S. Ridout, L. J. Byrne, and M. F. Tuite. Estimating the number of prions in yeast cells. Mathematical Medicine and Biology, 21(4):369–395, December 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/21/4/369/741364.

Carichino:2019:ETD

[CO19] Lucia Carichino and Sarah D. Olson. Emergent threedimensional sperm motility: coupling calcium dynamics and preferred curvature in a Kirchhoff rod model. Mathematical Medicine and Biology, 36(4):439–469, December 2019. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
36/4/439/5132952.

Cox:2006:AMC

[Cox06] Steven J. Cox. An adjoint method for channel localization. Mathematical Medicine and Biology, 23(2):139–152, June 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/2/139/671410.

Chauviere:2010:MCM

[CPB10] A. Chauviere, L. Preziosi, and H. Byrne. A model of cell migration within the extracellular matrix based on a phenotypic switching mechanism. Mathematical Medicine and Biology, 27 (3):255–281, September 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/27/3/255/791818.

Chaplain:1990:MMP

[CS90] M. A. J. Chaplain and B. D. Sleeman. A mathematical model for the production and secration of tumour angiogenesis factor in tumours. Mathematical Medicine and Biology, 7(2):93–108, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/7/2/93/698808.

Chaplain:1991:MMD

[CS91] M. A. J. Chaplain and A. M. Stuart. A mathematical model for the diffusion of tumour angiogenesis factor into the surrounding host tissue. Mathematical Medicine and Biology, 8(3):191–220, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/8/3/191/689707.

Chaplain:1993:MMC

[CS93] M. A. J. Chaplain and A. M. Stuart. A model mechanism for the chemotactic response of endothelial cells to tumour angiogenesis factor. Mathematical Medicine and Biology, 10(3): 149–168, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/3/149/666908.

Cain:2008:SCA

[CS08] John W. Cain and David G. Schaeffer. Shortening of cardiac action potential duration near an insulating boundary. Mathematical Medicine and Biology, 25(1):21–36, March 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/25/1/21/
745860.

Chattopadhyay:2002:DDE

[CSeA02] J. Chattopadhyay, R. R. Sarkar, and A. el Abdllaoui. A delay differential equation model on harmful algal blooms in the presence of toxic substances. Mathematical Medicine and Biology, 19(2):137–161, June 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/19/2/137/730723.

Choudhury:2020:MMF

[CTB20] M. Jabed A. Choudhury, Philip M. J. Trevelyan, and Graeme P. Boswell. Mathematical modelling of fungi-initiated siderophore-iron interactions. Mathematical Medicine and Biology, 37(4):515–550, December 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/37/4/515/5870745.

Cuzick:1984:RSP

[Cuz84] Jack Cuzick. A review of semi-parametric models for life histories. Mathematical Medicine and Biology, 1(4):323–332, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/1/4/323/768064.

Cardamone:2010:MCA

[CVEH10] L. Cardamone, A. Valentín, J. F. Eberth, and J. D. Humphrey. Modelling carotid artery adaptations to dynamic alterations in pressure and flow over the cardiac cycle. Mathematical Medicine and Biology, 27(4):343–371, December 2010. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/4/343/724982.

Cummings:2007:TGR

[CW07] L. J. Cummings and S. L. Waters. Tissue growth in a rotating bioreactor. Part II: fluid flow and nutrient transport problems. Mathematical Medicine and Biology, 24(2):169–208, June 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/2/169/685945.

Chen:2014:MMH

[CW14] C. Y. Chen and J. P. Ward. A mathematical model for the human menstrual cycle. Mathematical Medicine and Biology, 31(1):65–86, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/31/1/65/860035.

Dordevic:2022:SOC

[– DD22] Jasmina – Dordević and Kristina Rognlien Dahl. Stochastic optimal control of pre-exposure prophylaxis for HIV infection. Mathematical Medicine and Biology, 39(3):197–225, September 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/39/3/197/6590971.

deArazoza:2002:NLM

[dAL02] H. de Arazoza and R. Lounes. A non-linear model for a sexually transmitted disease with contact tracing. Mathematical Medicine and Biology, 19(3):221–234, September 2002. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
19/3/221/729874.

Dassios:2008:NCE

[Das08] George Dassios. Neuronal currents and EEG-MEG fields. Mathematical Medicine and Biology, 25(2):133–139, June 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/2/133/751926.

Devi:2015:SSP

[DAS15] Gurumayum Reenaroy Devi, Md. Jahoor Alam, and R. K. Brojen Singh. Synchronization in stress p53 network. Mathematical Medicine and Biology, 32(4):437–456, December 1, 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
32/4/437/2874298.

Dowty:2014:TED

[DBG14] James G. Dowty, Graham B. Byrnes, and Dorota M. Gertig. The time-evolution of DCIS size distributions with applications to breast cancer growth and progression. Mathematical Medicine and Biology, 31(4):353–364, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
31/4/353/646894.

DOrsi:2021:MMC

[DCC^{+}21] Laura D’Orsi, Luciano Curcio, Fabio Cibella, Alessandro Borri, Lilach Gavish, Arik Eisenkraft, and Andrea De Gaetano. A mathematical model of cardiovascular dynamics for the diagnosis and prognosis of hemorrhagic shock. Mathematical Medicine and Biology, 38(4):417–441, December 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/4/417/6367650.

Doyle:2021:LPM

[DCRK21] Matthew G. Doyle, Marina Chugunova, S. Lucy Roche, and James P. Keener. Lumped parameter models for two-ventricle and healthy and failing extracardiac Fontan circulations. Mathematical Medicine and Biology, 38(4):442–466, December 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/4/442/6371787.

Deshmukh:2006:EPM

[DCT06] Vikrant Deshmukh, Chris Cannings, and Alun Thomas. Estimating the parameters of a model for protein-protein interaction graphs. Mathematical Medicine and Biology, 23(4):279– 295, December 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/23/4/279/689797.

Demongeot:2025:E

[Dem25] Jacques Demongeot. Editorial. Mathematical Medicine and Biology, 42(1):1–3, March 2025. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/42/1/1/8078048.

Du:2018:TPM

[DF18] Jian Du and Aaron L. Fogelson. A two-phase mixture model of platelet aggregation. Mathematical Medicine and Biology, 35 (2):225–256, June 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/35/2/225/3038034.

DOnofrio:2009:FMA

[DG09] Alberto D’Onofrio and Alberto Gandolfi. A family of models of angiogenesis and anti-angiogenesis anti-cancer therapy. Mathematical Medicine and Biology, 26(1):63–95, March 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
26/1/63/853994.

Dobson:1987:GMD

[DGDW87] R. J. Dobson, D. A. Griffiths, A. D. Donald, and P. J. Waller. A genetic model describing the evolution of levamisole resistance in Trichostrongylus colubriformis, a nematode parasite of sheep. Mathematical Medicine and Biology, 4(4):279–293, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/4/4/279/751251.

Costanzo:2018:DCM

[DGM^{+}18] Ezio Di Costanzo, Alessandro Giacomello, Elisa Messina, Roberto Natalini, Giuseppe Pontrelli, Fabrizio Rossi, Robert Smits, and Monika Twarogowska. A discrete in continuous mathematical model of cardiac progenitor cells formation and growth as spheroid clusters (cardiospheres). Mathematical Medicine and Biology, 35(1):121–144, March 2018. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/1/121/2938064.

Dunster:2023:ECM

[DGN23] Joanne L. Dunster, Jonathan M. Gibbins, and Martin R. Nelson. Exploring the constituent mechanisms of hepatitis: a dynamical systems approach. Mathematical Medicine and Biology, 40(1):24–48, March 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/40/1/24/6748195.

DeRemigio:2008:CRS

[DGS08] Hilary DeRemigio, Jeffrey R. Groff, and Gregory D. Smith. Calcium release site ultrastructure and the dynamics of puffs and sparks. Mathematical Medicine and Biology, 25(1):65– 85, March 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/25/1/65/745869.

DeBoer:1987:SND

[DH87] Rob J. De Boer and Pauline Hogeweg. Self-nonself discrimination due to immunological nonlinearities: the analysis of a series of models by numerical methods. Mathematical Medicine and Biology, 4(1):1–32, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/4/1/1/875656.

Deng:2004:NAA

[DH04] Qingping Deng and Thomas G. Hallam. Numerical approximations for an age-structured model of a population dispersing in a spatially heterogeneous environment. Mathematical Medicine and Biology, 21(3):247–268, September 2004. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
21/3/247/660065.

Dallon:2006:MME

[DJG06] John Dallon, Wonhee Jang, and Richard H. Gomer. Mathematically modelling the effects of counting factor in Dictyostelium discoideum. Mathematical Medicine and Biology, 23 (1):45–62, March 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/23/1/45/809602.

DellaMarca:2023:IVV

[DLM23] Rossella Della Marca, Nadia Loy, and Marco Menale. Intransigent vs. volatile opinions in a kinetic epidemic model with imitation game dynamics. Mathematical Medicine and Biology, 40 (2):111–140, June 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/40/2/111/6881768.

Daley:1989:UVM

[DM89] D. J. Daley and J. H. Maindonald. A unified view of models describing the avoidance of superparasitism. Mathematical Medicine and Biology, 6(3):161–178, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/6/3/161/
684128.

Demongeot:2025:SPT

[DM25] Jacques Demongeot and Jean-Gabriel Minonzio. A signalprocessing tool adapted to the periodic biphasic phenomena: the Dynalet transform. Mathematical Medicine and Biology, 42(1):113–129, March 2025. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/42/1/113/7933690.

Ducrot:2020:INU

[DMNW20] A. Ducrot, P. Magal, T. Nguyen, and G. F. Webb. Identifying the number of unreported cases in SIR epidemic models. Mathematical Medicine and Biology, 37(2):243–261, June 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/2/243/5526758.

Demongeot:2025:FCB

[DMO25] Jacques Demongeot, Pierre Magal, and Kayode Oshinubi. Forecasting the changes between endemic and epidemic phases of a contagious disease, with the example of COVID-19. Mathematical Medicine and Biology, 42(1):98–112, March 2025. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/42/1/98/
7737412.

dOnofrio:2008:FSD

[dMS08] Alberto d’Onofrio, Piero Manfredi, and Ernesto Salinelli. Fatal SIR diseases and rational exemption to vaccination. Mathematical Medicine and Biology, 25(4):337–357, December 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/4/337/826449.

Ducrot:2025:LTB

[DMS25] A. Ducrot, D. Manceau, and A. Sylla. On the large time behaviour of the solutions of an evolutionary-epidemic system with spatial dispersal. Mathematical Medicine and Biology, 42 (1):38–70, March 2025. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/42/1/38/7899829.

Danforth:2009:IUB

[DOM^{+}09] Christopher M. Danforth, Thomas Orfeo, Kenneth G. Mann, Kathleen E. Brummel-Ziedins, and Stephen J. Everse. The impact of uncertainty in a blood coagulation model. Mathematical Medicine and Biology, 26(4):323–336, December 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
26/4/323/654367.

Donovan:2023:WAS

[Don23] Graham M. Donovan. Which airways should we treat? Structure-function relationships and estimation of the singular input modes from the forward model alone. Mathematical Medicine and Biology, 40(4):291–307, December 2023. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
40/4/291/7286684.

Leenheer:2008:MVD

[DP08] Patrick De Leenheer and Sergei S. Pilyugin. Multistrain virus dynamics with mutations: a global analysis. Mathematical Medicine and Biology, 25(4):285–322, December 2008. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/4/285/826470.

Dassios:2016:SST

[DP16] George Dassios and Vasiliki Christina Panagiotopoulou. On the stability of a spherical tumour. Mathematical Medicine and Biology, 33(3):273–293, September 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/33/3/273/1750089.

Dobreva:2018:ADM

[DPC18] Atanaska Dobreva, Ralf Paus, and N. G. Cogan. Analysing the dynamics of a model for alopecia areata as an autoimmune disorder of hair follicle cycling. Mathematical Medicine and Biology, 35(3):387–407, September 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/3/387/4071274.

Du:2015:TSI

[DPO^{+}15] Peng Du, Niranchan Paskaranandavadivel, Greg O’Grady, Shou-Jiang Tang, and Leo K. Cheng. A theoretical study of the initiation, maintenance and termination of gastric slow wave re-entry. Mathematical Medicine and Biology, 32(4):405– 423, December 1, 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/4/405/2874296.

Dvoriashyna:2018:AHF

[DRRT18] M. Dvoriashyna, R. Repetto, M. R. Romano, and J. H. Tweedy. Aqueous humour flow in the posterior chamber of the eye and its modifications due to pupillary block and iridotomy. Mathematical Medicine and Biology, 35(4):447–467, December 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/4/447/4582768.

Dassios:1991:NRE

[DS91] G. Dassios and B. D. Sleeman. A note on the reconstruction of ellipsoids from the X-Ray transform. Mathematical Medicine and Biology, 8(2):141–147, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/8/2/141/698759.

Costa:2020:IBF

[dSCD20] Michel Iskin da S. Costa and Lucas Dos Anjos. The interplay between fishery yield and top predator culling in a multispecies fishery context. Mathematical Medicine and Biology, 37(3):351– 363, September 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/37/3/351/5700594.

Costa:2016:AET

[dSCdA16] Michel Iskin da S. Costa and Lucas dos Anjos. Allee effects in tritrophic food chains: some insights in pest biological control. Mathematical Medicine and Biology, 33(4):461–474, December 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/33/4/461/2674335.

Costa:2006:ATP

[dSCM06] Michel Iskin da Silveira Costa and Magno Enrique Mendoza Meza. Application of a threshold policy in the management of multispecies fisheries and predator culling. Mathematical Medicine and Biology, 23(1):63–75, March 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/23/1/63/
809618.

Dallon:2000:BID

[DSMF00] John Dallon, Jonathan Sherratt, Philip Maini, and Mark Ferguson. Biological implications of a discrete mathematical model for collagen deposition and alignment in dermal wound repair. Mathematical Medicine and Biology, 17(4):379–393, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/4/379/664742.

Dirks:2021:MMB

[DSW^{+}21] Carolin Dirks, Paul Striewski, Benedikt Wirth, Anne Aalto, and Adan Olguin-Olguin. A mathematical model for bleb regulation in zebrafish primordial germ cells. Mathematical Medicine and Biology, 38(2):218–254, June 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/2/218/6144681.

Downham:1987:DDF

[DWLS87] D. Y. Downham, B. C. Wilson, J. Lexell, and M. Siöström. Distribution of different fibre types in human skeletal muscles: a method for the detection of neurogenic disorders. Mathematical Medicine and Biology, 4(1):81–91, ???? 1987. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
4/1/81/875680.

Ellingsrud:2021:ANS

[EBFR21] Ada J. Ellingsrud, Nicolas Boullé, Patrick E. Farrell, and Marie E. Rognes. Accurate numerical simulation of electrodiffusion and water movement in brain tissue. Mathematical Medicine and Biology, 38(4):516–551, December 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/4/516/6430038.

El-Doma:2000:SAG

[ED00] M. El-Doma. Stability analysis of a general age-dependent vaccination model for a vertically transmitted disease under the proportionate mixing assumption. Mathematical Medicine and Biology, 17(2):119–136, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/17/2/119/736858.

Eglen:2006:DRC

[Egl06] Stephen J. Eglen. Development of regular cellular spacing in the retina: theoretical models. Mathematical Medicine and Biology, 23(2):79–99, June 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/23/2/79/671407.

Engelhardt:2018:MMA

[EHE^{+}18] Benjamin Engelhardt, Janine Holze, Christina Elliott, George S. Baillie, Maik Kschischo, and Holger Fröhlich. Modelling and mathematical analysis of the M—_{2} — receptor-dependent joint signalling and secondary messenger network in CHO cells. Mathematical Medicine and Biology, 35(3):279–297, September 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/35/3/279/3827653.

Enderling:2012:PRT

[EHH12] Heiko Enderling, Lynn Hlatky, and Philip Hahnfeldt. The promoting role of a tumour-secreted chemorepellent in selfmetastatic tumour progression. Mathematical Medicine and Biology, 29(1):21–29, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/29/1/21/753885.

El-Hachem:2022:NVS

[EHMS22] Maud El-Hachem, Scott W. McCue, and Matthew J. Simpson. Non-vanishing sharp-fronted travelling wave solutions of the Fisher–Kolmogorov model. Mathematical Medicine and Biology, 39(3):226–250, September 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/39/3/226/6603417.

Engwer:2016:EEA

[EHS16] Christian Engwer, Alexander Hunt, and Christina Surulescu. Effective equations for anisotropic glioma spread with proliferation: a multiscale approach and comparisons with previous settings. Mathematical Medicine and Biology, 33(4):435– 459, December 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/33/4/435/2674337.

Echavarria-Heras:2012:SAP

[EHSALR12] Hector Echavarria-Heras, Elena Solana-Arellano, and Cecilia Leal-Ramirez. Surface aggregation patterns of LDL receptors near coated pits II. The retrograde membrane flow-diffusion and generalized plaque-form insertion mechanism. Mathematical Medicine and Biology, 29(2):109–130, June 2012. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/2/109/662195.

Evans:2001:CTA

[EP01] N. D. Evans and A. J. Pritchard. A control theoretic approach to containing the spread of rabies. Mathematical Medicine and Biology, 18(1):1–23, March 1, 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/18/1/1/727365.

Eisentrager:2013:PEC

[ESC13] Almut Eisenträger, Ian Sobey, and Marek Czosnyka. Parameter estimations for the cerebrospinal fluid infusion test. Mathematical Medicine and Biology, 30(2):157–174, June 2013. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
30/2/157/890413.

Eshima:2007:WDH

[ETO07] Nobuoki Eshima, Minoru Tabata, and Tadashige Okada. Why is the distribution of HTLV-I carriers geographically biased? An answer through a mathematical epidemic model. Mathematical Medicine and Biology, 24(2):149–167, June 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/2/149/685948.

Eshima:2003:PDH

[ETOK03] Nobuoki Eshima, Minoru Tabata, Tadashige Okada, and Shigeru Karukaya. Population dynamics of HTLV-I infection: a discrete-time mathematical epidemic model approach. Mathematical Medicine and Biology, 20(1):29–45, March 2003. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/1/29/663518.

Evans:2000:AME

[Eva00] J. D. Evans. Analysis of a multiple equivalent cylinder model with generalized taper. Mathematical Medicine and Biology, 17 (4):347–377, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/17/4/347/664730.

Eymard:2018:MMC

[EVK^{+}18] N. Eymard, V. Volpert, P. Kurbatova, V. Volpert, N. Bessonov, K. Ogungbenro, L. Aarons, P. Janiaud, P. Nony, A. Bajard, S. Chabaud, Y. Bertrand, B. Kassaï, C. Cornu, and P. Nony. Mathematical model of T-cell lymphoblastic lymphoma: disease, treatment, cure or relapse of a virtual cohort of patients. Mathematical Medicine and Biology, 35(1):25–47, March 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/1/25/2706309.

Faddy:1985:NSC

[Fad85] M. J. Faddy. Nonlinear stochastic compartmental models. Mathematical Medicine and Biology, 2(4):287–297, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
2/4/287/833991.

Faddy:1993:DAS

[Fad93] M. J. Faddy. Diffusion approximations for stochastic compartmental models. Mathematical Medicine and Biology, 10(3): 179–186, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/3/179/666910.

Fozard:2010:CAI

[FBJK10] J. A. Fozard, H. M. Byrne, O. E. Jensen, and J. R. King. Continuum approximations of individual-based models for epithelial monolayers. Mathematical Medicine and Biology, 27 (1):39–74, March 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/27/1/39/731502.

Franks:2003:MMC

[FBM^{+}03] S. J. Franks, H. M. Byrne, H. S. Mudhar, J. C. E. Underwood, and C. E. Lewis. Mathematical modelling of comedo ductal carcinoma in situ of the breast. Mathematical Medicine and Biology, 20(3):277–308, September 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/20/3/277/714221.

Franks:2024:MIV

[FDC^{+}24] S. J. Franks, J. L. Dunster, S. R. Carding, J. M. Lord, M. Hewison, P. C. Calder, and J. R. King. Modelling the influence of vitamin D and probiotic supplementation on the microbiome and immune response. Mathematical Medicine and Biology, 41 (4):304–345, December 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/41/4/304/7800638.

Frascoli:2017:MEC

[FFK17] Federico Frascoli, Emelie Flood, and Peter S. Kim. A model of the effects of cancer cell motility and cellular adhesion properties on tumour-immune dynamics. Mathematical Medicine and Biology, 34(2):215–240, June 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/34/2/215/2885294.

Fogelson:2004:PWI

[FG04] Aaron L. Fogelson and Robert D. Guy. Platelet-wall interactions in continuum models of platelet thrombosis: formulation and numerical solution. Mathematical Medicine and Biology, 21 (4):293–334, December 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/21/4/293/741355.

Fallon:2009:ITD

[FHC09] Marissa S. Fallon, Brett A. Howell, and Anuj Chauhan. Importance of Taylor dispersion in pharmacokinetic and multiple indicator dilution modelling. Mathematical Medicine and Biology, 26(4):263–296, December 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/26/4/263/654356.

Frei:2020:SMC

[FHR20] Christoph Frei, Thomas Hillen, and Adam Rhodes. A stochastic model for cancer metastasis: branching stochastic process with settlement. Mathematical Medicine and Biology, 37(2): 153–182, June 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/37/2/153/5510090.

Fowler:2000:RCC

[FK00] A. C. Fowler and G. P. Kalamangalam. The role of the central chemoreceptor in causing periodic breathing. Mathematical Medicine and Biology, 17(2):147–167, ???? 2000. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/2/147/736866.

Fowler:2002:PBH

[FK02] Andrew C. Fowler and G. P. Kalamangalam. Periodic breathing at high altitude. Mathematical Medicine and Biology, 19(4): 293–313, December 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/19/4/293/753209.

Franks:2003:IBU

[FK03] S. J. Franks and J. R. King. Interactions between a uniformly proliferating tumour and its surroundings: uniform material properties. Mathematical Medicine and Biology, 20(1): 47–89, March 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/20/1/47/663520.

Frank:2017:TFM

[FKCK17] Till D. Frank, Anatoly Kiyatkin, Alex Cheong, and Boris N. Kholodenko. Three-factor models versus time series models: quantifying time-dependencies of interactions between stimuli in cell biology and psychobiology for short longitudinal data. Mathematical Medicine and Biology, 34(2):177–191, June 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
34/2/177/2885291.

Flower:1993:MAG

[FKK93] A. C. Flower, G. P. Kalamangalam, and G. Kember. A mathematical analysis of the Grodins model of respiratory control. Mathematical Medicine and Biology, 10(4):249–280, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
10/4/249/678655.

Fister:2008:ICS

[FM08] K. Renee Fister and Maeve L. McCarthy. Identification of a chemotactic sensitivity in a coupled system. Mathematical Medicine and Biology, 25(3):215–232, September 2008. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/3/215/682275.

Fatehi:2010:MCA

[FMHF10] Hasnaa Fatehi, Michael Meyer-Hermann, and Marc Thilo Figge. Modelling cellular aggregation induced by chemotaxis and phototaxis. Mathematical Medicine and Biology, 27(4): 373–384, December 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/27/4/373/724984.

Ferreira:2016:CNF

[FNdO16] J. A. Ferreira, J. Naghipoor, and Paula de Oliveira. A coupled non-Fickian model of a cardiovascular drug delivery system. Mathematical Medicine and Biology, 33(3):329–357, September 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/33/3/329/1750278.

Fok:2012:GNC

[Fok12] Pak-Wing Fok. Growth of necrotic cores in atherosclerotic plaque. Mathematical Medicine and Biology, 29(4):301– 327, December 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/29/4/301/820997.

Fabregas:2014:MMP

[FR14] L. Rene I. Fabregas and J. Rubinstein. A mathematical model for the progression of dental caries. Mathematical Medicine and Biology, 31(4):319–337, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/4/319/646891.

Fok:2017:BMM

[FS17] Pak-Wing Fok and Rebecca Sanft. A biochemical and mechanical model of injury-induced intimal thickening. Mathematical Medicine and Biology, 34(1):77–108, March 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/34/1/77/
2885289.

Fassoni:2019:MDO

[FY19] Artur C. Fassoni and Hyun M. Yang. Modeling dynamics for oncogenesis encompassing mutations and genetic instability. Mathematical Medicine and Biology, 36(2):241–267, June 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/36/2/241/5043024.

Garnett:1994:BSP

[GA94] Geoffrey P. Garnett and Roy M. Anderson. Balancing sexual partnership in an age and activity stratified model of HIV transmission in heterosexual populations. Mathematical Medicine and Biology, 11(3):161–192, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/11/3/161/708086.

Gani:1984:SPE

[Gan84] J. Gani. Some population and epidemic models revisited. Mathematical Medicine and Biology, 1(3):277–287, ???? 1984. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
1/3/277/846446.

Gerrard:1986:MNO

[GB86] R. Gerrard and A. D. Barbour. Measures of niche overlap, II. Mathematical Medicine and Biology, 3(2):115–127, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
3/2/115/752656.

Go:2019:WWH

[GBBT19] Natacha Go, Catherine Belloc, Caroline Bidot, and Suzanne Touzeau. Why, when and how should exposure be considered at the within-host scale? A modelling contribution to PRRSv infection. Mathematical Medicine and Biology, 36(2):179–206, June 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/36/2/179/5000061.

Ge:2016:IEC

[GBPF16] Peinan Ge, William J. Bottega, Jonathan L. Prenner, and Howard F. Fine. On the influence of an equatorial cerclage on closure of posterior retinal detachment. Mathematical Medicine and Biology, 33(4):417–433, December 2016. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
33/4/417/2674336.

Gromov:2018:NOC

[GBSRS18] Dmitry Gromov, Ingo Bulla, Oana Silvia Serea, and Ethan O. Romero-Severson. Numerical optimal control for HIV prevention with dynamic budget allocation. Mathematical Medicine and Biology, 35(4):469–491, December 2018. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/4/469/4585730.

Gonzalez:2008:GTC

[GC08] Sara J. González and Daniel G. Carando. A general tumour control probability model for non-uniform dose distributions. Mathematical Medicine and Biology, 25(2):171–184, June 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/2/171/751948.

Gray:2015:RSS

[GC15] Catheryn W. Gray and Adelle C. F. Coster. A receptor state space model of the insulin signalling system in glucose transport. Mathematical Medicine and Biology, 32(4):457–473, December 1, 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/4/457/2874299.

Gu:2012:APS

[GCC^{+}12] S. Gu, G. Chakraborty, K. Champley, A. M. Alessio, J. Claridge, R. Rockne, M. Muzi, K. A. Krohn, A. M. Spence, E. C. Alvord, A. R. A. Anderson, P. E. Kinahan, and K. R. Swanson. Applying a patient-specific bio-mathematical model of glioma growth to develop virtual [18f]-FMISO-PET images. Mathematical Medicine and Biology, 29(1):31–48, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/1/31/753901.

Grist:1998:HST

[GD98] Eric P. M. Grist and Sophie Des Clers. How seasonal temperature variations may influence the structure of annual squid populations. Mathematical Medicine and Biology, 15(2):187–209, June 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/15/2/187/656883.

Gong:2013:MCT

[GDFH13] Jiafen Gong, Mairon M. Dos Santos, Chris Finlay, and Thomas Hillen. Are more complicated tumour control probability models better? Mathematical Medicine and Biology, 30(1): 1–19, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/1/1/710594.

Greenhalgh:2001:MTA

[GDL01] David Greenhalgh, Murray Doyle, and Fraser Lewis. A mathematical treatment of AIDS and condom use. Mathematical Medicine and Biology, 18(3):225–262, September 1, 2001. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
18/3/225/657414.

Griffiths:2000:ASC

[GEW00] Jeff Griffiths, Tracey England, and Janet Williams. Analytic solutions to compartmental models of the HIV/AIDS epidemic. Mathematical Medicine and Biology, 17(4):295–310, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/4/295/664705.

Gatenby:2008:ICM

[GF08] Robert A. Gatenby and B. Roy Frieden. Inducing catastrophe in malignant growth. Mathematical Medicine and Biology, 25 (3):267–283, September 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/25/3/267/682272.

Guy:2007:FGF

[GFK07] Robert D. Guy, Aaron L. Fogelson, and James P. Keener. Fibrin gel formation in a shear flow. Mathematical Medicine and Biology, 24(1):111–130, March 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/24/1/111/692196.

Golinelli:2006:BIH

[GGA06] D. Golinelli, P. Guttorp, and J. A. Abkowitz. Bayesian inference in a hidden stochastic two-compartment model for feline hematopoiesis. Mathematical Medicine and Biology, 23(3):153– 172, September 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/23/3/153/672274.

Greenhalgh:1997:MMS

[GH97] David Greenhalgh and Gordon Hay. Mathematical modelling of the spread of HIV/AIDS amongst injecting drug users. Mathematical Medicine and Biology, 14(1):11–38, March 1997. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
14/1/11/659993.

Greenman:1999:MMS

[GH99] J. V. Greenman and P. J. Hudson. Multihost, multiparasite systems: an application of bifurcation theory. Mathematical Medicine and Biology, 16(4):333–367, December 1999. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
16/4/333/2874292.

Gleason:2005:CMM

[GH05] Rudolph L. Gleason, Jr. and Jay D. Humphrey. A 2D constrained mixture model for arterial adaptations to large changes in flow, pressure and axial stretch. Mathematical Medicine and Biology, 22(4):347–369, December 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/22/4/347/650317.

Gomez:1990:EAP

[GJ90] G. Gómez and O. Julià. Estimation and asymptotic properties of the distribution of tune-to-tumour in carcinogenesis experiments. Mathematical Medicine and Biology, 7(2):109–123, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/7/2/109/698791.

Greenhalgh:2000:TSA

[GL00] David Greenhalgh and Fraser Lewis. Three-stage AIDS incubation period: a best case scenario using addict-needle interaction assumptions. Mathematical Medicine and Biology, 17 (2):95–118, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/17/2/95/736873.

Gaffney:2015:STS

[GL15] E. A. Gaffney and S. Seirin Lee. The sensitivity of Turing self-organization to biological feedback delays: 2D models of fish pigmentation. Mathematical Medicine and Biology, 32(1): 57–79, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/1/57/661581.

Gao:2017:MDA

[GLDP17] Daozhou Gao, Thomas M. Lietman, Chao-Ping Dong, and Travis C. Porco. Mass drug administration: the importance of synchrony. Mathematical Medicine and Biology, 34(2):241–260, June 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/34/2/241/2885295.

Gomez:2019:UDM

[GM19] Miquel Gomez and Francisco Maduell. A unidimensional diffusion model applied to uremic toxin kinetics in haemodiafiltration treatments. Mathematical Medicine and Biology, 36(2): 223–240, June 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/36/2/223/5036512.

Gaffney:1999:MCE

[GMM^{+}99] E. A. Gaffney, P. K. Maini, C. D. Mccaig, M. Zhao, and J. V. Forrester. Modelling corneal epithelial wound closure in the presence of physiological electric fields via a moving boundary formalism. Mathematical Medicine and Biology, 16(4):369– 393, December 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/4/369/2874293.

Grindrod:1989:SSS

[GMS89] P. Grindrod, J. D. Murray, and S. Sinha. Steady-state spatial patterns in a cell-chemotaxis model. Mathematical Medicine and Biology, 6(2):69–79, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/6/2/69/651317.

Guttorp:1990:SMH

[GNA90] Peter Guttorp, Michael A. Newton, and Janis L. Abkowitz. A stochastic model for haematopoiesis in cats. Mathematical Medicine and Biology, 7(2):125–143, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/7/2/125/
698792.

Gani:1984:MGM

[GP84] J. Gani and P. Purdue. Matrix-geometric methods for the general stochastic epidemic. Mathematical Medicine and Biology, 1(4):333–342, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/1/4/333/768071.

Gorban:2011:CDW

[GP11] Alexander Gorban and Sergei Petrovskii. Collective dynamics: when one plus one does not make two. Mathematical Medicine and Biology, 28(2):85–88, June 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/28/2/85/679776.

Giverso:2012:MCR

[GP12] C. Giverso and L. Preziosi. Modelling the compression and reorganization of cell aggregates. Mathematical Medicine and Biology, 29(2):181–204, June 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/29/2/181/662212.

Gibson:1998:EPS

[GR98] Gavin J. Gibson and Eric Renshaw. Estimating parameters in stochastic compartmental models using Markov chain methods. Mathematical Medicine and Biology, 15(1):19–40, March 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
15/1/19/683936.

Greenhalgh:1987:ARS

[Gre87] David Greenhalgh. Analytical results on the stability of agestructured recurrent epidemic models. Mathematical Medicine and Biology, 4(2):109–144, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/4/2/109/873584.

Greenhalgh:1988:TSR

[Gre88] David Greenhalgh. Threshold and stability results for an epidemic model with an age-structured meeting rate. Mathematical Medicine and Biology, 5(2):81–100, ???? 1988. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
5/2/81/731024.

Greenhalgh:1990:EMD

[Gre90] David Greenhalgh. An epidemic model with a densitydependent death rate. Mathematical Medicine and Biology, 7 (1):1–26, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/7/1/1/715230.

Greenhalgh:1992:SRS

[Gre92] David Greenhalgh. Some results for an SEIR epidemic model with density dependence in the death rate. Mathematical Medicine and Biology, 9(2):67–106, ???? 1992. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
9/2/67/731784.

Grindrod:1984:CPA

[GS84a] P. Grindrod and B. D. Sleeman. Comparison principles in the analysis of reaction–diffusion systems modelling unmyelinated nerve fibres. Mathematical Medicine and Biology, 1(4):343–363, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/1/4/343/768076.

Grindrod:1984:QAR

[GS84b] P. Grindrod and B. D. Sleeman. Qualitative analysis of reaction–diffusion systems modelling coupled unmyelinated nerve axons. Mathematical Medicine and Biology, 1(3):289–307, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/1/3/289/846453.

Garnett:1992:GII

[GSBA92] Geoffrey P. Garnett, Jonathan Swinton, Robert C. Brunham, and Roy M. Anderson. Gonococcal infection, infertility, and population growth: II. The influence of heterogeneity in sexual behaviour. Mathematical Medicine and Biology, 9(2):127–144, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/9/2/127/731776.

Griffiths:1987:TDA

[GSW87] J. D. Griffiths, J. K. Smedley, and T. G. Weale. Terminal distributions along a ‘Knight’s Line’ for a stochastic epidemic. Mathematical Medicine and Biology, 4(1):69–79, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
4/1/69/875677.

Guo:1991:MCE

[GT91] Sun Wei Guo and Elizabeth A. Thompson. Monte Carlo estimation of variance component models for large complex pedigrees. Mathematical Medicine and Biology, 8(3):171–189, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
8/3/171/689703.

Ghosh:1999:DJE

[GT99] Asit K. Ghosh and P. K. Tapaswi. Dynamics of Japanese encephalitis — a study in mathematical epidemiology. Mathematical Medicine and Biology, 16(1):1–27, March 1999. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
16/1/1/729495.

Green:2018:PFM

[GWO^{+}18] J. E. F. Green, J. P. Whiteley, J. M. Oliver, H. M. Byrne, and S. L. Waters. Pattern formation in multiphase models of chemotactic cell aggregation. Mathematical Medicine and Biology, 35(3):319–346, September 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/3/319/3829521.

Gani:1993:MSH

[GY93] J. Gani and S. Yakowitz. Modelling the spread of HIV among intravenous drug users. Mathematical Medicine and Biology, 10(1):51–65, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/1/51/749598.

Gomez:2019:SMM

[GY19] Miller Cerón Gómez and Hyun Mo Yang. A simple mathematical model to describe antibody-dependent enhancement in heterologous secondary infection in dengue. Mathematical Medicine and Biology, 36(4):411–438, December 2019. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
36/4/411/5123755.

Huynh:2012:MMA

[HA12] Giao T. Huynh and Frederick R. Adler. Mathematical modelling the age dependence of Epstein-Barr virus associated infectious mononucleosis. Mathematical Medicine and Biology, 29 (3):245–261, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/29/3/245/798004.

Harbron:1995:PBA

[Har95] Chris Harbron. A pedigree-based algorithm for finding efficient peeling sequences. Mathematical Medicine and Biology, 12(1):13–27, ???? 1995. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/12/1/13/707653.

Hassell:1984:PPE

[Has84] Michael P. Hassell. Parasitism in patchy environments: Inverse density dependence can be stabilizing. Mathematical Medicine and Biology, 1(1):123–133, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/1/1/123/754925.

Heryudono:2007:SEM

[HBD^{+}07] A. Heryudono, R. J. Braun, T. A. Driscoll, K. L. Maki, L. P. Cook, and P. E. King-Smith. Single-equation models for the tear film in a blink cycle: realistic lid motion. Mathematical Medicine and Biology, 24(4):347–377, December 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/4/347/661376.

Hughes:2022:PEE

[HBHJ22] Elliott Hughes, Rachelle Binny, Shaun Hendy, and Alex James. Predicting elimination of evolving virus variants. Mathematical Medicine and Biology, 39(4):410–424, December 2022. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
39/4/410/6668241.

Hatzikirou:2012:GGK

[HBS^{+}12] H. Hatzikirou, D. Basanta, M. Simon, K. Schaller, and A. Deutsch. ‘go or grow’: the key to the emergence of invasion in tumour progression? Mathematical Medicine and Biology, 29 (1):49–65, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/29/1/49/753882.

Hanis:1984:NSH

[HC84] Craig L. Hanis and Ranajit Chakraborty. Nonrandom sampling in human genetics: Familial correlations. Mathematical Medicine and Biology, 1(2):193–213, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/1/2/193/
643669.

Hsieh:2000:BCT

[HC00] Ying-Hen Hsieh and Kenneth Cooke. Behaviour change and treatment of core groups: its effect on the spread of HIV/AIDS. Mathematical Medicine and Biology, 17(3):213–241, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/3/213/667730.

Hill:2017:UTC

[HCWK17] Lydia Hill, Mark A. J. Chaplain, Roland Wolf, and Yury Kapelyukh. The usage of a three-compartment model to investigate the metabolic differences between hepatic reductase null and wild-type mice. Mathematical Medicine and Biology, 34 (1):1–13, March 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/34/1/1/2885283.

Houy:2020:IUE

[HF20] Nicolas Houy and Julien Flaig. Informed and uninformed empirical therapy policies. Mathematical Medicine and Biology, 37 (3):334–350, September 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/37/3/334/5686327.

Haque:2010:WPA

[HG10] Mainul Haque and David Greenhalgh. When a predator avoids infected prey: a model-based theoretical study. Mathematical Medicine and Biology, 27(1):75–94, March 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/27/1/75/
731489.

Hsieh:2008:PPM

[HH08] Ying-Hen Hsieh and Chin-Kuei Hsiao. Predator–prey model with disease infection in both populations. Mathematical Medicine and Biology, 25(3):247–266, September 2008. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/3/247/682283.

Hiby:1985:AEP

[Hib85] A. R. Hiby. An approach to estimating population densities of great whales from sighting surveys. Mathematical Medicine and Biology, 2(3):201–220, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/2/3/201/682029.

Higham:2007:SAT

[HKV07] Desmond J. Higham, Gabriela Kalna, and J. Keith Vass. Spectral analysis of two-signed microarray expression data. Mathematical Medicine and Biology, 24(2):131–148, June 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/2/131/685956.

Hartmann:2007:MAF

[HM07] Dirk Hartmann and Takashi Miura. Mathematical analysis of a free-boundary model for lung branching morphogenesis.

Mathematical Medicine and Biology, 24(2):209–224, June 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/2/209/685952.

Hyndman:2023:STM

[HMM23] Lauren Hyndman, Sean McKee, and Sean McGinty. Solute transport with Michaelis–Menten kinetics for in vitro cell culture. Mathematical Medicine and Biology, 40(1):49–72, March 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/40/1/49/6748977.

Hughes:1997:VMM

[HMMN97] Gareth Hughes, Neil Mcroberts, Laurence V. Madden, and Scot C. Nelson. Validating mathematical models of plantdisease progress in space and time. Mathematical Medicine and Biology, 14(2):85–112, June 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/14/2/85/726607.

Heyden:2019:FOS

[HO19] S. Heyden and M. Ortiz. Functional optimality of the sulcus pattern of the human brain. Mathematical Medicine and Biology, 36(2):207–221, June 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/36/2/207/5017347.

Hoppensteadt:1984:SON

[Hop84] F. C. Hoppensteadt. Synchronized oscillations in networks of neuron analogue circuits. Mathematical Medicine and Biology, 1(2):135–148, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/1/2/135/643661.

Horwood:1984:VRB

[Hor84a] J. W. Horwood. The variance and response of biological systems to variability in births and survivals. Mathematical Medicine and Biology, 1(3):309–322, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/1/3/309/
846459.

Horwood:1984:FRE

[Hor84b] Joseph W. Horwood. The frequency response of exploited fish Stocks to external perturbations. Mathematical Medicine and Biology, 1(2):215–231, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/1/2/215/643674.

Horwood:1990:NOR

[Hor90] Joseph W. Horwood. Near-optimal rewards from multiple species harvested by several fishing fleets. Mathematical Medicine and Biology, 7(1):55–68, ???? 1990. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
7/1/55/715243.

Horwood:1996:RSO

[Hor96] J. W. Horwood. Risk-sensitive optimal harvesting and control of biological populations. Mathematical Medicine and Biology, 13(1):35–71, March 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/13/1/35/644513.

Hentzel:1990:KTO

[HPH90] Irvin Roy Hentzel, Luiz Antonio Peresi, and Philip Holgate. On k th-order Bernstein algebras and stability at the k + 1 generation in polyploids. Mathematical Medicine and Biology, 7(1):33–40, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/7/1/33/715235.

Han:2023:CSS

[HPR^{+}23] Seong Woo Han, Charles Puelz, Craig G. Rusin, Daniel J. Penny, Ryan Coleman, and Charles S. Peskin. Computer simulation of surgical interventions for the treatment of refractory pulmonary hypertension. Mathematical Medicine and Biology, 40(1):1–23, March 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/40/1/1/6671730.

Hanin:2016:UNH

[HR16] Leonid Hanin and Jason Rose. Uncovering the natural history of cancer from post-mortem cross-sectional diameters of hepatic metastases. Mathematical Medicine and Biology, 33(4):397– 416, December 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/33/4/397/2674334.

Hsieh:1996:TSMa

[Hsi96a] Ying-Hen Hsieh. A two-sex model for treatment of HIV/AIDS and behaviour change in a population of varying size. Mathematical Medicine and Biology, 13(3):151–173, September 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/3/151/651304.

Hsieh:1996:TSMb

[Hsi96b] Ying-Hen Hsieh. A two-sex model for treatment of HIV/AIDS and behaviour change in a population of varying size. Mathematical Medicine and Biology, 13(4):317, December 1996. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/4/317/771445.

Hussain:2024:IST

[HSV24] Saddam Hussain, Moitri Sen, and Vitaly Volpert. Infection spreading in tissue as a reaction–diffusion wave. Mathematical Medicine and Biology, 41(3):169–191, September 2024. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
41/3/169/7713939.

Harbron:1994:AGR

[HT94] Chris Harbron and Alun Thomas. Alternative graphical representations of genotypes in a pedigree. Mathematical Medicine and Biology, 11(4):217–228, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/11/4/217/649635.

Hoogenveen:2010:CDP

[HvBB10] Rudolf T. Hoogenveen, Pieter H. M. van Baal, and Hendriek C. Boshuizen. Chronic disease projections in heterogeneous ageing populations: approximating multi-state models of joint distributions by modelling marginal distributions. Mathematical Medicine and Biology, 27(1):1–19, March 2010. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/1/1/731498.

Hermange:2024:MMS

[HVPC24] Gurvan Hermange, William Vainchenker, Isabelle Plo, and Paul-Henry Cournède. Mathematical modelling, selection and hierarchical inference to determine the minimal dose in IFNα therapy against myeloproliferative neoplasms. Mathematical Medicine and Biology, 41(2):110–134, June 2024. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
41/2/110/7693621.

Horwood:1986:OCN

[HW86a] J. W. Horwood and P. Whittle. Optimal control in the neighbourhood of an optimal equilibrium with examples from fisheries models. Mathematical Medicine and Biology, 3(2): 129–142, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/2/129/752666.

Horwood:1986:OHM

[HW86b] J. W. Horwood and P. Whittle. The optimal harvest from a multicohort stock. Mathematical Medicine and Biology, 3(2): 143–155, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/2/143/752675.

Hussaini:2011:QAR

[HWG11] N. Hussaini, M. Winter, and A. B. Gumel. Qualitative assessment of the role of public health education program on HIV transmission dynamics. Mathematical Medicine and Biology, 28(3):245–270, September 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/28/3/245/664315.

Haydon:1997:AFM

[HWK97] D. T. Haydon, M. E. J. Woolhouse, and R. P. Kitching. An analysis of foot-and-mouth-disease epidemics in the UK. Mathematical Medicine and Biology, 14(1):1–9, March 1997. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
14/1/1/659985.

Han:2024:DOR

[HYM^{+}24] Lifeng Han, Osman N. Yogurtcu, Marisabel Rodriguez Messan, Wencel Valega-Mackenzie, Ujwani Nukala, and Hong Yang. Dosage optimization for reducing tumor burden using a phenotype-structured population model with a drugresistance continuum. Mathematical Medicine and Biology, 41 (1):35–52, March 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/41/1/35/7614350.

Hamel:2025:KTF

[HZ25] François Hamel and Mingmin Zhang. KPP transition fronts in a one-dimensional two-patch habitat. Mathematical Medicine and Biology, 42(1):71–97, March 2025. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/42/1/71/7724036.

Iles:2004:SGW

[IC04] Mark M. Iles and Chris Cannings. Sequential genotyping within TDT families. Mathematical Medicine and Biology, 21(2): 115–127, June 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/21/2/115/698800.

Ismail:2013:FME

[IFP13] Z. Ismail, A. D. Fitt, and C. P. Please. A fluid mechanical explanation of the spontaneous reattachment of a previously detached Descemet membrane. Mathematical Medicine and Biology, 30(4):339–355, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/30/4/339/718704.

Iwasa:1989:AME

[ILA89] Yoh Iwasa, Simon A. Levin, and Viggo Andreasen. Aggregation in model ecosystems II. Approximate aggregation. Mathematical Medicine and Biology, 6(1):1–23, ???? 1989. CODEN MM-BABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL

http://academic.oup.com/imammb/article/6/1/1/666278.

Irvine:1994:SMG

[IMM94] Victor Irvine, Sally Mcclean, and Peter Millard. Stochastic models for geriatric in-patient behaviour. Mathematical Medicine and Biology, 11(3):207–216, ???? 1994. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
11/3/207/708104.

Ibragimov:2005:MMA

[IMRW05] A. I. Ibragimov, C. J. McNeal, L. R. Ritter, and J. R. Walton. A mathematical model of atherogenesis as an inflammatory response. Mathematical Medicine and Biology, 22(4):305– 333, December 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/22/4/305/650308.

Iima:2012:PTM

[IN12] Makoto Iima and Toshiyuki Nakagaki. Peristaltic transport and mixing of cytosol through the whole body of Physarum plasmodium. Mathematical Medicine and Biology, 29(3):263– 281, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/29/3/263/798015.

Iwasa:1985:EMS

[IR85] Yoh Iwasa and Jonathan Roughgarden. Evolution in a metapopolation with space-limited subpopulations. Mathematical Medicine and Biology, 2(2):93–107, ???? 1985. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
2/2/93/765079.

Jager:1986:SPG

[Jäg86] Edgar Jäger. Stable patterns generated by activator–inhibitor systems. Mathematical Medicine and Biology, 3(3):179–190, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/3/3/179/716898.

Jarrett:2019:MMT

[JBG^{+}19] Angela M. Jarrett, Meghan J. Bloom, Wesley Godfrey, Anum K. Syed, David A. Ekrut, Lauren I. Ehrlich, Thomas E. Yankeelov, and Anna G. Sorace. Mathematical modelling of trastuzumab-induced immune response in an in vivo murine model of HER2+ breast cancer. Mathematical Medicine and Biology, 36(3):381–410, September 2019. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/36/3/381/5101740.

Jarrett:2019:UDA

[JC19] Angela M. Jarrett and Nicholas G. Cogan. The ups and downs of S. aureus nasal carriage. Mathematical Medicine and Biology, 36(2):157–177, June 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/36/2/157/4995147.

Jarrett:2015:MIB

[JCS15] Angela M. Jarrett, N. G. Cogan, and M. E. Shirtliff. Modelling the interaction between the host immune response, bacterial dynamics and inflammatory damage in comparison with immunomodulation and vaccination experiments. Mathematical Medicine and Biology, 32(3):285–306, September 2015. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
32/3/285/646313.

Jensen:2003:E

[JKK03] Oliver E. Jensen, John R. King, and James P. Keener. Editorial. Mathematical Medicine and Biology, 20(1):??, March 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/20/1/i/663537.

Jensen:2013:MMA

[JKK13] Oliver Jensen, Jim Keener, and John King. Mathematical Medicine & Biology: a journal of the IMA Best Paper Prize. Mathematical Medicine and Biology, 30(3):??, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
30/3/NP/684901.

Jung:2008:OCA

[JLPB08] Eunok Jung, Suzanne Lenhart, Vladimir Protopopescu, and Charles Babbs. Optimal control applied to a thoraco– abdominal CPR model. Mathematical Medicine and Biology, 25 (2):157–170, June 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/25/2/157/751939.

Jepson:2023:PFT

[JOBF23] Jacob M. Jepson, Reuben D. O’Dea, John Billingham, and Nabil T. Fadai. Pattern formation and travelling waves in a multiphase moving boundary model of tumour growth. Mathematical Medicine and Biology, 40(4):327–347, December 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
40/4/327/7444987.

James:2021:SMC

[JPB^{+}21] A. James, M. J. Plank, R. N. Binny, A. Lustig, K. Hannah, S. C. Hendy, and N. Steyn. A structured model for COVID-19 spread: modelling age and healthcare inequities. Mathematical Medicine and Biology, 38(3):299–313, September 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/3/299/6276864.

Jones:2005:DTF

[JPM^{+}05] M. B. Jones, C. P. Please, D. L. S. McElwain, G. R. Fulford, A. P. Roberts, and M. J. Collins. Dynamics of tear film deposition and draining. Mathematical Medicine and Biology, 22 (3):265–288, September 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/22/3/265/652731.

Jeger:2002:MPV

[JvdBD02] M. J. Jeger, F. van den Bosch, and M. Y. Dutmer. Modelling plant virus epidemics in a plantation-nursery system. Mathematical Medicine and Biology, 19(2):75–94, June 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/19/2/75/
730715.

Jeger:1998:MAP

[JvdBMH98] M. J. Jeger, F. van den Bosch, L. V. Madden, and J. Holt. A model for analysing plant-virus transmission characteristics and epidemic development. Mathematical Medicine and Biology, 15 (1):1–18, March 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/15/1/1/683935.

Kakeshashi:1998:MAS

[Kak98] Masayuki Kakeshashi. A mathematical analysis of the spread of HIV/AIDS in Japan. Mathematical Medicine and Biology, 15 (4):299–311, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/15/4/299/678298.

Kakehashi:1999:MAS

[Kak99] Masayuki Kakehashi. A mathematical analysis of the spread of HIV/AIDS in Japan. Mathematical Medicine and Biology, 16 (1):111–112, March 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/1/111/729503.

Kaplan:2001:EPH

[Kap01] Edward H. Kaplan. Evaluating plasma holds in the presence of multiple infections. Mathematical Medicine and Biology, 18(3): 215–224, September 1, 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/18/3/215/657409.

Krzyzanski:2007:MGF

[KB07] Wojciech Krzyzanski and Jonathan Bell. Modelling the geometric features and investigating electrical properties of dendrites in a fish thalamic neuron. Mathematical Medicine and Biology, 24(3):271–286, September 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/24/3/271/712080.

Krause:2019:LCM

[KBVW19] Andrew L. Krause, Dmitry Beliaev, Robert A. Van Gorder, and Sarah L. Waters. Lattice and continuum modelling of a bioactive porous tissue scaffold. Mathematical Medicine and Biology, 36(3):325–360, September 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/36/3/325/5068240.

Kenber:1995:ASM

[KE95] G. C. Kenber and J. D. Evans. Analytical solutions to a multicyEinder somatic shunt cable model for passive neurones with spines. Mathematical Medicine and Biology, 12(2):137–157, June 1995. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/12/2/137/666447.

Keener:2006:SCO

[Kee06] James P. Keener. Stochastic calcium oscillations. Mathematical Medicine and Biology, 23(1):1–25, March 2006. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/1/1/809614.

Khan:1999:HBE

[KG99] Q. J. A. Khan and David Greenhalgh. Hopf bifurcation in epidemic models with a time delay in vaccination. Mathematical Medicine and Biology, 16(2):113–142, June 1999. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
16/2/113/668974.

Kelly:1999:UDD

[KGG^{+}99] L. A. Kelly, G. Gibson, G. Gettinby, W. Donachie, and J. C. Low. The use of dummy data points when fitting bacterial growth curves. Mathematical Medicine and Biology, 16(2): 155–170, June 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/2/155/668986.

Kiemel:1987:MPS

[KH87] Tim Kiemel and Philip Holmes. A model for the periodic synaptic inhibition of a neuronal oscillator. Mathematical Medicine and Biology, 4(2):145–169, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/4/2/145/2268134.

Kimura:1984:EAT

[Kim84] Motoo Kimura. Evolution of an altruistic trait through group selection as studied by the diffusion equation method. Mathematical Medicine and Biology, 1(1):1–15, ???? 1984. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
1/1/1/754908.

Karmakar:1988:LDL

[KJ88] N. Karmakar and G. Jayaraman. Linear diffusion of lead in the intestinal wall: a theoretical study. Mathematical Medicine and Biology, 5(1):33–43, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/5/1/33/763977.

Kuznetsov:2015:CBD

[KK15] I. A. Kuznetsov and A. V. Kuznetsov. A comparison between the diffusion–reaction and slow axonal transport models for predicting tau distribution along an axon. Mathematical Medicine and Biology, 32(3):263–283, September 2015. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
32/3/263/646312.

Kuznetsov:2017:WMT

[KK17] I. A. Kuznetsov and A. V. Kuznetsov. What mechanisms of tau protein transport could be responsible for the inverted tau concentration gradient in degenerating axons? Mathematical Medicine and Biology, 34(1):125–150, March 2017. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
34/1/125/2885290.

Kuznetsov:2022:ASS

[KK22] Ivan A. Kuznetsov and Andrey V. Kuznetsov. An analytical solution simulating growth of Lewy bodies. Mathematical Medicine and Biology, 39(3):299–312, September 2022. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
39/3/299/6598653.

King:2003:MHT

[KKC^{+}03] J. R. King, A. J. Koerber, J. M. Croft, J. P. Ward, P. Williams, and R. E. Sockett. Modelling host tissue degradation by extracellular bacterial pathogens. Mathematical Medicine and Biology, 20(3):227–260, September 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/20/3/227/714217.

Kopocinska:2004:CDA

[KKO04] I. Kopocińska, B. Kopociński, and A. Okulewicz. A class of distributions applicable to the description of the number of nematodes parasitizing birds. Mathematical Medicine and Biology, 21(1):35–48, March 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/21/1/35/689379.

Kalachev:2011:ATT

[KKO^{+}11] L. V. Kalachev, T. C. Kelly, M. J. O’Callaghan, A. V. Pokrovskii, and A. V. Pokrovskiy. Analysis of threshold-type behaviour in mathematical models of the intrusion of a novel macroparasite in a host colony. Mathematical Medicine and Biology, 28(4):287–333, December 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/28/4/287/658301.

Khader:2014:IHM

[KLG^{+}14] Karim Khader, Molly Leecaster, Tom Greene, Matthew Samore, and Alun Thomas. Improved hidden Markov model for nosocomial infections. Mathematical Medicine and Biology, 31 (4):338–352, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/31/4/338/646892.

Korobeinikov:2005:NLI

[KM05] Andrei Korobeinikov and Philip K. Maini. Non-linear incidence and stability of infectious disease models. Mathematical Medicine and Biology, 22(2):113–128, June 2005. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
22/2/113/770946.

Kaslik:2018:SHB

[KN18] Eva Kaslik and Mihaela Neamtu. Stability and Hopf bifurcation analysis for the hypothalamic-pituitary-adrenal axis model with memory. Mathematical Medicine and Biology, 35(1): 49–78, March 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/35/1/49/2739353.

Karev:2011:ABS

[KNB11] Georgy P. Karev, Artem S. Novozhilov, and Faina S. Berezovskaya. On the asymptotic behaviour of the solutions to the replicator equation. Mathematical Medicine and Biology, 28 (2):89–110, June 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/28/2/89/679755.

Kohlmann:2013:AMM

[Koh13] Martin Kohlmann. Analysis of a mathematical model for the growth of cancer cells. Mathematical Medicine and Biology, 30 (2):175–189, June 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/2/175/890423.

Korobeinikov:2004:LFG

[Kor04] Andrei Korobeinikov. Lyapunov functions and global properties for SEIR and SEIS epidemic models. Mathematical Medicine and Biology, 21(2):75–83, June 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/21/2/75/
698794.

Korobeinikov:2009:GAP

[Kor09a] Andrei Korobeinikov. Global asymptotic properties of virus dynamics models with dose-dependent parasite reproduction and virulence and non-linear incidence rate. Mathematical Medicine and Biology, 26(3):225–239, September 2009. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
26/3/225/808439.

Korobeinikov:2009:SEG

[Kor09b] Andrei Korobeinikov. Stability of ecosystem: global properties of a general predator–prey model. Mathematical Medicine and Biology, 26(4):309–321, December 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/26/4/309/654364.

Karpouzas:1991:MNO

[KP91] I. Karpouzas and Y. Pouliquen. Modelling and numerical optimization of corneal rotation. Mathematical Medicine and Biology, 8(1):73–82, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/8/1/73/788908.

Keeling:2009:MMD

[KPSW09] Stephen L. Keeling, Georg Propst, Georg Stadler, and Werner Wackernagel. A mathematical model for the deformation of the eyeball by an elastic band. Mathematical Medicine and Biology, 26(2):165–185, June 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/26/2/165/727676.

Konar:1990:PDC

[KR90] A. Konar and A. B. Roy. Population drift and control of damping in a predator–prey system. Mathematical Medicine and Biology, 7(4):245–259, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/7/4/245/744276.

Kroon:2010:CMS

[Kro10] Martin Kroon. A constitutive model for smooth muscle including active tone and passive viscoelastic behaviour. Mathematical Medicine and Biology, 27(2):129–155, June 2010. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/2/129/695264.

Krzyzanski:2001:AWS

[Krz01] Wojciech Krzyzanski. An application of a weak solution of the cable equation to the Rall model of a nerve cell. Mathematical Medicine and Biology, 18(4):377–407, December 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
18/4/377/747275.

Korobeinikov:2016:PES

[KSS16] Andrei Korobeinikov, Elena Shchepakina, and Vladimir Sobolev. Paradox of enrichment and system order reduction: bacteriophages dynamics as case study. Mathematical Medicine and Biology, 33(3):359–369, September 2016. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
33/3/359/1750308.

Krawczyk:1993:CST

[KT93] Jacek B. Krawczyk and Bolesław Tolwinski. A cooperative solution for the three-nation problem of exploitation of the southern bluefin tuna. Mathematical Medicine and Biology, 10(2): 135–147, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/2/135/654147.

Kuznetsov:2013:MAC

[Kuz13] A. V. Kuznetsov. Modelling of axonal cargo rerouting in a dendrite. Mathematical Medicine and Biology, 30(3):273– 285, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/3/273/684868.

Kuznetsov:2024:EDA

[Kuz24] Andrey V. Kuznetsov. Effect of diffusivity of amyloid beta monomers on the formation of senile plaques. Mathematical Medicine and Biology, 41(4):346–362, December 2024. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
41/4/346/7822270.

Kimpton:2013:MTW

[KWW^{+}13] L. S. Kimpton, J. P. Whiteley, S. L. Waters, J. R. King, and J. M. Oliver. Multiple travelling-wave solutions in a minimal model for cell motility. Mathematical Medicine and Biology, 30 (3):241–272, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/30/3/241/684877.

Li:2016:CTF

[LBD^{+}16] Longfei Li, Richard J. Braun, Tobin A. Driscoll, William D. Henshaw, Jeffrey W. Banks, and P. Ewen King-Smith. Computed tear film and osmolarity dynamics on an eye-shaped domain. Mathematical Medicine and Biology, 33(2):123–157, June 1, 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/33/2/123/2875345.

Li:2018:CFF

[LBHKS18] Longfei Li, R. J. Braun, W. D. Henshaw, and P. E. King-Smith. Computed flow and fluorescence over the ocular surface. Mathematical Medicine and Biology, 35(S1):??, April 2018. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL https://academic.oup.com/imammb/article/
35/Supplement_1/i51/4259013.

Lakawicz:2015:AMB

[LBPF15] Joseph M. Lakawicz, William J. Bottega, Jonathan L. Prenner, and Howard F. Fine. An analysis of the mechanical behaviour of a detaching retina. Mathematical Medicine and Biology, 32 (2):137–161, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/2/137/830129.

Leite:2000:BRR

[LBY00] Maria Beatriz Ferreira Leite, Rodney Carlos Bassanezi, and Hyun Mo Yang. The basic reproduction ratio for a model of directly transmitted infections considering the virus charge and the immunological response. Mathematical Medicine and Biology, 17(1):15–31, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/17/1/15/674979.

Liu:1998:SAB

[LC98] Junli Liu and John W. Crawford. Stability of an autocatalytic biochemical system in the presence of noise perturbations. Mathematical Medicine and Biology, 15(4):339–350, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/15/4/339/678306.

Liu:2004:PCL

[LC04] Bing Liu and Lansun Chen. The periodic competing Lotka– Volterra model with impulsive effect. Mathematical Medicine and Biology, 21(2):129–145, June 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/21/2/129/698806.

Lefebvre:2017:SMT

[LCC^{+}17] Guillaume Lefebvre, François Cornelis, Patricio Cumsille, Thierry Colin, Clair Poignard, and Olivier Saut. Spatial modelling of tumour drug resistance: the case of GIST liver metastases. Mathematical Medicine and Biology, 34(2):151–176, June 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/34/2/151/2885292.

Lounes:1999:TTM

[LdA99] Rachid Lounes and Hector de Arazoza. A two-type model for the Cuban national programme on HIV/AIDS. Mathematical Medicine and Biology, 16(2):143–154, June 1999. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
16/2/143/668980.

Lewis:1992:UGT

[Lew92] Cathryn M. Lewis. The use of graph theory techniques to investigate genealogical structure. Mathematical Medicine and Biology, 9(3):145–159, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/9/3/145/649538.

Leiderman:2011:GFS

[LF11] Karin Leiderman and Aaron L. Fogelson. Grow with the flow: a spatial-temporal model of platelet deposition and blood coagulation under flow. Mathematical Medicine and Biology, 28 (1):47–84, March 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/28/1/47/679981.

Lindwall:2024:IID

[LG24] Gustav Lindwall and Philip Gerlee. Inference on an interacting diffusion system with application to in vitro glioblastoma migration (publication template). Mathematical Medicine and Biology, 41(3):250–276, September 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/41/3/250/7732137.

Li:1987:DVS

[LHZ87] Jai Li, Thomas G. Hallam, and Ma Zhien. Demographic variation and survival in discrete population models. Mathematical Medicine and Biology, 4(3):237–246, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/4/3/237/
874279.

Le:2019:MMW

[LJR^{+}19] Thuy T. T. Le, Felix Jost, Thomas Raupach, Jakob Zierk, Manfred Rauh, Meinolf Suttorp, Martin Stanulla, Markus Metzler, and Sebastian Sager. A mathematical model of white blood cell dynamics during maintenance therapy of childhood acute lymphoblastic leukemia. Mathematical Medicine and Biology, 36 (4):471–488, December 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/36/4/471/5136501.

Lemon:2007:MMC

[LK07] G. Lemon and J. R. King. Multiphase modelling of cell behaviour on artificial scaffolds: effects of nutrient depletion and spatially nonuniform porosity. Mathematical Medicine and Biology, 24(1):57–83, March 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/24/1/57/692191.

Lee:1998:MMR

[LKH98] A. J. Lee, J. R. King, and S. Hibberd. Mathematical modelling of the release of drug from porous, nonswelling transdermal drug-delivery devices. Mathematical Medicine and Biology, 15 (2):135–163, June 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/15/2/135/656873.

Lenbury:1996:DSF

[LKN96] Yongwimon Lenbury, Kanchana Kumnungkit, and Boriboon Novaprateep. Detection of slow-fast limit cycles in a model for electrical activity in the pancreatic β-cell. Mathematical Medicine and Biology, 13(1):1–21, March 1996. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/1/1/644501.

Lee:1996:MPM

[LKR96] A. J. Lee, J. R. King, and T. G. Rogers. A multiple-pathway model for the diffusion of drugs in skin. Mathematical Medicine and Biology, 13(2):127–150, June 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/13/2/127/655284.

Lawson:2000:AST

[LL00] Andrew B. Lawson and Petra Leimich. Approaches to the space-time modelling of infectious disease behaviour. Mathematical Medicine and Biology, 17(1):1–13, ???? 2000. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/1/1/674972.

Lavi:2009:WCC

[LL09] Orit Lavi and Yoram Louzoun. What cycles the cell? — Robust autonomous cell cycle models. Mathematical Medicine and Biology, 26(4):337–359, December 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/26/4/337/654369.

Lu:2017:MST

[LLZ^{+}17] Yichen Lu, Mei Yan Lee, Shu Zhu, Talid Sinno, and Scott L. Diamond. Multiscale simulation of thrombus growth and vessel occlusion triggered by collagen/tissue factor using a datadriven model of combinatorial platelet signalling. Mathematical Medicine and Biology, 34(4):523–546, December 2017. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
34/4/523/2885305.

Lauwerier:1986:HBH

[LM86] H. A. Lauwerier and J. A. J. Metz. Hopf bifurcation in hostparasitoid models. Mathematical Medicine and Biology, 3(3): 191–210, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/3/191/716915.

Lane:1987:AWP

[LMM87] D. C. Lane, J. D. Murray, and V. S. Manoranjan. Analysis of wave phenomena in a morphogenetic mechanochemical model and an application to post-fertilization waves on eggs. Mathematical Medicine and Biology, 4(4):309–331, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
4/4/309/751263.

Ledzewicz:2010:RPO

[LMMS10] Urszula Ledzewicz, John Marriott, Helmut Maurer, and Heinz Schättler. Realizable protocols for optimal administration of drugs in mathematical models for anti-angiogenic treatment. Mathematical Medicine and Biology, 27(2):157–179, June 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/2/157/695256.

Lenbury:2000:HDS

[LOT00] Yongwimon Lenbury, Rujira Ouncharoen, and Nardtida Tumrasvin. Higher-dimensional separation principle for the analysis of relaxation oscillations in nonlinear systems: application to a model of HIV infection. Mathematical Medicine and Biology, 17 (3):243–261, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/17/3/243/667734.

Lounes:1989:TTP

[Lou89] Rachid Lounes. A two-type population epidemic problem. Mathematical Medicine and Biology, 6(3):205–208, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
6/3/205/684139.

Landman:2001:TDN

[LP01] K. A. Landman and C. P. Please. Tumour dynamics and necrosis: surface tension and stability. Mathematical Medicine and Biology, 18(2):131–158, June 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/18/2/131/672234.

Lenbury:2005:DDE

[LP05] Yongwimon Lenbury and Pornsarp Pornsawad. A delaydifferential equation model of the feedback-controlled hypothalamuspituitary-adrenal axis in humans. Mathematical Medicine and Biology, 22(1):15–33, March 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/22/1/15/657677.

Lunelli:2008:EEA

[LP08] Antonella Lunelli and Andrea Pugliese. Evaluating the effectiveness of antiviral treatment in models for influenza pandemic. Mathematical Medicine and Biology, 25(4):359– 372, December 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/25/4/359/826442.

Loy:2021:SNL

[LP21] Nadia Loy and Luigi Preziosi. Stability of a non-local kinetic model for cell migration with density-dependent speed. Mathematical Medicine and Biology, 38(1):83–105, March 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/1/83/6041785.

Li:2024:ICC

[LPSR24] Li Li, Yi-Zhi Pang, Gui-Quan Sun, and Shigui Ruan. Impact of climate change on vegetation patterns in Altay Prefecture, China. Mathematical Medicine and Biology, 41(1): 53–80, March 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/41/1/53/7616119.

Louie:1993:TSA

[LRW93] K. Louie, M. G. Roberts, and G. C. Wake. Thresholds and stability analysis of models for the spatial spread of a fatal disease. Mathematical Medicine and Biology, 10(3):207–226, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/10/3/207/666915.

Louie:1994:RAS

[LRW94] K. Louie, M. G. Roberts, and G. C. Wake. The regulation of an age-structured population by a fatal disease. Mathematical Medicine and Biology, 11(4):229–244, ???? 1994. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
11/4/229/649655.

Lin:1993:AIM

[LTW93] Shili Lin, Elizabeth Thompson, and Ellen Wijsman. Achieving irreducibility of the Markov chain Monte Carlo method applied to pedigree data. Mathematical Medicine and Biology, 10(1):1–17, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/10/1/1/749586.

Li:2015:TDI

[LTY15] Yanyan Li, Moxun Tang, and Jianshe Yu. Transcription dynamics of inducible genes modulated by negative regulations. Mathematical Medicine and Biology, 32(2):115–136, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
32/2/115/830120.

Lu:2011:OCI

[Lu11] Lili Lu. Optimal control of input rates of Stein’s models. Mathematical Medicine and Biology, 28(1):31–46, March 2011. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
28/1/31/679976.

Ludwig:1989:IUE

[Lud89] Donald Ludwig. Irreducible uncertainty in estimation from catch and effort data. Mathematical Medicine and Biology, 6 (4):269–275, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/6/4/269/794064.

Lutscher:2006:ESM

[Lut06] Frithjof Lutscher. Book review: Exactly Solvable Models for Biological Invasion by S. Petrovskii and B.-L. Li. Mathematical Medicine and Biology, 23(4):389–390, December 2006. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/4/389/689801.

Lamour:1998:SBS

[LvdB98] A. Lamour and F. van den Bosch. The stability of the biocontrol system in an epizootic insect-pathogen model incorporating three routes of infection. Mathematical Medicine and Biology, 15(4):313–320, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/15/4/313/678302.

Lamour:2002:QSS

[LvdBTJ02] A. Lamour, F. van den Bosch, A. J. Termorshuizen, and M. J. Jeger. Quasi-steady state approximation to a fungal growth model. Mathematical Medicine and Biology, 19(3):163– 183, September 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/19/3/163/729865.

Lamour:2000:MGS

[LVTJ00] A. Lamour, F. Van Den Bosch, A. J. Termorshuizen, and M. J. Jeger. Modelling the growth of soil-borne fungi in response to carbon and nitrogen. Mathematical Medicine and Biology, 17 (4):329–346, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/17/4/329/664723.

Lamour:2001:MGS

[LVTJ01] A. Lamour, F. Van Den Bosch, A. J. Termorshuizen, and M. J. Jeger. Modelling the growth of soil-borne fungi in response to carbon and nitrogen. Mathematical Medicine and Biology, 18(2):213, June 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/18/2/213/672241.

Lang:2017:MMB

[LVWG17] Georgina E. Lang, Dominic Vella, Sarah L. Waters, and Alain Goriely. Mathematical modelling of blood-brain barrier failure and oedema. Mathematical Medicine and Biology, 34(3):391– 414, September 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/34/3/391/2885299.

Moghadas:2006:BEM

[MA06] Seyed M. Moghadas and Murray E. Alexander. Bifurcations of an epidemic model with non-linear incidence and infectiondependent removal rate. Mathematical Medicine and Biology, 23(3):231–254, September 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/23/3/231/672278.

McCormack:2007:DEM

[MA07] Robert K. McCormack and Linda J. S. Allen. Disease emergence in multi-host epidemic models. Mathematical Medicine and Biology, 24(1):17–34, March 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/24/1/17/692193.

Magnusson:1992:FPS

[Mag92] Kjartan G. Magnusson. A feedback and probing strategy to regulate harvesting from a renewable resource. Mathematical Medicine and Biology, 9(1):43–65, ???? 1992. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
9/1/43/664828.

Maini:2003:EMB

[Mai03] Philip K. Maini. Essential mathematical biology. Mathematical Medicine and Biology, 20(2):225–226, June 2003. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/2/225/658774.

Morozov:2011:NRP

[MANS11] Andrew Morozov, Elena Arashkevich, Anastasia Nikishina, and Konstantin Solovyev. Nutrient-rich plankton communities stabilized via predator–prey interactions: revisiting the role of vertical heterogeneity. Mathematical Medicine and Biology, 28 (2):185–215, June 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/28/2/185/679764.

Matteson:2008:EKM

[Mat08] Andrew Matteson. An enzyme kinetic model of blood island formation. Mathematical Medicine and Biology, 25(3):233– 245, September 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/25/3/233/682278.

Maki:2020:ILR

[MBB20] Kara L. Maki, Richard J. Braun, and Gregory A. Barron. The influence of a lipid reservoir on the tear film formation. Mathematical Medicine and Biology, 37(3):363–388, September 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/3/363/5716121.

Miller:2014:MMP

[MBC^{+}14] J. K. Miller, J. S. Brantner, C. Clemons, K. L. Kreider, A. Milsted, P. Wilber, Y. H. Yun, W. J. Youngs, G. Young, H. T. Badawy, A. Milsted, C. Clemons, K. L. Kreider, P. Wilber, G. Young, Y. H. Yun, P. O. Wagers, and W. J. Youngs. Mathematical modelling of Pseudomonas aeruginosa biofilm growth and treatment in the cystic fibrosis lung. Mathematical Medicine and Biology, 31(2):179–204, June 2014. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
31/2/179/660584.

Maki:2008:OGM

[MBDKS08] K. L. Maki, R. J. Braun, T. A. Driscoll, and P. E. King-Smith. An overset grid method for the study of reflex tearing. Mathematical Medicine and Biology, 25(3):187–214, September 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/3/187/682269.

Mode:1984:MOC

[MBEP84] C. J. Mode, R. C. Busby, D. C. Ewbank, and G. T. Pickens. A mathematical overview of a computer simulation model of maternity histories with illustrative examples. Mathematical Medicine and Biology, 1(1):107–121, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/1/1/107/
754917.

Maki:2010:TFD

[MBHKS10] Kara L. Maki, Richard J. Braun, William D. Henshaw, and P. Ewen King-Smith. Tear film dynamics on an eyeshaped domain I: pressure boundary conditions. Mathematical Medicine and Biology, 27(3):227–254, September 2010. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/3/227/791833.

Maini:1992:PFR

[MBS92] Philip K. Maini, Debbie L. Benson, and Jonathan A. Sherratt. Pattern formation in reaction–diffusion models with spatially inhomogeneous diffusion coefficients. Mathematical Medicine and Biology, 9(3):197–213, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/9/3/197/649554.

Milne:1998:SAF

[MC98] Alice E. Milne and Zaid S. Chalabi. Stability analysis of the FitzHugh–Nagumo differential equations driven by impulses: Applied to the electrical firing of magnocellular neurons. Mathematical Medicine and Biology, 15(4):367–385, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
15/4/367/678311.

Milne:2001:CAR

[MC01] Alice E. Milne and Zaid S. Chalabi. Control analysis of the Rose-Hindmarsh model for neural activity. Mathematical Medicine and Biology, 18(1):53–75, March 1, 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/18/1/53/
727373.

Matzavinos:2004:MMS

[MCK04] Anastasios Matzavinos, Mark A. J. Chaplain, and Vladimir A. Kuznetsov. Mathematical modelling of the spatio-temporal response of cytotoxic T-lymphocytes to a solid tumour. Mathematical Medicine and Biology, 21(1):1–34, March 2004. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
21/1/1/689371.

Markovic:2016:MCE

[MČM^{+}16] Vladimir M. Marković, Željko Čupić, Stevan Maćešić, Ana Stanojević, Vladana Vukojević, and Ljiljana Kolar-Anić. Modelling cholesterol effects on the dynamics of the hypothalamic– pituitary–adrenal (HPA) axis. Mathematical Medicine and Biology, 33(1):1–28, March 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/33/1/1/2363477.

Magal:2008:CIH

[MCRC08] Christelle Magal, Chris Cosner, Shigui Ruan, and J. Casas. Control of invasive hosts by generalist parasitoids. Mathematical Medicine and Biology, 25(1):1–20, March 2008. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/1/1/745857.

Manem:2018:MRS

[MD18] V. S. K. Manem and A. Dhawan. Modelling recurrence and second cancer risks induced by proton therapy. Mathematical Medicine and Biology, 35(3):347–361, September 2018. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
35/3/347/3925038.

Montcho:2024:AME

[MDS^{+}24] Yvette Montcho, Sidoine Dako, Valère Kolawole Salako, Chénangnon Frédéric Tovissodé, Martin Wolkewitz, and Romain Glèlè Kakaï. Assessing marginal effects of nonpharmaceutical interventions on the transmission of SARS-CoV-2 across Africa: a hybrid modeling study. Mathematical Medicine and Biology, 41(3):225–249, September 2024. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
41/3/225/7724873.

Mikhasev:2010:SSS

[MEB10] Gennady Mikhasev, Sergey Ermochenko, and Matthias Bornitz. On the strain-stress state of the reconstructed middle ear after inserting a malleus-incus prosthesis. Mathematical Medicine and Biology, 27(4):289–312, December 2010. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/4/289/724973.

Mirzaei:2021:SMA

[MF21] Navid Mohammad Mirzaei and Pak-Wing Fok. Simple model of atherosclerosis in cylindrical arteries: impact of anisotropic growth on Glagov remodeling. Mathematical Medicine and Biology, 38(1):59–82, March 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/38/1/59/5894145.

McLean:2006:SMM

[MG06] Douglas R. McLean and Bruce P. Graham. Stability in a mathematical model of neurite elongation. Mathematical Medicine and Biology, 23(2):101–117, June 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/23/2/101/671412.

Mbah:2014:OCD

[MG14] Martial L. Ndeffo Mbah and Christopher A. Gilligan. Optimal control of disease infestations on a lattice. Mathematical Medicine and Biology, 31(1):87–97, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/31/1/87/
860038.

Martinez-Gonzalez:2015:CTA

[MGDPC^{+}15] Alicia Martínez-González, Mario Durán-Prado, Gabriel F. Calvo, Francisco J. Alcaín, Luis A. Pérez-Romasanta, and Víctor M. Pérez-García. Combined therapies of antithrombotics and antioxidants delay in silico brain tumour progression. Mathematical Medicine and Biology, 32(3):239– 262, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/3/239/646309.

Mode:1989:MSN

[MGSS89] Charles J. Mode, Herman E. Gollwitzer, Michael A. Salsburg, and Candace K. Sleeman. A methodological study of a nonlinear stochastic model of an AIDS epidemic with recruitment. Mathematical Medicine and Biology, 6(3):179–203, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
6/3/179/684133.

Mortensen:2021:APP

[MGSS21a] Peter Mortensen, Hao Gao, Godfrey Smith, and Radostin D. Simitev. Action potential propagation and block in a model of atrial tissue with myocyte-fibroblast coupling. Mathematical Medicine and Biology, 38(1):106–131, March 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/1/106/6067662.

Mortensen:2021:AAP

[MGSS21b] Peter Mortensen, Hao Gao, Godfrey Smith, and Radostin D. Simitev. Addendum: Action potential propagation and block in a model of atrial tissue with myocyte-fibroblast coupling. Mathematical Medicine and Biology, 38(3):292–298, September 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/38/3/292/6271027.

Merlushkin:1997:SBI

[MH97] Anton Merlushkin and Alan G. Hawkes. Stochastic behaviour of ion channels in varying conditions. Mathematical Medicine and Biology, 14(2):125–149, June 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/14/2/125/726596.

Myers:2018:MMA

[MHGY18] Matthew R. Myers, Prasanna Hariharan, Suvajyoti Guha, and Jing Yan. A mathematical model for assessing the effectiveness of protective devices in reducing risk of infection by inhalable droplets. Mathematical Medicine and Biology, 35(1): 1–23, March 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/35/1/1/2433376.

Malunguza:2017:DPT

[MHMDM17] N. J. Malunguza, S. D. Hove-Musekwa, S. Dube, and Z. Mukandavire. Dynamical properties and thresholds of an HIV model with super-infection. Mathematical Medicine and Biology, 34 (4):493–522, December 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/34/4/493/2885304.

Meyer-Hermann:2006:ABC

[MHMI06] Michael E. Meyer-Hermann, Philip K. Maini, and Dagmar Iber. An analysis of B cell selection mechanisms in germinal centers. Mathematical Medicine and Biology, 23(3):255–277, September 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/23/3/255/672282.

Malchow:2011:CDI

[MJB11] Horst Malchow, Alex James, and Richard Brown. Competition and diffusive invasion in a noisy environment. Mathematical Medicine and Biology, 28(2):153–163, June 2011. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
28/2/153/679759.

Malice:1989:RVI

[MK89] Marie-Pierre Malice and Richard J. Kryscio. On the role of variable incubation periods in simple epidemic models. Mathematical Medicine and Biology, 6(4):233–242, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/6/4/233/
794002.

Maass:2020:MDP

[MK20] Kelsey Maass and Minsun Kim. A Markov decision process approach to optimizing cancer therapy using multiple modalities. Mathematical Medicine and Biology, 37(1):22–39, March 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/1/22/5376663.

Mahdi:2000:VCH

[ML00] Smail Mahdi and Sabin Lessard. Variability in centred house-of-cards mutation models. Mathematical Medicine and Biology, 17 (3):185–200, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/17/3/185/667723.

Maler:2010:CCT

[ML10] Adrian Maler and Frithjof Lutscher. Cell-cycle times and the tumour control probability. Mathematical Medicine and Biology, 27(4):313–342, December 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/27/4/313/724979.

Mardia:1992:PHF

[MLH92] K. V. Mardia, Q. Li, and T. J. Hainsworth. On the Penrose hypothesis on fingerprint patterns. Mathematical Medicine and Biology, 9(4):289–294, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/9/4/289/730285.

Mena-Lorca:1999:DDD

[MLVHCC99] Jaime Mena-Lorca, Jorge X. Velasco-Hernandez, and Carlos Castillo-Chavez. Density-dependent dynamics and superinfection in an epidemic model. Mathematical Medicine and Biology, 16(4):307–317, December 1999. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/16/4/307/652974.

Magnus:1987:EUS

[MM87] R. J. Magnus and K. G. Magnusson. Existence and uniqueness of solutions to the multispecies virtual population analysis equations. Mathematical Medicine and Biology, 4(3):247–263, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/4/3/247/874287.

Maier:2021:OAV

[MMA^{+}21] Sandra B. Maier, Eduardo Massad, Marcos Amaku, Marcelo N. Burattini, and David Greenhalgh. The optimal age of vaccination against dengue in Brazil based on serotype-specific forces of infection derived from serological data. Mathematical Medicine and Biology, 38(1):1–27, March 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/38/1/1/5863710.

Mukherjee:2024:GIT

[MMF24] Avishek Mukherjee, Navid Mohammad Mirzaei, and Pak-Wing Fok. Genesis of intimal thickening due to hemodynamical shear stresses. Mathematical Medicine and Biology, 41(4):363– 381, December 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/41/4/363/7822271.

McGinty:2015:RMP

[MMMW15] Sean McGinty, Sean McKee, Christopher McCormick, and Marcus Wheel. Release mechanism and parameter estimation in drug-eluting stent systems: analytical solutions of drug release and tissue transport. Mathematical Medicine and Biology, 32(2):163–186, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/32/2/163/830140.

Mourad:2022:SMM

[MMT22] Ayman Mourad, Fatima Mroue, and Zahraa Taha. Stochastic mathematical models for the spread of COVID-19: a novel epidemiological approach. Mathematical Medicine and Biology, 39 (1):49–76, March 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/39/1/49/6457811.

McGinty:2011:MDE

[MMWM11] Sean McGinty, Sean McKee, Roger M. Wadsworth, and Christopher McCormick. Modelling drug-eluting stents. Mathematical Medicine and Biology, 28(1):1–29, March 2011. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
28/1/1/679971.

Marchant:2006:BBM

[MNB06] Ben P. Marchant, John Norbury, and Helen M. Byrne. Biphasic behaviour in malignant invasion. Mathematical Medicine and Biology, 23(3):173–196, September 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/23/3/173/672270.

Murray:1984:GBP

[MO84] J. D. Murray and G. F. Oster. Generation of biological pattern and form. Mathematical Medicine and Biology, 1(1):51–75, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/1/1/51/754940.

Mode:1988:SAC

[Mod88] Charles J. Mode. On statistically assessing critical population size of an endangered species in a random environment. Mathematical Medicine and Biology, 5(2):147–166, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
5/2/147/731016.

Mode:1997:TPS

[Mod97] Charles J. Mode. Threshold parameters for a simple stochastic partnership model of sexually transmitted diseases formulated as a two-type CMJ process. Mathematical Medicine and Biology, 14(4):251–260, December 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/14/4/251/655353.

Moneim:2007:SVE

[Mon07] I. A. Moneim. Seasonally varying epidemics with and without latent period: a comparative simulation study. Mathematical Medicine and Biology, 24(1):1–15, March 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/1/1/692198.

Mode:1987:DHU

[MPE87] C. J. Mode, G. T. Pickens, and D. C. Ewbank. Demographic heterogeneity and uncertainty in population projections. Mathematical Medicine and Biology, 4(3):223–236, ???? 1987. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
4/3/223/874272.

MacArthur:2005:MMD

[MPP05] Ben D. MacArthur, Colin P. Please, and Graham J. Pettet. A mathematical model of dynamic glioma-host interactions: receptor-mediated invasion and local proteolysis. Mathematical Medicine and Biology, 22(3):247–264, September 2005. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
22/3/247/652730.

Maddalena:2020:ESN

[MR20] Lucia Maddalena and Stefania Ragni. Existence of solutions and numerical approximation of a non-local tumor growth model. Mathematical Medicine and Biology, 37(1): 58–82, March 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/37/1/58/5424065.

Marion:1998:SEM

[MRG98] Glenn Marion, Eric Renshaw, and Gavin Gibson. Stochastic effects in a model of nematode infection in ruminants. Mathematical Medicine and Biology, 15(2):97–116, June 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/15/2/97/
656886.

Matveev:2001:OCR

[MS01] Alexey S. Matveev and Andrey V. Savkin. Optimal chemotherapy regimens: influence of tumours on normal cells and several toxicity constraints. Mathematical Medicine and Biology, 18(1): 25–40, March 1, 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/18/1/25/727369.

Manfredi:2002:PIO

[MS02] Piero Manfredi and Ernesto Salinelli. Population-induced oscillations in blended SI-SEI epidemiological models. Mathematical Medicine and Biology, 19(2):95–112, June 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/19/2/95/
730718.

Matveev:2005:ITN

[MS05] Alexey S. Matveev and Andrey V. Savkin. Influence of tumours on normal cells and optimal chemotherapy regimens: the case of several drugs and toxicity constraints. Mathematical Medicine and Biology, 22(2):143–162, June 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/22/2/143/770967.

Moulton:2013:LOM

[MS13] Michael J. Moulton and Timothy W. Secomb. A low-order model for left ventricle dynamics throughout the cardiac cycle. Mathematical Medicine and Biology, 30(1):45–63, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/30/1/45/710600.

Musio:2010:BSP

[MSB10] Monica Musio, Erik A. Sauleau, and Antoine Buemi. Bayesian semi-parametric ZIP models with space-time interactions: an application to cancer registry data. Mathematical Medicine and Biology, 27(2):181–194, June 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/27/2/181/695272.

McInerney:2004:MFC

[MSBM04] D. McInerney, S. Schnell, R. E. Baker, and P. K. Maini. A mathematical formulation for the cell-cycle model in somitogenesis: analysis, parameter constraints and numerical solutions. Mathematical Medicine and Biology, 21(2):85–113, June 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
21/2/85/698797.

Mateos:1985:GCS

[MSjH^{+}85] J. C. Mateos, F. Sanchez, j. Horno, J. Ureña, J. R. Zaragoza, and G. Madurga. A geometric calculation of square and circular size equivalents for rectangular radiation fields in radiotherapy. Mathematical Medicine and Biology, 2(2):131–137, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/2/2/131/765070.

Mohseni-Salehi:2018:ESR

[MSZMGFS18] Fazeleh S. Mohseni-Salehi, Fatemeh Zare-Mirakabad, Soudeh Ghafouri-Fard, and Mehdi Sadeghi. The effect of stochasticity on repair of DNA double strand breaks throughout nonhomologous end joining pathway. Mathematical Medicine and Biology, 35(4):517–539, December 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/4/517/4718073.

Munoz:2016:NRM

[Muñ16] Ana I. Muñoz. Numerical resolution of a model of tumour growth. Mathematical Medicine and Biology, 33(1): 57–85, March 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/33/1/57/2363522.

Murray:1995:EED

[Mur95] J. M. Murray. An example of the effects of drug resistance on the optimal schedule for a single drug in cancer chemotherapy. Mathematical Medicine and Biology, 12(1):55–69, ???? 1995. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
12/1/55/707669.

Murray:1997:OST

[Mur97] J. M. Murray. The optimal scheduling of two drugs with simple resistance for a problem in cancer chemotherapy. Mathematical Medicine and Biology, 14(4):283–303, December 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
14/4/283/655359.

Mwambi:2002:TTB

[Mwa02] Henry G. Mwambi. Ticks and tick-borne diseases in Africa: a disease transmission model. Mathematical Medicine and Biology, 19(4):275–292, December 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/19/4/275/753208.

Naasell:1986:SM

[Nås86] Ingemar Nåsell. On superinfection in malaria. Mathematical Medicine and Biology, 3(3):211–227, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/3/3/211/
716917.

Nikin-Beers:2018:MOA

[NBC18] Ryan Nikin-Beers and Stanca M. Ciupe. Modelling original antigenic sin in dengue viral infection. Mathematical Medicine and Biology, 35(2):257–272, June 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/2/257/3056195.

Nisbet:1985:SSM

[NBGM85] R. M. Nisbet, S. P. Blythe, W. S. C. Gurney, and J. A. J. Metz. Stage-structure models of populations with distinct growth and development processes. Mathematical Medicine and Biology, 2 (1):57–68, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/2/1/57/815672.

Nestor-Bergmann:2018:RCS

[NBGWJ18] Alexander Nestor-Bergmann, Georgina Goddard, Sarah Woolner, and Oliver E. Jensen. Relating cell shape and mechanical stress in a spatially disordered epithelium using a vertexbased model. Mathematical Medicine and Biology, 35(S1):??, April 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL https://academic.oup.com/imammb/
article/35/Supplement_1/i1/4079751.

Nkwayep:2022:MMC

[NBT^{+}22] C. H. Nkwayep, S. Bowong, B. Tsanou, M. A. Aziz Alaoui, and J. Kurths. Mathematical modeling of COVID-19 pandemic in the context of sub-Saharan Africa: a short-term forecasting in Cameroon and Gabon. Mathematical Medicine and Biology, 39 (1):1–48, March 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/39/1/1/6511580.

Nudel:2017:CSI

[NDd17] Iftah Nudel, Luis Dorfmann, and Gal deBotton. The compartment syndrome: is the intra-compartment pressure a reliable indicator for early diagnosis? Mathematical Medicine and Biology, 34(4):547–558, December 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/34/4/547/2419538.

Nourollahi:2019:OMS

[NGK19] Sevnaz Nourollahi, Archis Ghate, and Minsun Kim. Optimal modality selection in external beam radiotherapy. Mathematical Medicine and Biology, 36(3):361–380, September 2019. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
36/3/361/5089942.

Nagorcka:1992:SSP

[NM92] B. N. Nagorcka and J. R. Mooney. From stripes to spots: Prepatterns which can be produced in the skin by a reaction– diffusion system. Mathematical Medicine and Biology, 9(4): 249–267, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/9/4/249/730276.

Nani:1994:MSR

[NO94] Frank K. Nani and M. Namik Oğuztöreli. Modelling and simulation of Rosenberg-type adoptive cellular immunotherapy. Mathematical Medicine and Biology, 11(2):107–147, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
11/2/107/771436.

Nani:1999:MSC

[NO99] Frank K. Nani and M. Namik Oğuztöreli. Modelling and simulation of chemotherapy of haematological and gynaecological cancers. Mathematical Medicine and Biology, 16(1): 39–91, March 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/1/39/729544.

Noori:2012:EAA

[Noo12] H. R. Noori. The effects of the acute administration of lowdosage ethanol on the phasic neurochemical oscillations of the basal ganglia. Mathematical Medicine and Biology, 29(3):231– 244, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/29/3/231/797998.

Nielsen:2014:BAE

[NPO14] Kenneth H. M. Nielsen, Flemming M. Pociot, and Johnny T. Ottesen. Bifurcation analysis of an existing mathematical model reveals novel treatment strategies and suggests potential cure for type 1 diabetes. Mathematical Medicine and Biology, 31(3):205–225, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/3/205/710024.

Nokes:1995:CCV

[NS95] D. James Nokes and Jonathan Swinton. The control of childhood viral infections by pulse vaccination. Mathematical Medicine and Biology, 12(1):29–53, ???? 1995. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/12/1/29/
707662.

Nevai:2014:MST

[NS14] Andrew L. Nevai and Edy Soewono. A model for the spatial transmission of dengue with daily movement between villages and a city. Mathematical Medicine and Biology, 31(2):150–178, June 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/31/2/150/660581.

Nguyen:1997:EGF

[NSCO97] H. T. Nguyen, A. G. Shannon, P. A. Coates, and D. R. Owens. Estimation of glomerular filtration rate in type II (non-insulin dependent) diabetes mellitus patients. Mathematical Medicine and Biology, 14(2):151–160, June 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/14/2/151/726601.

Norbury:1993:SPF

[NW93] J. Norbury and G. C. Wake. Spatial pattern formation for steady states of a population model. Mathematical Medicine and Biology, 10(1):19–30, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/10/1/19/749591.

Oleinick:2006:MRN

[OAG^{+}06] Alexander I. Oleinick, Christian Amatore, Manon Guille, Stephane Arbault, Oleksiy V. Klymenko, and Irina Svir. Modelling release of nitric oxide in a slice of rat’s brain: describing stimulated functional hyperemia with diffusion–reaction equations. Mathematical Medicine and Biology, 23(1):27–44, March 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/23/1/27/809607.

Orme:1996:MMF

[OC96] M. E. Orme and M. A. J. Chaplain. A mathematical model of the first steps of tumour-related angiogenesis: Capillary sprout formation and secondary branching. Mathematical Medicine and Biology, 13(2):73–98, June 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/13/2/73/655288.

Orme:1997:TDM

[OC97] M. E. Orme and M. A. J. Chaplain. Two-dimensional models of tumour angiogenesis and anti-angiogenesis strategies. Mathematical Medicine and Biology, 14(3):189–205, September 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
14/3/189/2874290.

Ohta:1984:PGT

[Oht84] Tomoka Ohta. Population genetics of transposable elements. Mathematical Medicine and Biology, 1(1):17–29, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
1/1/17/754930.

ODea:2013:ISP

[OK13] R. D. O’Dea and J. R. King. The isolation of spatial patterning modes in a mathematical model of juxtacrine cell signalling.

Mathematical Medicine and Biology, 30(2):95–113, June 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
30/2/95/890406.

Oliver:2005:TFT

[OKM^{+}05] J. M. Oliver, J. R. King, K. J. McKinlay, P. D. Brown, D. M. Grant, C. A. Scotchford, and J. V. Wood. Thin-film theories for two-phase reactive flow models of active cell motion. Mathematical Medicine and Biology, 22(1):53–98, March 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/22/1/53/
657681.

ONeill:1999:BMD

[O’N99] Philip D. O’Neill. On a branching model of division-within-division. Mathematical Medicine and Biology, 16(4):395– 407, December 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/4/395/652981.

ONeill:2013:MTU

[O’N13] Philip D. O’Neill. Book review: Mathematical Tools for Understanding Infectious Disease Dynamics by O. Diekmann, H. Heesterbeek and T. Britton, Princeton University Press, pp. 516, ISBN 978-0-691-15539-5. Mathematical Medicine and Biology, 30(4):383, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/30/4/383/718836.

ODea:2015:MAN

[ONE^{+}15] R. D. O’Dea, M. R. Nelson, A. J. El Haj, S. L. Waters, and H. M. Byrne. A multiscale analysis of nutrient transport and biological tissue growth in vitro. Mathematical Medicine and Biology, 32(3):345–366, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/32/3/345/646322.

Orme:2001:EEF

[OOB01] B. A. A. Orme, S. R. Otto, and J. R. Blake. Enhanced efficiency of feeding and mixing due to chaotic flow patterns around choanoflagellates. Mathematical Medicine and Biology, 18(3):293–325, September 1, 2001. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/18/3/293/657419.

Owen:1998:MMI

[OS98] Markus R. Owen and Jonathan A. Sherratt. Modelling the macrophage invasion of tumours: Effects on growth and composition. Mathematical Medicine and Biology, 15(2):165–185, June 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/15/2/165/656878.

Olsen:1997:MMC

[OSMA97] Luke Olsen, Jonathan A. Sherratt, Philip K. Maini, and Frank Arnold. A mathematical model for the capillary endothelial cell-extracellular matrix interactions in wound-healing angiogenesis. Mathematical Medicine and Biology, 14(4):261– 281, December 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/14/4/261/655355.

Oelz:2005:MNL

[OSS05] Dietmar Oelz, Christian Schmeiser, and Alexander Soreff. Multistep navigation of leukocytes: a stochastic model with memory effects. Mathematical Medicine and Biology, 22(4):291– 303, December 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/22/4/291/650304.

ODea:2010:MMT

[OWB10] R. D. O’Dea, S. L. Waters, and H. M. Byrne. A multiphase model for tissue construct growth in a perfusion bioreactor. Mathematical Medicine and Biology, 27(2):95–127, June 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/2/95/695249.

Oliver:1998:BCE

[OWL^{+}98] M. A. Oliver, R. Webster, C. Lajaunie, K. R. Muir, S. E. Parkes A. H. Cameron, M. C. G. Stevens, and J. R. Mann. Binomial cokriging for estimating and mapping the risk of childhood cancer. Mathematical Medicine and Biology, 15(3):279– 297, September 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/15/3/279/795661.

Panaretos:1989:PMI

[Pan89] J. Panaretos. A probability model involving the use of the zero-truncated Yule distribution for analysing surname data. Mathematical Medicine and Biology, 6(2):133–136, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
6/2/133/651315.

Parthasarathy:1997:ESD

[Par97] P. R. Parthasarathy. The effect of superinfection on the distribution of the infectious period — a continued fraction approximation. Mathematical Medicine and Biology, 14(2):113–123, June 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/14/2/113/726588.

Proctor:2000:SMA

[PB00] C. J. Proctor and M. Broom. A spatial model of antipredator vigilance. Mathematical Medicine and Biology, 17(1):75–93, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/17/1/75/675003.

Picchini:2008:MLE

[PDD08] Umberto Picchini, Susanne Ditlevsen, and Andrea De Gaetano. Maximum likelihood estimation of a time-inhomogeneous stochastic differential model of glucose dynamics. Mathematical Medicine and Biology, 25(2):141–155, June 2008. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/2/141/751944.

Puelz:2021:MPK

[PDR^{+}21] Charles Puelz, Zach Danial, Jay S. Raval, Jonathan L. Marinaro, Boyce E. Griffith, and Charles S. Peskin. Models for plasma kinetics during simultaneous therapeutic plasma exchange and extracorporeal membrane oxygenation. Mathematical Medicine and Biology, 38(2):255–271, June 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/2/255/6149285.

Papa:2020:IAI

[PFF^{+}20] Federico Papa, Giovanni Felici, Marco Franzetti, Alberto Gandolfi, and Carmela Sinisgalli. Impact of ART-induced viral suppression on the HIV epidemic in Italy. Mathematical Medicine and Biology, 37(2):183–211, June 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/37/2/183/5510091.

Perthame:2010:STM

[PG10] Benoît Perthame and Mathias Gauduchon. Survival thresholds and mortality rates in adaptive dynamics: conciliating deterministic and stochastic simulations. Mathematical Medicine and Biology, 27(3):195–210, September 2010. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/3/195/791802.

Perez-Garcia:2015:DER

[PGBMG^{+}15] Víctor M. Pérez-García, Magdalena Bogdanska, Alicia Martínez-González, Juan Belmonte-Beitia, Philippe Schucht, and Luis A. Pérez-Romasanta. Delay effects in the response of low-grade gliomas to radiotherapy: a mathematical model and its therapeutical implications. Mathematical Medicine and Biology, 32 (3):307–329, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/32/3/307/646317.

Paulo:2000:MDV

[PGC^{+}00] A. C. Paulo, M. C. Gomes, A. C. Casinhas, A. Horta, and T. Domingos. Multiple dose vaccination against childhood diseases: high coverage with the first dose remains crucial for eradication. Mathematical Medicine and Biology, 17(3):201–212, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/17/3/201/667725.

Poehler:2022:EDC

[PGL^{+}22] Elliot Poehler, Liam Gibson, Audrey Lustig, Nicole J. Moreland, Reuben McGregor, and Alex James. Estimating decay curves of neutralizing antibodies to SARS-CoV-2 infection. Mathematical Medicine and Biology, 39(4):368–381, December 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/39/4/368/6618579.

Perez-Garcia:2016:EPL

[PGPR16] Víctor M. Pérez-García and Luis A. Pérez-Romasanta. Extreme protraction for low-grade gliomas: theoretical proof of concept of a novel therapeutical strategy. Mathematical Medicine and Biology, 33(3):253–271, September 2016. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
33/3/253/1750136.

Phelps:1991:MRS

[Phe91a] Iv F. M. Phelps. Multicomponent rank selection as an alternative to Haldane’s dilemma. Mathematical Medicine and Biology, 8(1):57–72, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/8/1/57/788900.

Phelps:1991:UMS

[Phe91b] Iv F. M. Phelps. A unifying model for the substitutional genetic load. Mathematical Medicine and Biology, 8(1):31–56, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/8/1/31/788893.

Phelps:1995:MEG

[Phe95] Iv F. M. Phelps. A model for the evolution of the genome: The effect of stochasticity on genetic loads. Mathematical Medicine and Biology, 12(1):1–11, ???? 1995. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/12/1/1/707640.

Plank:2022:PRT

[PJL^{+}22] Michael J. Plank, Alex James, Audrey Lustig, Nicholas Steyn, Rachelle N. Binny, and Shaun C. Hendy. Potential reduction in transmission of COVID-19 by digital contact tracing systems: a modelling study. Mathematical Medicine and Biology, 39(2): 156–168, June 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/39/2/156/6548784.

Patrone:2021:CUU

[PK21] Paul N. Patrone and Anthony J. Kearsley. Classification under uncertainty: data analysis for diagnostic antibody testing. Mathematical Medicine and Biology, 38(3):396–416, September 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/38/3/396/6336022.

Puri:2023:GOM

[PK23] Ashishi Puri and Sanjeev Kumar. A generalized order mixture model for tracing connectivity of white matter fascicles complexity in brain from diffusion MRI. Mathematical Medicine and Biology, 40(3):223–237, September 2023. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
40/3/223/7112766.

Pontrelli:2017:ITD

[PLFP17] Giuseppe Pontrelli, Marco Lauricella, José A. Ferreira, and Gonçalo Pena. Iontophoretic transdermal drug delivery: a multi-layered approach. Mathematical Medicine and Biology, 34(4):559–576, December 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/34/4/559/2417354.

Pickens:1986:PMV

[PM86] Gary T. Pickens and Charles J. Mode. Projection of mean and variance functions for population processes with timehomogeneous laws of evolution. Mathematical Medicine and Biology, 3(1):1–22, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/3/1/1/684210.

Pantha:2023:EVT

[PMAV23] Buddhi Pantha, Jemal Mohammed-Awel, and Naveen K. Vaidya. Effects of vaccination on the two-strain transmission dynamics of COVID-19: Dougherty County, Georgia, USA, as a case study. Mathematical Medicine and Biology, 40(4):308– 326, December 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/40/4/308/7420502.

Pandiyan:2014:PSM

[PMB14] Balamurugan Pandiyan, Stephen J. Merrill, and Salvatore Benvenga. A patient-specific model of the negative-feedback control of the hypothalamus–pituitary–thyroid (HPT) axis in autoimmune (Hashimoto’s) thyroiditis. Mathematical Medicine and Biology, 31(3):226–258, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/3/226/710017.

Perrillat-Mercerot:2021:PDM

[PMMA^{+}21] Angélique Perrillat-Mercerot, Alain Miranville, Abramo Agosti, Elisabetta Rocca, Pasquale Ciarletta, and Rémy Guillevin. Partial differential model of lactate neuro-energetics: analytic results and numerical simulations. Mathematical Medicine and Biology, 38(2):178–201, June 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/38/2/178/6105780.

Pettet:2000:LVE

[PMN00] G. J. Pettet, D. L. S. Mcelwain, and J. Norbury. Lotka–Volterra equations with chemotaxis: walls, barriers and travelling waves. Mathematical Medicine and Biology, 17(4):395–413, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/4/395/664754.

Pybus:2023:DMT

[POB23] Hannah J. Pybus, Reuben D. O’Dea, and Bindi S. Brook. A dynamical model of TGF-β activation in asthmatic airways. Mathematical Medicine and Biology, 40(3):238–265, September 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/40/3/238/7191675.

Porco:1999:MME

[Por99] Travis C. Porco. A mathematical model of the ecology of Lyme disease. Mathematical Medicine and Biology, 16(3):261– 296, September 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/3/261/2874289.

PoznaNski:1988:MVC

[Poz88] R. R. Pozna’Nski. Membrane voltage changes in passive dendritic trees: a tapering equivalent cylinder model. Mathematical Medicine and Biology, 5(2):113–145, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/5/2/113/
731015.

Poznanski:1990:APS

[Poz90] Roman R. Poznański. Analysis of a postsynaptic scheme based on a tapering equivalent cable model. Mathematical Medicine and Biology, 7(3):175–197, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/7/3/175/733179.

Pozrikidis:2005:RSS

[Poz05] C. Pozrikidis. Resting shape and spontaneous membrane curvature of red blood cells. Mathematical Medicine and Biology, 22(1):34–52, March 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/22/1/34/657680.

PoznaNski:1996:SRW

[PP96] Roman R. Pozna’Nski and M. Shelton Peiris. Subthreshold response to white-noise current input in a tapering cable model of a neuron. Mathematical Medicine and Biology, 13(3):207– 222, September 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/13/3/207/651316.

Portillo:2021:MMA

[PP21] A. M. Portillo and C. Peláez. Mathematical modelling of ageing acceleration of the human follicle due to oxidative stress and other factors. Mathematical Medicine and Biology, 38(3):273– 291, September 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/38/3/273/6200054.

Paul:2016:SSD

[PR16] Subhadip Paul and Prasun Kumar Roy. Strategy for stochastic dose-rate induced enhanced elimination of malignant tumour without dose escalation. Mathematical Medicine and Biology, 33(3):319–328, September 2016. CODEN MMBABK.

ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/33/3/319/1749988.

Puwal:2005:APS

[PRK05] Steffan Puwal, Bradley J. Roth, and Serge Kruk. Automating phase singularity localization in mathematical models of cardiac tissue dynamics. Mathematical Medicine and Biology, 22 (4):335–346, December 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/22/4/335/650310.

Pop:2007:SFN

[PRWJ07] S. R. Pop, G. Richardson, S. L. Waters, and O. E. Jensen. Shock formation and non-linear dispersion in a microvascular capillary network. Mathematical Medicine and Biology, 24(4): 379–400, December 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/24/4/379/661378.

Plank:2003:RRW

[PS03] M. J. Plank and B. D. Sleeman. A reinforced random walk model of tumour angiogenesis and anti-angiogenic strategies. Mathematical Medicine and Biology, 20(2):135–181, June 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/2/135/658769.

Perelson:1985:MSC

[PSS85] Alan S. Perelson, Lee A. Segel, and John L. Spouge. On a model for the structure of circular mitochondrial genomes in higher plants. Mathematical Medicine and Biology, 2(1):41–56, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/2/1/41/815662.

Pearson:2014:MMI

[PSWO14] Natalie C. Pearson, Rebecca J. Shipley, Sarah L. Waters, and James M. Oliver. Multiphase modelling of the influence of fluid flow and chemical concentration on tissue growth in a hollow fibre membrane bioreactor. Mathematical Medicine and Biology, 31(4):393–430, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/31/4/393/646897.

Pstras:2017:MMC

[PTW^{+}17] Leszek Pstras, Karl Thomaseth, Jacek Waniewski, Italo Balzani, and Federico Bellavere. Mathematical modelling of cardiovascular response to the Valsalva manoeuvre. Mathematical Medicine and Biology, 34(2):261–292, June 2017. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
34/2/261/2885298.

Pugliese:1991:CMM

[Pug91] Andrea Pugliese. Contact matrices for multipopulation epidemic models: How to build a consistent matrix close to data. Mathematical Medicine and Biology, 8(4):249–271, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
8/4/249/666754.

Pielaat:1998:MDP

[PvdB98] A. Pielaat and F. van den Bosch. A model for dispersal of plant pathogens by rainsplash. Mathematical Medicine and Biology, 15(2):117–134, June 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/15/2/117/656869.

Pei:2009:PMP

[PYLC09] Yongzhen Pei, Yong Yang, Changguo Li, and Lansun Chen. Pest management of a prey-predator model with sexual favoritism. Mathematical Medicine and Biology, 26(2):97–115, June 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/26/2/97/727668.

Rejniak:2012:SAC

[RA12] Katarzyna A. Rejniak and Alexander R. A. Anderson. State of the art in computational modelling of cancer. Mathematical Medicine and Biology, 29(1):1–2, March 2012. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/1/1/753905.

Rosenberg:1987:HIS

[RB87] A. A. Rosenberg and J. R. Beddington. Harvesting interacting species of different lifespans. Mathematical Medicine and Biology, 4(3):265–278, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/4/3/265/874292.

Ross:2015:CTA

[RB15] Joshua V. Ross and Andrew J. Black. Contact tracing and antiviral prophylaxis in the early stages of a pandemic: the probability of a major outbreak. Mathematical Medicine and Biology, 32(3):331–343, September 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/32/3/331/646320.

Ruiz-Baier:2014:MMA

[RBGR^{+}14] Ricardo Ruiz-Baier, Alessio Gizzi, Simone Rossi, Christian Cherubini, Aymen Laadhari, Simonetta Filippi, and Alfio Quarteroni. Mathematical modelling of active contraction in isolated cardiomyocytes. Mathematical Medicine and Biology, 31(3):259–283, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/3/259/710028.

Ruck:2017:IMM

[RBMH17] T. Ruck, S. Bittner, S. G. Meuth, and M. Herty. Insights from mathematical modelling for T cell migration into the central nervous system. Mathematical Medicine and Biology, 34 (1):39–58, March 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/34/1/39/2885287.

Reed:1988:OHF

[Ree88] William J. Reed. Optimal harvesting of a fishery subject to random catastrophic collapse. Mathematical Medicine and Biology, 5(3):215–235, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/5/3/215/692266.

Reed:1997:EHF

[Ree97] William J. Reed. Estimating historical forest-fire frequencies from time-since-last-fire-sample data. Mathematical Medicine and Biology, 14(1):71–83, March 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/14/1/71/660012.

Reluga:2005:SMA

[Rel05] Timothy C. Reluga. Simple models of antibiotic cycling. Mathematical Medicine and Biology, 22(2):187–208, June 2005. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
22/2/187/770954.

Renshaw:1985:CSS

[Ren85] Eric Renshaw. Computer simulation of Sitka spruce: Spatial branching models for canopy growth and root structure. Mathematical Medicine and Biology, 2(3):183–200, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
2/3/183/2871037.

Renshaw:1994:CB

[Ren94] Eric Renshaw. Chaos in biometry. Mathematical Medicine and Biology, 11(1):17–44, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/11/1/17/667153.

Renshaw:1998:SAS

[Ren98] Eric Renshaw. Saddlepoint approximations for stochastic processes with truncated cumulant generating functions. Mathematical Medicine and Biology, 15(1):41–52, March 1998. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
15/1/41/683938.

Roberts:1991:PDN

[RG91] M. G. Roberts and B. T. Grenfell. The population dynamics of nematode infections of ruminants: Periodic perturbations as a model for management. Mathematical Medicine and Biology, 8(2):83–93, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/8/2/83/698767.

Roberts:1992:PDN

[RG92] M. G. Roberts and B. T. Grenfell. The population dynamics of nematode infections of ruminants: The effect of seasonally in the free-living stages. Mathematical Medicine and Biology, 9(1):29–41, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/9/1/29/664825.

Ruxton:1995:EGF

[RG95] G. D. Ruxton and C. A. Glasbey. Energetics of group foraging: Analysis of a random-walk model. Mathematical Medicine and Biology, 12(1):71–81, ???? 1995. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/12/1/71/707678.

Richardson:2009:MAM

[Ric09] G. Richardson. A multiscale approach to modelling electrochemical processes occurring across the cell membrane with application to transmission of action potentials. Mathematical Medicine and Biology, 26(3):201–224, September 2009. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
26/3/201/808420.

Rinaldi:1990:GSR

[Rin90] F. Rinaldi. Global stability results for epidemic models with latent period. Mathematical Medicine and Biology, 7(2):69–75, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/7/2/69/698795.

Roberts:1996:SMD

[RJ96] M. G. Roberts and J. Jowett. An SEI model with densitydependent demographics and epidemiology. Mathematical Medicine and Biology, 13(4):245–257, December 1996. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/4/245/771414.

Robertson:2001:PFW

[RK01] Malcolm B. Robertson and Uwe Köhler. Physiological flow waveform in a rigid elliptical vessel. Mathematical Medicine and Biology, 18(1):77–98, March 1, 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/18/1/77/727376.

Ruxton:2002:SIE

[RKAL02] Graeme D. Ruxton, Q. J. A. Khan, and Mohamed Al-Lawatia. The stability of internal equilibria in predator–prey models with breeding suppression. Mathematical Medicine and Biology, 19 (3):207–219, September 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/19/3/207/729871.

Ryu:2013:ETI

[RL13] Hwayeon Ryu and Anita T. Layton. Effect of tubular inhomogeneities on feedback-mediated dynamics of a model of a thick ascending limb. Mathematical Medicine and Biology, 30 (3):191–212, September 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/30/3/191/684865.

Rlngwood:1985:SPA

[Rln85] G. A. Rlngwood. The structure of Poisson algebras. Mathematical Medicine and Biology, 2(1):69–73, ???? 1985. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
2/1/69/815677.

Romero-Mendez:2010:ASP

[RMJLSG10] Ricardo Romero-Méndez, Joel N. Jiménez-Lozano, Mihir Sen, and F. Javier González. Analytical solution of the Pennes equation for burn-depth determination from infrared thermographs. Mathematical Medicine and Biology, 27(1):21–38, March 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/1/21/731492.

Rohani:2004:CNS

[RMK04] Pejman Rohani, Octavio Miramontes, and Matt J. Keeling. The colour of noise in short ecological time series data. Mathematical Medicine and Biology, 21(1):63–72, March 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/21/1/63/
689398.

Ribassin-Majed:2014:DMT

[RMLC14] Laureen Ribassin-Majed, Rachid Lounes, and Stephan Clemençon. Deterministic modelling for transmission of Human Papillomavirus 6/11: impact of vaccination. Mathematical Medicine and Biology, 31(2):125–149, June 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/2/125/660580.

Ratti:2020:MMH

[RNE^{+}20] Vardayani Ratti, Seema Nanda, Susan K. Eszterhas, Alexandra L. Howell, and Dorothy I. Wallace. A mathematical model of HIV dynamics treated with a population of geneedited haematopoietic progenitor cells exhibiting threshold phenomenon. Mathematical Medicine and Biology, 37(2): 212–242, June 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/37/2/212/5526927.

Roberts:1992:DCB

[Rob92] M. G. Roberts. The dynamics and control of bovine tuberculosis in possums. Mathematical Medicine and Biology, 9(1):19–28, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/9/1/19/664819.

Roberts:1999:KMM

[Rob99] M. G. Roberts. A Kermack-McKendrick model applied to an infectious disease in a natural population. Mathematical Medicine and Biology, 16(4):319–332, December 1999. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
16/4/319/652975.

Rodriguez-Perez:2007:ACI

[RPSGR^{+}07] D. Rodríguez-Pérez, Oscar Sotolongo-Grau, Ramón Espinosa Riquelme, Oscar Sotolongo-Costa, J. Antonio Santos Miranda, and J. C. Antoranz. Assessment of cancer immunotherapy outcome in terms of the immune response time features. Mathematical Medicine and Biology, 24(3):287–300, September 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/3/287/712089.

Rejniak:2012:CII

[RQA12] Katarzyna A. Rejniak, Vito Quaranta, and Alexander R. A. Anderson. Computational investigation of intrinsic and extrinsic mechanisms underlying the formation of carcinoma. Mathematical Medicine and Biology, 29(1):67–84, March 2012. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/1/67/753892.

Rohani:1999:DSM

[RR99] Pejman Rohani and Graeme D. Ruxton. Dispersal and stability in metapopulations. Mathematical Medicine and Biology, 16(3): 297–306, September 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/16/3/297/657151.

Reuben:1990:SPM

[RS90] A. J. Reuben and A. G. Shannon. Some problems in the mathematical modelling of erythrocyte sedimentation. Mathematical Medicine and Biology, 7(3):145–156, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/7/3/145/
733162.

Remien:2014:MMC

[RSA14] Christopher H. Remien, Norman L. Sussman, and Frederick R. Adler. Mathematical modelling of chronic acetaminophen metabolism and liver injury. Mathematical Medicine and Biology, 31(3):302–317, September 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/31/3/302/710035.

Roberts:2001:UML

[RT01] M. G. Roberts and M. I. Tobias. The use of multistate lifetable models for improving population health. Mathematical Medicine and Biology, 18(2):119–130, June 2001. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
18/2/119/672231.

Rivas:2021:CPR

[RTG21] Samantha R. Rivas, Alex C. Tessner, and Eli E. Goldwyn. Calculating prescription rates and addiction probabilities for the four most commonly prescribed opioids and evaluating their impact on addiction using compartment modelling. Mathematical Medicine and Biology, 38(2):202–217, June 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/2/202/6133992.

Ridenti:2023:MMI

[RTMT23] Marco A. Ridenti, Lara K. Teles, Alexandre Maranhão, and Vladimir K. Teles. Mathematical modeling and investigation on the role of demography and contact patterns in social distancing measures effectiveness in COVID-19 dissemination. Mathematical Medicine and Biology, 40(1):73–95, March 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/40/1/73/
6824420.

Ruan:2001:ZTD

[RW01] Shigui Ruan and Junjie Wei. On the zeros of a third degree exponential polynomial with applications to a delayed model for the control of testosterone secretion. Mathematical Medicine and Biology, 18(1):41–52, March 1, 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/18/1/41/727370.

Rubio:2022:MMD

[RY22] Felipe Alves Rubio and Hyun Mo Yang. A mathematical model to describe antibody-dependent enhancement and assess the effect of limiting cloning for plasma cells in heterologous secondary dengue infection. Mathematical Medicine and Biology, 39(2):169–196, June 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/39/2/169/6563011.

Robbins:2007:ACA

[RZBR07] K. R. Robbins, W. Zhang, J. K. Bertrand, and R. Rekaya. The ant colony algorithm for feature selection in high-dimension gene expression data for disease classification. Mathematical Medicine and Biology, 24(4):413–426, December 2007. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
24/4/413/661381.

Swinton:2002:MMA

[SA02] Jonathan Swinton and Bill Amos. Is microsatellite mutational asymmetry detectable in allele frequency distributions? Mathematical Medicine and Biology, 19(4):257–273, December 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
19/4/257/753205.

Solana-Arellano:1998:SAP

[SAEHLR98] Elena Solana-Arellano, Hector Echaverría-Heras, and Cecilia Leal-Ramírez. Surface aggregation patterns of LDL receptors near coated pits I. The radially convective diffusion and generalized insertion mechanism. Mathematical Medicine and Biology, 15(4):351–366, December 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/15/4/351/678309.

Solana-Arellano:2003:ILA

[SAEHM03] Elena Solana-Arellano, Hector Echavarria-Heras, and Margarita Gallegos Martinez. Improved leaf area index based biomass estimations for Zostera marina L. Mathematical Medicine and Biology, 20(4):367–375, December 2003. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/4/367/714034.

Schley:2004:MVH

[SBM04] D. Schley, J. Billingham, and R. J. Marchbanks. A model of in-vivo hydrocephalus shunt dynamics for blockage and performance diagnostics. Mathematical Medicine and Biology, 21(4): 347–368, December 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/21/4/347/741361.

Schenzle:1984:ASM

[Sch84] Dieter Schenzle. An age-structured model of pre- and postvaccination measles transmission. Mathematical Medicine and Biology, 1(2):169–191, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/1/2/169/643666.

Schaffer:1985:CND

[Sch85a] W. M. Schaffer. Can nonlinear dynamics elucidate mechanisms in ecology and epidemiology? Mathematical Medicine and Biology, 2(4):221–252, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/2/4/221/2268102.

Schumacher:1985:MCL

[Sch85b] Konrad Schumacher. Mathematical considerations on the larchlarch-bud-moth hypothesis. Mathematical Medicine and Biology, 2(4):253–285, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/2/4/253/833985.

Simitev:2023:PAS

[SDA^{+}23] Radostin D. Simitev, Antesar Al Dawoud, Muhamad H. N. Aziz, Rachel Myles, and Godfrey L. Smith. Phenomenological analysis of simple ion channel block in large populations of uncoupled cardiomyocytes. Mathematical Medicine and Biology, 40(2):175–198, June 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/40/2/175/6998547.

Secomb:2016:GFM

[Sec16] Timothy W. Secomb. A Green’s function method for simulation of time-dependent solute transport and reaction in realistic microvascular geometries. Mathematical Medicine and Biology, 33(4):475–494, December 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/33/4/475/2674338.

Segel:2003:BRJ

[Seg03] Lee A. Segel. Book review: J. D. Murray: Mathematical Biology (3rd Ed), Volume I (An Introduction) and Volume II (Spatial Models and Biomedical Applications). Mathematical Medicine and Biology, 20(4):377–378, December 2003. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/4/377/714036.

Sgouralis:2017:RMU

[SEL17] Ioannis Sgouralis, Roger G. Evans, and Anita T. Layton. Renal medullary and urinary oxygen tension during cardiopulmonary bypass in the rat. Mathematical Medicine and Biology, 34(3): 313–333, September 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/34/3/313/2885300.

Seymour:1990:VLP

[Sey90] R. M. Seymour. Very long period cycles in a near-optimal model of the population dynamics of Acanthaster Planci. Mathematical Medicine and Biology, 7(3):157–174, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/7/3/157/
733172.

Seymour:1995:SAC

[Sey95] R. M. Seymour. Some aspects of the coevolution of virulence and resistance in contact transmission disease processes with ecological constraints. Mathematical Medicine and Biology, 12 (2):83–136, June 1995. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/12/2/83/666450.

Strychalski:2013:CMB

[SG13] Wanda Strychalski and Robert D. Guy. A computational model of bleb formation. Mathematical Medicine and Biology, 30(2): 115–130, June 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/2/115/890410.

Schatzmann:1986:MNO

[SGB86] E. Schatzmann, R. Gerrard, and A. D. Barbour. Measures of niche overlap, I. Mathematical Medicine and Biology, 3(2): 99–113, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/2/99/752748.

Swinton:1992:GII

[SGBA92] Jonathan Swinton, Geoffrey P. Garnett, Robert C. Brunham, and Roy M. Anderson. Gonococcal infection, infertility, and population growth: I. Endemic states in behaviourally homogeneous growing populations. Mathematical Medicine and Biology, 9(2):107–126, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/9/2/107/731772.

Sepulveda-Galvez:2018:FTP

[SGBCC18] Alfonso Sepúlveda-Gálvez, Jesús Agustín Badillo-Corona, and Isaac Chairez. Finite-time parametric identification for the model representing the metabolic and genetic regulatory effects of sequential aerobic respiration and anaerobic fermentation processes in Escherichia coli. Mathematical Medicine and Biology, 35(3):299–317, September 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/3/299/3074186.

Saberian:2016:OFR

[SGK16] Fatemeh Saberian, Archis Ghate, and Minsun Kim. Optimal fractionation in radiotherapy with multiple normal tissues. Mathematical Medicine and Biology, 33(2):211–252, June 1, 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/33/2/211/2875347.

Sotolongo-Grau:2009:IST

[SGRPSM^{+}09] O. Sotolongo-Grau, D. Rodríguez-Pérez, J. A. Santos-Miranda, O. Sotolongo-Costa, and J. C. Antoranz. Immune systemtumour efficiency ratio as a new oncological index for radiotherapy treatment optimization. Mathematical Medicine and Biology, 26(4):297–307, December 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/26/4/297/654362.

Smith:2019:AEI

[SGZ19] Sean G. Smith, Boyce E. Griffith, and David A. Zaharoff. Analyzing the effects of instillation volume on intravesical delivery using biphasic solute transport in a deformable geometry. Mathematical Medicine and Biology, 36(2):139–156, June 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
36/2/139/4961298.

Seymour:2001:PIA

[SH01] R. M. Seymour and B. Henderson. Pro-inflammatory-anti-inflammatory cytokine dynamics mediated by cytokinereceptor dynamics in monocytes. Mathematical Medicine and Biology, 18(2):159–192, June 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/18/2/159/672236.

Sheehan:1992:SGC

[She92] Nuala Sheehan. Sampling genotypes on complex pedigrees with phenotypic constraints: The origin of the B allele among the Polar Eskimos. Mathematical Medicine and Biology, 9(1):1–18, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/9/1/1/664815.

Stocks:2017:SMN

[SHGB17] Theresa Stocks, Thomas Hillen, Jiafen Gong, and Martin Burger. A stochastic model for the normal tissue complication probability (NTCP) and applications. Mathematical Medicine and Biology, 34(4):469–492, December 2017. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
34/4/469/2885303.

Shannon:1994:MMI

[SHO^{+}94] A. G. Shannon, J. M. Hogg, R. L. Ollerton, S. Luzio, and D. R. Owens. A mathematical model of insulin secretion. Mathematical Medicine and Biology, 11(4):245–266, ???? 1994. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
11/4/245/649670.

Smith:2003:AMM

[SJ03] F. T. Smith and M. A. Jones. AVM modelling by multibranching tube flow: large flow rates and dual solutions. Mathematical Medicine and Biology, 20(2):183–204, June 2003. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/2/183/658771.

Shaikhet:2024:APL

[SK24] Leonid Shaikhet and Andrei Korobeinikov. Asymptotic properties of the Lotka–Volterra competition and mutualism model under stochastic perturbations. Mathematical Medicine and Biology, 41(1):19–34, March 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/41/1/19/7591973.

Sazonov:2011:TWN

[SKG11] Igor Sazonov, Mark Kelbert, and Michael B. Gravenor. Travelling waves in a network of SIR epidemic nodes with an approximation of weak coupling. Mathematical Medicine and Biology, 28(2):165–183, June 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/28/2/165/679769.

Schaffer:1993:TPE

[SKTO93] W. M. Schaffer, B. E. Kendall, C. W. Tidd, and L. F. Olsen. Transient periodicity and episodic predictability in biological dynamics. Mathematical Medicine and Biology, 10(4):227–247, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/10/4/227/678649.

Sgouralis:2016:CFM

[SL16] Ioannis Sgouralis and Anita T. Layton. Conduction of feedback-mediated signal in a computational model of coupled nephrons. Mathematical Medicine and Biology, 33(1): 87–106, March 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/33/1/87/2363532.

Sleeman:1988:PDB

[Sle88] B. D. Sleeman. Period-doubling bifurcations leading to chaos in discrete models of biology. Mathematical Medicine and Biology, 5(1):21–31, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/5/1/21/763970.

Shakeel:2013:CMC

[SMGW13] Muhammad Shakeel, Paul C. Matthews, Richard S. Graham, and Sarah L. Waters. A continuum model of cell proliferation and nutrient transport in a perfusion bioreactor. Mathematical Medicine and Biology, 30(1):21–44, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/30/1/21/
710589.

Smith:2020:AOR

[Smi20] Gregory Douglas Conradi Smith. Allostery in oligomeric receptor models. Mathematical Medicine and Biology, 37(3):313– 333, September 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/37/3/313/5666166.

Sherratt:1992:MMW

[SMML92] Jonathan A. Sherratt, Paul Martin, J. D. Murray, and Julian Lewis. Mathematical models of wound healing in embryonic and adult epidermis. Mathematical Medicine and Biology, 9 (3):177–196, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/9/3/177/649550.

Sleeman:1998:FTV

[SN98] B. D. Sleeman and H. R. Nlmmo. Fluid transport in vascularized tumours and metastasis. Mathematical Medicine and Biology, 15(1):53–63, March 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/15/1/53/683940.

Shafer:2011:SMS

[SNB^{+}11] Ilari Shafer, Rachel Nancollas, Morgan Boes, Alisha L. Sieminski, and John B. Geddes. Stability of a microvessel subject to structural adaptation of diameter and wall thickness. Mathematical Medicine and Biology, 28(3):271–286, September 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/28/3/271/664321.

Sneyd:2003:CCB

[Sne03] James Sneyd. Computational cell biology. Mathematical Medicine and Biology, 20(1):131–133, March 2003. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/1/131/663534.

Showa:2016:EBA

[SNHMM16] S. P. Showa, F. Nyabadza, S. D. Hove-Musekwa, and G. Magombedze. Exploring the benefits of antibody immune response in HIV-1 infection using a discrete model. Mathematical Medicine and Biology, 33(2):189–210, June 1, 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
33/2/189/2875346.

Stevens:2007:IIH

[SPL^{+}07] Scott A. Stevens, Michelle Previte, William D. Lakin, Nimish J. Thakore, Paul L. Penar, and Brandon Hamschin. Idiopathic intracranial hypertension and transverse sinus stenosis: a modelling study. Mathematical Medicine and Biology, 24(1):85– 109, March 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/24/1/85/692200.

Sonnino:2022:MSS

[SPN22] Giorgio Sonnino, Philippe Peeters, and Pasquale Nardone. Modelling the spreading of the SARS-CoV-2 in presence of the lockdown and quarantine measures by a kinetic-type reactions approach. Mathematical Medicine and Biology, 39(2): 105–125, June 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/39/2/105/6454152.

Stefanini:2012:CMV

[SQGP12] Marianne O. Stefanini, Amina A. Qutub, Feilim Mac Gabhann, and Aleksander S. Popel. Computational models of VEGF-associated angiogenic processes in cancer. Mathematical Medicine and Biology, 29(1):85–94, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/29/1/85/
753898.

Singh:1989:DMF

[SSA89] M. P. Singh, Maithili Sharan, and A. Aminataei. Development of mathematical formulae for O_{2} and CO_{2} dissociation curves in the blood. Mathematical Medicine and Biology, 6(1):25–46, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/6/1/25/666284.

Singh:1988:PGE

[SSS88] M. P. Singh, Maithili Sharan, and Indu Sud. The process of gas exchange in systemic circulation in a hyperbaric environment: an analytical approach. Mathematical Medicine and Biology, 5 (4):281–301, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/5/4/281/700655.

Sharan:1991:MMR

[SSS91] Maithili Sharan, M. P. Singh, and Balbir Singh. A mathematical model for the rate of oxygenation of blood in pulmonary capillaries using n th-order one-step kinetics of oxygen uptake by haemoglobin. Mathematical Medicine and Biology, 8(2): 125–140, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/8/2/125/698755.

Shipley:2020:HDC

[SSS^{+}20] Rebecca J. Shipley, Amy F. Smith, Paul W. Sweeney, Axel R. Pries, and Timothy W. Secomb. A hybrid discrete-continuum approach for modelling microcirculatory blood flow. Mathematical Medicine and Biology, 37(1):40–57, March 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/37/1/40/
5393647.

Sharan:1991:FEA

[SSSK91] Maithili Sharan, Balbir Singh, M. P. Singh, and Pawan Kumar. Finite-element analysis of oxygen transport in the systemic capillaries. Mathematical Medicine and Biology, 8(2): 107–123, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/8/2/107/698748.

Smith:1985:TTD

[ST85] P. Smith and E. H. Twizell. A transient two-dimensional model of thermoregulation in a human subject. Mathematical Medicine and Biology, 2(3):161–181, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/2/3/161/
682025.

Szczesny:1989:GER

[ST89] Stanisław Szczesny and Boleslaw Turczyński. The generalized equation for the radiotracer clearance method of measurement of regional microcirculation. Mathematical Medicine and Biology, 6(4):257–267, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/6/4/257/794038.

Stocker:1998:BME

[SW98] Sabine Stöcker and Daniel Weihs. Bird migration — an energybased analysis of costs and benefits. Mathematical Medicine and Biology, 15(1):65–85, March 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/15/1/65/683943.

Sobey:2006:ENL

[SW06] Ian Sobey and Benedikt Wirth. Effect of non-linear permeability in a spherically symmetric model of hydrocephalus. Mathematical Medicine and Biology, 23(4):339–361, December 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/4/339/689787.

Shipley:2012:FMT

[SW12] Rebecca J. Shipley and Sarah L. Waters. Fluid and mass transport modelling to drive the design of cell-packed hollow fibre bioreactors for tissue engineering applications. Mathematical Medicine and Biology, 29(4):329–359, December 2012. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/4/329/821011.

Swan:1985:OCA

[Swa85] George W. Swan. Optimal control applications in the chemotherapy of multiple myeloma. Mathematical Medicine and Biology, 2(3):139–160, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/2/3/139/682020.

Swan:1987:OCA

[Swa87] George W. Swan. Optimal control analysis of a cancer chemotherapy problem. Mathematical Medicine and Biology, 4 (2):171–184, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/4/2/171/873595.

Swan:1988:GAO

[Swa88] G. W. Swan. General applications of optimal control theory in cancer chemotherapy. Mathematical Medicine and Biology, 5 (4):303–316, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/5/4/303/700661.

Siggers:2009:FDS

[SWWC09] Jennifer H. Siggers, Sarah Waters, Jonathan Wattis, and Linda Cummings. Flow dynamics in a stented ureter. Mathematical Medicine and Biology, 26(1):1–24, March 2009. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
26/1/1/853980.

Scott:2021:MFM

[SZB21] M. Scott, K. Zychaluk, and R. N. Bearon. A mathematical framework for modelling 3D cell motility: applications to glioblastoma cell migration. Mathematical Medicine and Biology, 38(3):333–354, September 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/38/3/333/6311226.

Tam:1999:DEM

[Tam99] Judy Tam. Delay effect in a model for virus replication. Mathematical Medicine and Biology, 16(1):29–37, March 1999. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
16/1/29/729524.

Tang:1987:GMG

[Tan87] Betty Tang. Growth of microorganisms in gradient-like enviroments. Mathematical Medicine and Biology, 4(1):47–58, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/4/1/47/875668.

Thomas:2003:ESM

[TC03] Alun Thomas and Chris Cannings. Enumeration and simulation of marriage node graphs on zero-loop pedigrees. Mathematical Medicine and Biology, 20(3):261–275, September 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/3/261/714219.

Thomas:2004:SRZ

[TC04] Alun Thomas and Chris Cannings. Simulating realistic zero loop pedigrees using a bipartite Prüfer code and graphical modelling. Mathematical Medicine and Biology, 21(4):335– 345, December 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/21/4/335/741359.

Thomas:2006:SCP

[TC06] Peter J. Thomas and Jack D. Cowan. Simultaneous constraints on pre- and post-synaptic cells couple cortical feature maps in a 2D geometric model of orientation preference. Mathematical Medicine and Biology, 23(2):119–138, June 2006. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/2/119/671411.

Tchuenche:2011:MMA

[TCC^{+}11] Jean M. Tchuenche, Christinah Chiyaka, David Chan, Amanda Matthews, and Ghislaine Mayer. A mathematical model for antimalarial drug resistance. Mathematical Medicine and Biology, 28(4):335–355, December 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/28/4/335/658330.

Thompson:1991:ELR

[TG91] E. A. Thompson and Sun Wei Guo. Evaluation of likelihood ratios for complex genetic models. Mathematical Medicine and Biology, 8(3):149–169, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/8/3/149/689694.

Tavares:2023:CIM

[TG23] J. N. Tavares and Emilie Gomes. COVID-19 immunotherapy: a mathematical model. Mathematical Medicine and Biology, 40 (2):199–221, June 2023. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/40/2/199/7112767.

Trowbridge:1988:SMP

[TH88] E. A. Trowbridge and P. J. Harley. A stochastic model of pulmonary platelet production. Mathematical Medicine and Biology, 5(1):45–63, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/5/1/45/763999.

Thornton:1996:DSI

[TH96] B. S. Thornton and W. T. Hung. Dynamic stiffness and implications for assisting the operation of the left ventricle. Mathematical Medicine and Biology, 13(4):275–295, December 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/4/275/771421.

Thornton:1992:DML

[THH92] B. S. Thornton, W. T. Hung, and Cherrell Hirst. Diagnostic model for local temporal thermal change at the skin of the breast during extended application of diagnostic ultrasound. Mathematical Medicine and Biology, 9(3):161–175, ???? 1992. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
9/3/161/649542.

Thornton:1991:RDF

[THI91] B. S. Thornton, W. T. Hung, and J. Irving. Relaxation distribution function of intracellular dielectric zones as an indicator of tumorous transition of living cells. Mathematical Medicine and Biology, 8(2):95–106, ???? 1991. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/8/2/95/698772.

Thilagam:2016:MMD

[Thi16] Alagu Thilagam. Mathematical modelling of decline in follicle pool during female reproductive ageing. Mathematical Medicine and Biology, 33(1):107–121, March 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/33/1/107/2363548.

Thompson:1984:IGJ

[Tho84] E. A. Thompson. Information gain in joint linkage analysis. Mathematical Medicine and Biology, 1(1):31–49, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
1/1/31/754936.

Thompson:1985:HLX

[Tho85] E. A. Thompson. The history of a lethal X-Linked mutant. Mathematical Medicine and Biology, 2(2):75–92, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
2/2/75/765074.

Thomas:1986:ACP

[Tho86a] Alun Thomas. Approximate computation of probability functions for pedigree analysis. Mathematical Medicine and Biology, 3(3):157–166, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/3/157/716848.

Thomas:1986:OCP

[Tho86b] Alun Thomas. Optimal computation of probability functions for pedigree analysis. Mathematical Medicine and Biology, 3 (3):167–178, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/3/167/716868.

Thomas:1987:AMR

[Tho87a] G. E. Thomas. An approach to modelling radiation damage by fast ionizing particles. Mathematical Medicine and Biology, 4 (4):363–377, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/4/4/363/751290.

Thompson:1987:CCL

[Tho87b] E. A. Thompson. Crossover counts and likelihood in multipoint linkage analysis. Mathematical Medicine and Biology, 4 (2):93–108, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/4/2/93/873603.

Thomas:1988:DP

[Tho88a] Alun Thomas. Drawing pedigrees. Mathematical Medicine and Biology, 5(3):201–213, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/5/3/201/692263.

Thompson:1988:TLT

[Tho88b] E. A. Thompson. Two-locus and three-locus gene identity by descent in pedigrees. Mathematical Medicine and Biology, 5 (4):261–279, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/5/4/261/700649.

Thomas:1993:CPG

[Tho93] Alun Thomas. A class of perfect graphs in genetics. Mathematical Medicine and Biology, 10(2):77–81, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/10/2/77/
654151.

Thomas:1994:LAC

[Tho94] Alun Thomas. Linkage analysis on complex pedigrees by simulation. Mathematical Medicine and Biology, 11(2):79–93, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/11/2/79/771443.

Thomas:2010:CIB

[Tho10] Alun Thomas. The conditional independences between variables derived from two independent identically distributed Markov random fields when pairwise order is ignored. Mathematical Medicine and Biology, 27(3):283–288, September 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
27/3/283/791825.

Tyurin:2005:OPD

[TK05] Kirill V. Tyurin and Michael A. Khanin. Optimality principle and determination of kinetic constants for biochemical reactions. Mathematical Medicine and Biology, 22(1):1–14, March 2005. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/22/1/1/657675.

Taylor-King:2019:SAD

[TKBD^{+}19] Jake P. Taylor-King, Etienne Baratchart, Andrew Dhawan, Elizabeth A. Coker, Inga Hansine Rye, Hege Russnes, S. Jon Chapman, David Basanta, and Andriy Marusyk. Simulated ablation for detection of cells impacting paracrine signalling in histology analysis. Mathematical Medicine and Biology, 36(1): 93–112, March 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/36/1/93/4857315.

Thomas:2018:EMN

[TKR^{+}18] Alun Thomas, Karim Khader, Andrew Redd, Molly Leecaster, Yue Zhang, Makoto Jones, Tom Greene, and Matthew Samore. Extended models for nosocomial infection: parameter estimation and model selection. Mathematical Medicine and Biology, 35(S1):??, April 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL https://academic.oup.

com/imammb/article/35/Supplement_1/i29/4508809.

Taylor:1996:GPF

[TMM96] Gordon Taylor, Sally Mcclean, and Peter Millard. Geriatricpatient flow-rate modelling. Mathematical Medicine and Biology, 13(4):297–307, December 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/13/4/297/771427.

Trewenack:2009:CMD

[TPL09] Abbey J. Trewenack, Colin P. Please, and Kerry A. Landman. A continuum model for the development of tissue-engineered cartilage around a chondrocyte. Mathematical Medicine and Biology, 26(3):241–262, September 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/26/3/241/808446.

Tweedy:2018:FAC

[TPRS18] Jennifer H. Tweedy, Jan O. Pralits, Rodolfo Repetto, and Paolo Soleri. Flow in the anterior chamber of the eye with an implanted iris-fixated artificial lens. Mathematical Medicine and Biology, 35(3):363–385, September 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/3/363/4071238.

Tindall:2008:UPO

[TPS^{+}08] M. J. Tindall, M. A. Peletier, N. M. W. Severens, D. J. Veldman, and B. A. J. M. de Mol. Understanding post-operative temperature drop in cardiac surgery: a mathematical model. Mathematical Medicine and Biology, 25(4):323–335, December 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/25/4/323/826460.

Thomas:2015:EPE

[TRK^{+}15] Alun Thomas, Andrew Redd, Karim Khader, Molly Leecaster, Tom Greene, and Matthew Samore. Efficient parameter estimation for models of healthcare-associated pathogen transmission in discrete and continuous time. Mathematical Medicine and Biology, 32(1):81–100, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/32/1/81/661586.

Thompson:2019:IDG

[TRK19] Travis B. Thompson, Beatrice M. Riviere, and Matthew G. Knepley. An implicit discontinuous Galerkin method for modeling acute edema and resuscitation in the small intestine. Mathematical Medicine and Biology, 36(4):513–548, December 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
36/4/513/5306244.

Trujillo:2009:ATO

[TRMB09] Macarena Trujillo, María J. Rivera, Juan A. López Molina, and Enrique J. Berjano. Analytical thermal-optic model for laser heating of biological tissue using the hyperbolic heat transfer equation. Mathematical Medicine and Biology, 26(3):187– 200, September 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/26/3/187/808425.

Thomas:1994:PMD

[TS94] Alun Thomas and Mark H. Skolnick. A probabilistic model for detecting coding regions in DNA sequences. Mathematical Medicine and Biology, 11(3):149–160, ???? 1994. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
11/3/149/708077.

Tanaka:2024:PIT

[TSE24] Martin L. Tanaka, David M. Saylor, and Robert M. Elder. Polymer-interface-tissue model to estimate leachable release from medical devices. Mathematical Medicine and Biology, 41 (4):382–403, December 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/41/4/382/7825954.

Tsiliyannis:2024:BSM

[Tsi24] Christos Tsiliyannis. Beyond SIRD models: a novel dynamic model for epidemics, relating infected with entries to health care units and application for identification and restraining policy. Mathematical Medicine and Biology, 41(3):192– 224, September 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/41/3/192/7735399.

Thomas:1994:GMS

[TSL94] Alun Thomas, Mark H. Skolnick, and Cathryn M. Lewis. Genomic mismatch scanning in pedigrees. Mathematical Medicine and Biology, 11(1):1–16, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/11/1/1/667150.

Teytsa:2021:BAP

[TTBL21] H. M. Ndongmo Teytsa, B. Tsanou, S. Bowong, and J. M-S. Lubuma. Bifurcation analysis of a phage-bacteria interaction model with prophage induction. Mathematical Medicine and Biology, 38(1):28–58, March 2021. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/38/1/28/5877074.

Tuckwell:2000:NEE

[TW00] Henry C. Tuckwell and Frederic Y. M. Wan. Nature of equilibria and effects of drug treatments in some simple viral population dynamical models. Mathematical Medicine and Biology, 17 (4):311–327, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/17/4/311/664716.

Talbott:2015:MET

[TXAS15] Kevin Talbott, Amber Xu, Daniel M. Anderson, and Padmanabhan Seshaiyer. Modelling the evaporation of a tear film over a contact lens. Mathematical Medicine and Biology, 32 (2):209–238, June 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/2/209/830146.

Uche:1996:MMN

[UA96] Chinma O. Uche and Roy M. Anderson. Mixing matrices: Necessary constraints in populations of finite size. Mathematical Medicine and Biology, 13(1):23–33, March 1996. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/13/1/23/
644508.

Usher:1996:SDR

[UH96] J. R. Usher and D. Henderson. Some drug-resistant models for cancer chemotherapy. Part 1: Cycle-nonspecific drugs. Mathematical Medicine and Biology, 13(2):99–126, June 1996. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/2/99/655291.

Ushakova:2015:BEH

[UPM^{+}15] Anastasia Ushakova, Frank Olav Pettersen, Arild Mæland, Bo Henry Lindqvist, and Dag Kvale. Bayesian estimation of HIV-1 dynamics in vivo. Mathematical Medicine and Biology, 32(1):39–56, March 2015. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/32/1/39/661583.

vandenBosch:1986:IBE

[vdBD86] F. van den Bosch and O. Diekmann. Interactions between eggeating predator and prey: The effect of the functional response and of age structure. Mathematical Medicine and Biology, 3 (1):53–69, ???? 1986. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/3/1/53/684222.

vandenBoogaard:1984:MEN

[vdBJ84] Henk F. P. van den Boogaard and Peter I. M. Johannesma. The master equation for neural interaction. Mathematical Medicine and Biology, 1(4):365–389, ???? 1984. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/1/4/365/768085.

Vargas:2001:GAU

[VdC01] J. A. Vargas and Rafael F. del Castillo. Genetic associations under mixed mating systems: the Bennett-Binet effect. Mathematical Medicine and Biology, 18(4):327–341, December 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
18/4/327/747270.

Vargas-De-Leoon:2013:GSP

[VDLK13] Cruz Vargas-De-Leóon and Andrei Korobeinikov. Global stability of a population dynamics model with inhibition and negative feedback. Mathematical Medicine and Biology, 30(1): 65–72, March 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/1/65/710607.

Venema:1993:EPB

[Ven93] Henk W. Venema. Estimation of the parameters of a binary Markov random field on a graph with application to fibre type distributions in a muscle cross-section. Mathematical Medicine and Biology, 10(2):115–133, ???? 1993. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/10/2/115/654143.

Venturino:2002:EPP

[Ven02] Ezio Venturino. Epidemics in predator–prey models: disease in the predators. Mathematical Medicine and Biology, 19(3):185– 205, September 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/19/3/185/729868.

Velasco-Hernandez:1996:ETP

[VHBCC96] Jorge X. Velasco-Hernández, Fred Brauer, and Carlos Castillo-Chavez. Effects of treatment and prevalence-dependent recruitment on the dynamics of a fatal disease. Mathematical Medicine and Biology, 13(3):175–192, September 1996. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/3/175/651310.

Vaidya:2007:MPM

[VHT07] Naveen K. Vaidya, Huaxiong Huang, and Shu Takagi. Modelling HA protein-mediated interaction between an influenza virus and a healthy cell: pre-fusion membrane deformation. Mathematical Medicine and Biology, 24(3):251–270, September 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/24/3/251/712085.

Varadharajan:1994:SIT

[VJ94] Minal Varadharajan and Girija Jayaraman. Sodium ion transport in the intestinal wall: a mathematical model. Mathematical Medicine and Biology, 11(3):193–205, ???? 1994. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
11/3/193/708096.

Vaillant:2000:MCP

[VL00] Jean Vaillant and Petr Lánský. Multidimensional counting processes and evoked neuronal activity. Mathematical Medicine and Biology, 17(1):53–73, ???? 2000. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/17/1/53/674994.

Vo:2013:MPR

[VM13] Tuoi T. N. Vo and Martin G. Meere. Minimizing the passive release of heparin-binding growth factors from an affinity-based delivery system. Mathematical Medicine and Biology, 30(4): 357–382, December 2013. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/30/4/357/718687.

Vrdoljak:2024:SAA

[VV24] Anton Vrdoljak and Damir Vukicević. Selector of amino-acid scales set. Mathematical Medicine and Biology, 41(3):157– 168, September 2024. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/41/3/157/7709487.

Veliov:2016:AAA

[VW16] V. M. Veliov and A. Widder. Aggregation and asymptotic analysis of an SI-epidemic model for heterogeneous populations. Mathematical Medicine and Biology, 33(3):295–318, September 2016. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/33/3/295/1750241.

Wilkie:2020:MMI

[WA20] Kathleen P. Wilkie and Farjana Aktar. Mathematically modelling inflammation as a promoter of tumour growth. Mathematical Medicine and Biology, 37(4):491–514, December 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/4/491/5840464.

Winter:2010:MWE

[WAB10] Katlyn N. Winter, Daniel M. Anderson, and Richard J. Braun. A model for wetting and evaporation of a post-blink precorneal tear film. Mathematical Medicine and Biology, 27(3):211– 225, September 2010. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/27/3/211/791810.

Whiteley:2022:MAC

[WBG22] Jonathan P. Whiteley, Cameron P. Brown, and Eamonn A. Gaffney. Modelling articular cartilage: the relative motion of two adjacent poroviscoelastic layers. Mathematical Medicine and Biology, 39(3):251–298, September 2022. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
39/3/251/6604802.

Wood:1989:IME

[WBGN89] S. N. Wood, S. P. Blythe, W. S. C. Gurney, and R. M. Nisbet. Instability in mortality estimation schemes related to stagestructure population models. Mathematical Medicine and Biology, 6(1):47–68, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/6/1/47/666290.

Wang:2012:ITT

[WBSD12] Zhihui Wang, Veronika Bordas, Jonathan Sagotsky, and Thomas S. Deisboeck. Identifying therapeutic targets in a combined EGFR-TGFβR signalling cascade using a multiscale agent-based cancer model. Mathematical Medicine and Biology, 29(1):95–108, March 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/29/1/95/753895.

White:1998:CIV

[WCM98] L. J. White, M. J. Cox, and G. F. Medley. Cross immunity and vaccination against multiple microparasite strains. Mathematical Medicine and Biology, 15(3):211–233, September 1998. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
15/3/211/795634.

Waters:2006:TGR

[WCSR06] S. L. Waters, L. J. Cummings, K. M. Shakesheff, and F. R. A. J. Rose. Tissue growth in a rotating bioreactor. Part I: mechanical stability. Mathematical Medicine and Biology, 23 (4):311–337, December 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/23/4/311/689775.

Ward:2017:PTM

[WDD^{+}17] John P. Ward, Joanne L. Dunster, Gianne Derks, Pratibha Mistry, and José D. Salazar. Predicting tyrosinaemia: a mathematical model of 4-hydroxyphenylpyruvate dioxygenase inhibition by nitisinone in rats. Mathematical Medicine and Biology, 34 (3):335–390, September 2017. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/34/3/335/2885296.

Wilson:1988:SPM

[WDLS88] B. C. Wilson, D. Y. Downham, J. Lexell, and M. Sjöström. Some probability models for diagnosing neurogenic disorders. Mathematical Medicine and Biology, 5(3):167–179, ???? 1988. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
5/3/167/692250.

Wilkie:2012:AIB

[WDS12] K. P. Wilkie, C. S. Drapaca, and S. Sivaloganathan. Aging impact on brain biomechanics with applications to hydrocephalus. Mathematical Medicine and Biology, 29(2):145–161, June 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
29/2/145/662200.

Weir:1985:SAM

[Wei85] B. S. Weir. Statistical analysis of molecular genetic data. Mathematical Medicine and Biology, 2(1):1–39, ???? 1985. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
2/1/1/815647.

Whiteley:2003:ECG

[WG03] Jonathan P. Whiteley and David J. Gavaghan. Efficient computations of gas transport in the respiratory tree. Mathematical Medicine and Biology, 20(1):91–104, March 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/20/1/91/
663526.

Whiteley:2020:MIS

[WG20] Jonathan P. Whiteley and Eamonn A. Gaffney. Modelling the inclusion of swelling pressure in a tissue level poroviscoelastic model of cartilage deformation. Mathematical Medicine and Biology, 37(3):389–428, September 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/37/3/389/5739945.

Whiteley:2007:NLM

[WGCB07] Jonathan P. Whiteley, David J. Gavaghan, S. Jonathan Chapman, and J. Michael Brady. Non-linear modelling of breast tissue. Mathematical Medicine and Biology, 24(3):327–345, September 2007. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/24/3/327/712100.

Whiteley:2003:PBI

[WGH03] J. P. Whiteley, D. J. Gavaghan, and C. E. W. Hahn. Periodic breathing induced by arterial oxygen partial pressure oscillations. Mathematical Medicine and Biology, 20(2):205–224, June 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/20/2/205/658773.

Whiteley:2020:ESA

[Whi20] Jonathan P. Whiteley. An evaluation of some assumptions underpinning the bidomain equations of electrophysiology. Mathematical Medicine and Biology, 37(2):262–302, June 2020. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
37/2/262/5610879.

Wang:2012:GAS

[WHT12] Jinliang Wang, Gang Huang, and Yasuhiro Takeuchi. Global asymptotic stability for HIV-1 dynamics with two distributed delays. Mathematical Medicine and Biology, 29(3):283–300, September 2012. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/29/3/283/798008.

Ward:1997:MMA

[WK97] J. P. Ward and J. R. King. Mathematical modelling of avascular-tumour growth. Mathematical Medicine and Biology, 14(1):39–69, March 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/14/1/39/660000.

Ward:1999:MMA

[WK99] J. P. Ward and J. R. King. Mathematical modelling of avascular-tumour growth II: Modelling growth saturation. Mathematical Medicine and Biology, 16(2):171–211, June 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
16/2/171/668993.

Ward:2001:MMQ

[WKK^{+}01] J. P. Ward, J. R. King, A. J. Koerber, P. Williams, J. M. Croft, and R. E. Sockett. Mathematical modelling of quorum sensing in bacteria. Mathematical Medicine and Biology, 18(3):263– 292, September 1, 2001. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/18/3/263/657415.

Ward:2004:CSR

[WKK^{+}04] J. P. Ward, J. R. King, A. J. Koerber, J. M. Croft, R. E. Sockett, and P. Williams. Cell-signalling repression in bacterial quorum sensing. Mathematical Medicine and Biology, 21(3): 169–204, September 2004. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.

oup.com/imammb/article/21/3/169/660051.

Weeks:1989:TTT

[WL89] Daniel E. Weeks and Kenneth Lange. Trials, tribulations, and triumphs of the EM algorithm in pedigree analysis. Mathematical Medicine and Biology, 6(4):209–232, ???? 1989. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/6/4/209/
793982.

Wood:1994:SMB

[Woo94] S. N. Wood. Spline models of biological population dynamics: How to estimate mortality rates for stage structured populations with dimorphic life histories. Mathematical Medicine and Biology, 11(1):61–78, ???? 1994. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/11/1/61/667159.

Wall:1985:RSM

[WP85] C. Wall and J. N. Perry. The role of a simulation model in understanding the behaviour of male pea moths attracted to pheromone traps. Mathematical Medicine and Biology, 2(4): 299–309, ???? 1985. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/2/4/299/833992.

Waters:2018:EII

[WPH^{+}18] Ryan S. Waters, Justin S. A. Perry, SunPil Han, Bibiana Bielekova, and Tomas Gedeon. The effects of interleukin-2 on immune response regulation. Mathematical Medicine and Biology, 35(1):79–119, March 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/35/1/79/3038032.

Wirth:2006:AFP

[WS06] Benedikt Wirth and Ian Sobey. An axisymmetric and fully 3D poroelastic model for the evolution of hydrocephalus. Mathematical Medicine and Biology, 23(4):363–388, December 2006. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
23/4/363/689781.

Wirth:2008:MIP

[WS08] Benedikt Wirth and Ian Sobey. A model for an inverse power constitutive law for cerebral compliance. Mathematical Medicine and Biology, 25(2):113–131, June 2008. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
25/2/113/751920.

Wirth:2009:ASD

[WS09] Benedikt Wirth and Ian Sobey. Analytic solution during an infusion test of the linear unsteady poroelastic equations in a spherically symmetric model of the brain. Mathematical Medicine and Biology, 26(1):25–61, March 2009. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/26/1/25/
853986.

Watton:2009:MGS

[WVH09] Paul N. Watton, Yiannis Ventikos, and Gerhard A. Holzapfel. Modelling the growth and stabilization of cerebral aneurysms. Mathematical Medicine and Biology, 26(2):133–164, June 2009.

CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
26/2/133/727673.

White:2005:ABC

[WW05] Steven M. White and K. A. Jane White. Applications of biological control in resistant host-pathogen systems. Mathematical Medicine and Biology, 22(3):227–245, September 2005. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
22/3/227/652724.

Williams:2014:PSM

[WWWW^{+}14] Nakeya D. Williams, Øistein Wind-Willassen, Andrew A. Wright, Jesper Mehlsen, Johnny T. Ottesen, and Mette S. Olufsen. Patient-specific modelling of head-up tilt. Mathematical Medicine and Biology, 31(4):365–392, December 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
31/4/365/646887.

Xiaoming:1987:CSB

[XDJ87] Ding Xiaoming, Ding Dafu, and Xu Jinghua. Chiral symmetry breaking in living systems. Mathematical Medicine and Biology, 4(1):33–46, ???? 1987. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/4/1/33/875663.

Xiao:2011:MDS

[XZT11] Yanni Xiao, Yicang Zhou, and Sanyi Tang. Modelling disease spread in dispersal networks at two levels. Mathematical Medicine and Biology, 28(3):227–244, September 2011. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
28/3/227/664311.

Yamazaki:2018:GWP

[Yam18] Kazuo Yamazaki. Global well-posedness of infectious disease models without life-time immunity: the cases of cholera and avian influenza. Mathematical Medicine and Biology, 35(4): 427–445, December 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/35/4/427/4568333.

Yang:2002:ERI

[Yan02] Hyun Mo Yang. The effects of re-infection in directly transmitted infections modelled with vaccination. Mathematical Medicine and Biology, 19(2):113–135, June 2002. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
19/2/113/730721.

Yuan:2003:ETH

[YB03] Ao Yuan and George E. Bonney. Exact test of Hardy-Weinberg equilibrium by Markov chain Monte Carlo. Mathematical Medicine and Biology, 20(4):327–340, December 2003. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
20/4/327/714031.

Young:2018:ETG

[YB18] Todd R. Young and Erik M. Boczko. Early treatment gains for antibiotic administration and within human host time series data. Mathematical Medicine and Biology, 35(2):203–224, June 2018. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/35/2/203/3056187.

Yakowitz:1996:CME

[YBG96] S. Yakowitz, M. Blount, and J. Gani. Computing marginal expectations for large compartmentalized models with application to AIDS evolution in a prison system. Mathematical Medicine and Biology, 13(4):223–244, December 1996. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
13/4/223/771409.

Yang:1999:PMA

[YC99] Hyun Mo Yang and Wilson Castro Ferreira, Jr. A population model applied to HIV transmission considering protection and treatment. Mathematical Medicine and Biology, 16(3):237– 259, September 1999. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/16/3/237/657146.

Yates:2000:PGA

[YF00] Tony L. Yates and Leslie R. Fletcher. Prediction of a glucose appearance function from foods using deconvolution. Mathematical Medicine and Biology, 17(2):169–184, ???? 2000. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
17/2/169/736868.

Yang:2021:MTN

[YJY21] Hyun Mo Yang, Luis Pedro Lombardi Junior, and Ariana Campos Yang. Modeling the transmission of the new coronavirus in São Paulo State, Brazil — assessing the epidemiological impacts of isolating young and elder persons. Mathematical Medicine and Biology, 38(2):137–177, June 2021. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
38/2/137/6090329.

Yang:2020:MCP

[YLS^{+}20] Yingqiao Yang, Kewei Li, Gerhard Sommer, Kai-Leung Yung, and Gerhard A. Holzapfel. Mechanical characterization of porcine liver properties for computational simulation of indentation on cancerous tissue. Mathematical Medicine and Biology, 37(4):469–490, December 2020. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/37/4/469/5829769.

Young:1990:LE

[You90] Alan Young. The landfill ecosystem. Mathematical Medicine and Biology, 7(3):199–217, ???? 1990. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://

academic.oup.com/imammb/article/7/3/199/733197.

Young:2014:MMI

[YRCU14] Jennifer Young, Béatrice Rivière, Charles S. Cox, Jr., and Karen Uray. A mathematical model of intestinal oedema formation. Mathematical Medicine and Biology, 31(1):1–15, March 2014. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/31/1/1/860020.

Zaric:2002:DRA

[ZB02] Gregory S. Zaric and Margaret L. Brandeau. Dynamic resource allocation for epidemic control in multiple populations. Mathematical Medicine and Biology, 19(4):235–255, December 2002. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
19/4/235/753203.

Zemskov:2011:OPF

[ZEM11] E. P. Zemskov, I. R. Epstein, and A. Muntean. Oscillatory pulses in FitzHugh–Nagumo type systems with cross-diffusion. Mathematical Medicine and Biology, 28(2):217–226, June 2011. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
28/2/217/679766.

Zhang:2022:DSH

[Zha22] Wenjing Zhang. Deterministic and stochastic in-host tuberculosis models for bacterium-directed and host-directed therapy combination. Mathematical Medicine and Biology, 39(2): 126–155, June 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/39/2/126/6540670.

Zierenberg:2013:AMP

[ZHF^{+}13] Jennifer R. Zierenberg, David Halpern, Marcel Filoche, Bernard Sapoval, and James B. Grotberg. An asymptotic model of particle deposition at an airway bifurcation. Mathematical Medicine and Biology, 30(2):131–156, June 2013. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
30/2/131/890418.

Zhang:2019:MCD

[ZJJA19] Wenjing Zhang, Sophia Jang, Colleen B. Jonsson, and Linda J. S. Allen. Models of cytokine dynamics in the inflammatory response of viral zoonotic infectious diseases. Mathematical Medicine and Biology, 36(3):269–295, September 2019. CO-DEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (elec-

tronic). URL http://academic.oup.com/imammb/article/
36/3/269/5046615.

Zhong:2019:MMG

[ZKB^{+}19] L. Zhong, C. F. Ketelaar, R. J. Braun, C. G. Begley, and P. E. King-Smith. Mathematical modelling of glob-driven tear film breakup. Mathematical Medicine and Biology, 36 (1):55–91, March 2019. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/36/1/55/4828068.

Zhang:2003:EPT

[ZLC03] Yujuan Zhang, Bing Liu, and Lansun Chen. Extinction and permanence of a two-prey one-predator system with impulsive effect. Mathematical Medicine and Biology, 20(4):309– 325, December 2003. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http://academic.oup.

com/imammb/article/20/4/309/714029.

Zarzoso:1997:MFE

[ZNB97] V. Zarzoso, A. K. Nandi, and E. Bacharakis. Maternal and foetal ECG separation using blind source separation methods. Mathematical Medicine and Biology, 14(3):207–225, September 1997. CODEN MMBABK. ISSN 1477-8599 (print), 1477-

8602 (electronic). URL http://academic.oup.com/imammb/
article/14/3/207/2874291.

Zanca:2022:HQD

[ZOZ^{+}22] Adriana Zanca, James M. Osborne, Sophie G. Zaloumis, Carolina D. Weller, and Jennifer A. Flegg. How quickly does a wound heal? Bayesian calibration of a mathematical model of venous leg ulcer healing. Mathematical Medicine and Biology, 39(4):313–331, December 2022. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic). URL http:

//academic.oup.com/imammb/article/39/4/313/6607792.

Zhu:2008:IDC

[ZZ08] Huiyan Zhu and Xingfu Zou. Impact of delays in cell infection and virus production on HIV-1 dynamics. Mathematical Medicine and Biology, 25(2):99–112, June 2008. CODEN MMBABK. ISSN 1477-8599 (print), 1477-8602 (electronic).

URL http://academic.oup.com/imammb/article/25/2/99/
751914.